### v1

Abstract
We analyze the convergence properties of gradient descent algorithms on Riemannian manifolds. We study randomization of the tangent space directions of Riemannian gradient flows for minimizing smooth cost functions (of Morse–Bott type) to obtain convergence to local optima. We prove that through randomly projecting Riemannian gradients according to the Haar measure, convergence to local optima can be obtained almost surely despite the existence of saddle points. As an application we consider ground state preparation through quantum optimization over the unitary group. In this setting one can efficiently approximate the Haar-random projections by implementing unitary 2-designs on quantum computers. We prove that the respective algorithm almost surely converges to the global minimum that corresponds to the ground state of a desired Hamiltonian. Finally, we discuss the time required by the algorithm to pass a saddle point in a simple two-dimensional setting.
keywords:
Riemannian gradient system, randomized gradient descent, Morse–Bott systems, saddle-point escape, quantum optimization
pacs:
MSC Classification] 65K10, 53B21, 37H10
1 Introduction
Both gradient systems and numerical linear algebra come with a long history as well as a vibrant research activity. For instance, gradient systems have been advanced lately from Hilbert manifolds [47] to general metric spaces (e.g., in the context of optimal mass transport [2, 45]). In contrast, numerical linear algebra has faced factorization problems of huge tensor structures [6] (tensor SVD) using “randomization strategies” for their classical algorithms in order to handle these amounts of data. Only a few decades ago a systematic study of their interplay started by analyzing the QR-algorithm and interior point methods from a (Riemannian) geometric point of view, cf. the collection of [11]. In this development the seminal work of [12, 13] on the so-called “double-bracket flow” on orbits of the orthogonal group (and likewise of the unitary group in [14]), followed by independent similar ideas of [16], has sparked a plethora of applications. For an early overview on optimization techniques on Riemannian manifolds (including higher-order methods) see [56] in [11]. A mathematical account in all detail can be found in the monograph by [28] including discretization schemes, suitable step sizes, and proofs of convergence. For the latter, the authors heavily exploited the Morse-Bott structure of the cost function (i.e. the particular stucture of the Hessian at possibly non-isolated critical points, cf. Subsec. 2.3 and [22] to obtain convergence to a single critical point. Based on these developments, higher-order methods with applications to further cost functions on classical matrix manifolds are treated in [1]. In [54] these ideas have been extended to flows and their discretization schemes on more general homogeneous spaces. There they were also connected to applications in quantum dynamics as well as to applications in numerical (multi-)linear algebra like higher-rank tensor approximations—as elaborated by [19] to complement standard methods (like higher-order powers of [34] or quasi Newton methods of [53]).
Independently, in the physics community, [62] (re)devised gradient flows of Brockett type, e.g., to (band-)diagonalize Hamiltonians, which the monograph by [31] elaborated to address further quantum many-body applications.
Often gradient-flow approaches for deriving optimization schemes on Riemannian manifolds hinge on the computability of the Riemannian exponential ensuring to take the gradient from the tangent space back on the manifold. To be precise, this is crucial whenever the manifold cannot be identified with its tangent spaces such as for the unitary group or unitary orbits.
With these stipulations, Riemannian optimization techniques constitute a versatile toolbox, since a cost function can readily be tailored to the particular application such as finding extremal eigen- or singular values (i.e. problems otherwise addressed by variational approaches). These instances also connect to other branches of mathematics such as numerical and $C$-numerical ranges and their extremal values, the numerical and $C$-numerical radii (see [27, 38]).
Over the last decade, techniques from Riemannian optimization have also been adopted in quantum information science. For example, in quantum computing, hybrid quantum-classical algorithms [8], such as variational quantum algorithms (VQAs) [15, 59], have been developed to solve ground state problems [49] (i.e. typically minimal eigenvalue problems), combinatorial optimization [23], and quantum machine learning problems [9]. In VQAs a fixed parameterized quantum circuit, i.e. a parameterized set of unitary transformations (usually implemented by a fixed set of unitary gates), is iteratively optimized in tandem with a classical computer to minimize a cost function whose global optima encode the solution(s) to the desired problem. More precisely, a quantum processor, including a measurement device, is used to output the current (e.g., expectation) value of the cost function in each iterative step. Interleaved, the classical computer component provides the optimization routine that determines the parameter update for the next iteration on the quantum processor. Since quantum computers can, in principle, handle some classically exponentially complex problems, the hope is that VQAs will be able to solve problems of large size a solely classical algorithm would be incapable of.
However, since the optimization problems VQAs aim to solve are typically non-convex, which can often be traced back to the quantum circuit parameterization [41, 36], the iterative search for the optimal parameters can get stuck in suboptimal solutions [10].
To overcome this challenge, adaptive quantum algorithms have been designed. Instead of fixing a quantum circuit and picking a parameterization as in VQAs, adaptive algorithms successively grow a quantum circuit informed by measurement data from the quantum processor. In this approach the cost function is minimized while the quantum circuit is grown, see [25, 63, 42, 43]. One of the most prominent examples of such a strategy is the so-called ADAPT-VQE algorithm [25], which was originally proposed to solve ground state problems in chemistry, and was further developed later to solve combinatorial optimization problems on quantum computers [58].
Unfortunately, since typically in adaptive quantum algorithms the circuit growth is informed by gradient estimates performed by the quantum computer, adaptive quantum algorithms face similar challenges as VQAs, i.e., the adaptive circuit growth can get stuck when gradients vanish. This issue has recently been addressed by identifying adaptive quantum algorithms as quantum computer implementations of Riemannian gradient flows on the special unitary group $\text{SU}(d)$ of dimension $d=2^{n}$ where $n$ is the number of qubits of the quantum computer [63]. Since implementing Riemannian gradient methods on quantum computers requires in general exponential resources (i.e., the number of gates and/or the number of iterations of the corresponding quantum circuit growths exponentially in $n$). [63] proposed projecting the Riemannian gradient into smaller dimensional subspaces that scale polynomially in $n$, which in turn yields scalable quantum computer implementations. However, the proposed “dimension reduction” strategy of the Riemannian gradient to obtain efficient quantum computer implementations comes with a cost: While it can be shown that in the quantum computing setting (i.e., the Riemannian gradient flow on the special unitary group for grounds state problems) the full gradient flow converges almost surely to the problem solution given by the global minimizer, despite the existence of strict saddle points (i.e., saddle points at which the Hessian has at least one negative eigenvalue) [37], for the projected versions no convergence guarantees can be made. Indeed, numerical simulations suggest that existing versions of gradient algorithms that are efficiently implementable on quantum computers also suffer from converging to suboptimal solutions. Inspired by the success of classical randomized gradient descent [52], this problem has most recently been addressed by some of the authors in [43] via “random projections” of the Riemannian gradient, as depicted in Fig. 1, which are efficiently implementable on quantum computers. An efficient quantum computer implementation of these random directions can be achieved through exact and approximate unitary 2-designs [21]. The numerical experiments presented in [43] indicate that such a randomization procedure of the Riemannian gradient indeed yields (almost surely) convergence to the problem solution.
In this paper we consider the randomized gradient algorithm presented in [43] in the more general setting of Morse–Bott cost functions on Riemannian manifolds. For both continuous and discrete probability distributions of the random gradient directions we prove that the algorithm almost surely converges to a local minimum even. Finally we show how the result can be applied to quantum optimization tasks, such as ground state preparation. In the quantum setting there are no local optima, thereby we prove that the randomized gradient algorithm converges almost surely to the global minimum.
Figure 1: Schematic representation of the randomized gradient descent algorithm on a Riemannian manifold $\mathcal{M}$. At a point $X_{k}$ on $\mathcal{M}$ the Riemannian gradient $\operatorname{grad}J(X_{k})$ of a cost function $J:\mathcal{M}\to\mathbb{R}$ lies in the tangent space $T_{X_{k}}\mathcal{M}$ depicted as plane spanned by $\Omega_{k}$ and its orthocomplement $\Omega_{k}^{\perp}$. Instead of iteratively following the full gradient in order to optimize $J$, here in each step we consider a projection of $\text{grad}J(X_{k})$ onto a randomly chosen tangent-space direction $\Omega_{k}$. This is key for efficient implementations of gradient flows in large-scale quantum optimization problems. In the case of the special unitary group the randomized directions are given by (traceless) skew-Hermitian operators $\Omega_{k}=iH_{k}$. The corresponding quantum circuit consists of a sequence of unitary transformations as shown in Fig. 2.
Figure 2: Schematic implementation of the discretized Riemannian gradient scheme by a quantum algorithm in case the manifold is $\mathcal{M}=SU(2^{n})$ or $\mathcal{M}=Ad_{SU(2^{n})}$. For both instances the dimension of the problem scales exponentially with the number $n$ of quantum particles (here qubits). Each quantum circuit $U_{k}$ is adaptively turned into $U_{k+1}=\exp\left(\frac{-i\eta\,dJ(0)}{d\theta_{k}}H_{k}\right)U_{k}$ by moving in a random tangent-space direction $iH_{k}$ whose corresponding projection $\frac{dJ_{k}(0)}{d\theta_{k}}=\langle\text{grad}J(U_{k}),iH_{k}U_{k}\rangle$ is measured on a quantum device, where $\eta$ denotes the learning rate and we define the cost-function value at the $k$-th iteration as $J_{k}(x)\equiv J(x,U_{k})$ where $J(x,U_{k})=\text{tr}\{A\exp(-ixH_{k})U_{k}\rho U_{k}^{*}\exp(ixH_{k})\}$ (see Sec. 5). The measurement projection is done on the quantum computer, and the statistics of repeated measurements for the expectation value is done on the classical computer. In this way a randomized gradient flow as in Fig. 1 is implemented. In this work we show that for sufficiently small $\eta$ such a (discretized) flow converges almost surely to the global optimum.
Structure and Main Results
In Sec. 2 we introduce the fundamental notions from Riemannian geometry to describe gradient flows (Sec. 2.1), recollecting some well-known results on the convergence of gradient flows (Sec. 2.2), and their algorithmic implementation as gradient descents (Sec. 2.4). In Sec. 3 we describe a gradient descent algorithm where in each step the full gradient of a high-dimensional problem is projected on just a randomly chosen direction of the tangent space, considering both continuous and discrete distributions.
The main main result is obtained in Secs. 3.2 and 3.3, where we prove that a randomly projected gradient descent algorithm to a smooth Morse–Bott function (with compact sublevel sets) on a Riemannian manifold comes with two vitally beneficial properties: (i) it almost surely escapes saddle points—as is shown in Lem. 3.11 and as a consequence (ii) it almost surely converges a local minimum—as shown in Thm. 3.14.
In Sec. 4 the case of a two-dimensional saddle point is studied approximately using analytical methods and also using numerical simulation. We obtain good approximations for the time necessary to pass the saddle point.
Finally Sec. 5 shows how the previous results can be applied in quantum optimization to the problem of ground state preparation.
2 Riemannian Gradient Flow and Descent
2.1 Basic Concepts
First we recall some basic notions and notations from Riemannian geometry. Let $M$ denote a finite dimensional smooth manifold of dimension $n$ with tangent and cotangent bundles ${\sf T}M$ and ${\sf T}^{*}M$, respectively. Moreover, let $M$ be endowed with a Riemannian metric $g$, i.e. a smoothly varying scalar product $g(\mathbf{x})=\langle\,\cdot,\cdot\,\rangle_{\mathbf{x}}$ on each tangent space ${\sf T}_{\mathbf{x}}M$, $\mathbf{x}\in M$. The pair $(M,g)$ is called a Riemannian manifold.
Every Riemannian manifold is equipped with a Riemannian density $\mu$ induced by the metric $g$, see [35, Prop. 2.44]. This density can be used to integrate functions, to induce a measure on $M$ and to define sets of measure zero. Equivalently (but without introduction a Riemannian density), one could say that $S\subset M$ has measure zero if there is an atlas such that $S\cap{\rm dom}\,\sigma$ has Lebesgue measure zero in every chart $\sigma$. In particular any submanifold of dimension strictly smaller than $N$ has measure zero.
2.2 Gradient Flows and Asymptotic Behaviour (I)
In the following general description, we depart from the above notation of ‘the’ cost function $J$ borrowed from quantum optimization and introduce $f:M\to\mathbb{R}$ for an arbitrary smooth (cost) function on $M$ with differential $df:M\to{\sf T}^{*}M$. Then the gradient of $f$ at $\mathbf{x}\in M$, denoted by $\operatorname{grad}f(\mathbf{x})$, is the unique vector in ${\sf T}_{\mathbf{x}}M$ determined by the identity
$df(\mathbf{x})\xi=\langle\operatorname{grad}f(\mathbf{x}),\xi\rangle_{\mathbf{x}}$ (1)
for all $\xi\in{\sf T}_{\mathbf{x}}M$. Equation (1) naturally defines a vector field on $M$ via $\operatorname{grad}f:M\to{\sf T}M$, $\mathbf{x}\to\operatorname{grad}f(\mathbf{x})$ called gradient vector field of $f$. The corresponding ordinary differential equation
$\dot{\mathbf{x}}=-\operatorname{grad}f(\mathbf{x})=:F(\mathbf{x})$ (2)
and its flow are referred to as the gradient system and the gradient flow of $f$, respectively.
In local coordinates the above reads as follows: Let $\sigma:U\to\mathbb{R}^{n}$ be a chart about $\mathbf{x}\in M$, denote $\mathbf{y}=\sigma(\mathbf{x})$, and let $\tilde{f}:=f\circ\sigma^{-1}$ be the chart representation of $f$. Moreover let $\tilde{g}:=(\sigma^{-1})^{*}g$ be the chart representation of the metric $g$ on $\sigma(U)$, i.e. $\tilde{g}(\mathbf{y})=\sum_{i,j=1}^{n}\tilde{g}_{ij}(\mathbf{y})\,dy^{i}dy^{j}$, where $\tilde{G}(y):=\big{(}\tilde{g}_{ij}(\mathbf{y})\big{)}$ is a positive definite matrix varying smoothly in $\mathbf{y}\in\sigma(U)$. Finally let $\tilde{F}:=(\sigma^{-1})^{*}F$ be the chart representation of the gradient vector field $\operatorname{grad}f$. Then it holds that
$\tilde{F}^{j}(\mathbf{y})=\sum_{i=1}^{n}\tilde{g}^{ij}(\mathbf{y})\frac{\partial\tilde{f}}{\partial y_{i}}(\mathbf{y}),$
where $\tilde{g}^{ij}(\mathbf{y})$ are the entries of the inverse of $\tilde{G}(\mathbf{y})$ and $\frac{\partial}{\partial y_{i}}$ denotes the $i$-th partial derivative.
Any convergence analysis of gradient systems is based on the following two observations: (i) the critical points of $f$ are the equilibria of (2), (ii) $f$ constitutes a Lyapunov function for (2), i.e. $f$ is monotonically decreasing along solutions. Despite their simplicity they immediately allow several non-trivial statements about the asymptotic behavior of (2) which can be characterized by its $\omega$-limit sets
$\omega(\mathbf{x}_{0}):=\bigcap_{s>0}\overline{\big{\{}\varphi(t,\mathbf{x}_{0})\;|\;t>s\big{\}}}\,,$
where $t\mapsto\varphi(t,\mathbf{x}_{0})$ denotes the unique solution of (2) with initial value $\varphi(0,\mathbf{x}_{0})=\mathbf{x}_{0}$.
Proposition 2.1 ([35]).
If f has compact sublevel sets, i.e. if the sets $\{\mathbf{x}\in M:f(\mathbf{x})\leq c\}$ are compact for all $c\in\mathbb{R}$, then every solution of (2) exists for $t\geq 0$ and its $\omega$-limit set is a non-empty, compact and connected subset of the set of critical points of $f$.
Although, Proposition 2.1 shows that $\omega(\mathbf{x})$ is contained in the set of critical points of $f$ it does not guarantee convergence to a single critical point and, indeed, there are smooth gradient systems the trajectories of which exhibit a non-trivial convergence behavior to the set of critical points (cf. “Mexican hat counter-example” by [18]). Yet for isolated critical points one has the following trivial consequence of Proposition 2.1.
Corollary 2.2.
If f has compact sublevel sets and all critical points are isolated, then any solution of (2) converges (for $t\to\infty$) to a single critical point of $f$. Moreover if additionally all saddle points are strict111see Subsec. 2.3 below then for almost all initial points (in the sense of Sec. 2.1) the flow converges to a local minimum.
Continua of critical points are much more subtle to handle. Some enhanced conditions guaranteeing convergence to a single critical point will be briefly discussed in the next subsection.
2.3 The Hessian, Morse–Bott Functions and Asymptotic Behaviour (II)
As in the Euclidian case, linearizing the vector field $F$ at its equilibria sheds light on its local stability. Clearly, since $F=\operatorname{grad}f$ constitutes a gradient vector field, its linearization at an equilibrium $\mathbf{x}\in M$ is given by the Hessian $\operatorname{H}_{f}(\mathbf{x})$ of $f$ at $\mathbf{x}\in M$. In general, if $M$ is non-Euclidian, the computation of $\operatorname{H}_{f}(\mathbf{x})$ can be rather involved. Yet, at critical points $\mathbf{x}_{0}\in M$ of $f$, the Hessian is given by the symmetric bilinear form
$\operatorname{H}_{f}(\mathbf{x}_{0})(\xi,\eta):=\sum_{i,j=1}^{n}\operatorname{\tilde{H}}(f\circ\sigma^{-1})\big{(}\sigma(\mathbf{x}_{0})\big{)}_{ij}(D\sigma(\mathbf{x}_{0})\xi)_{i}(D\sigma(\mathbf{x}_{0})\eta)_{j}\,,$ (3)
where $\sigma$ is any chart around $\mathbf{x}_{0}\in M$ and $\operatorname{\tilde{H}}(f\circ\sigma^{-1})$ denotes the ordinary Hesse matrix of the chart representation $f\circ\sigma^{-1}$. It is straightforward to show that (3) is independent of $\sigma$. Moreover, we will call $\mathbf{x}_{0}$ a strict saddle point if $\operatorname{H}_{f}(\mathbf{x}_{0})$ (or more precisely the associated symmetric operator) has at least one negative eigenvalue.
The above concepts allow a trivial generalization of a well-known result from elementary calculus which follows straightforwardly in local coordinates.
Proposition 2.3.
Let $(M,g)$ be a Riemannian manifold and let $\mathbf{x}$ be a critical point of $f:M\to\mathbb{R}$. If $\operatorname{H}_{f}(\mathbf{x})$ is positive definite, then $\mathbf{x}$ is a strict local minimum of $f$.
Now the question arises whether the (asymptotic) stability of an equilibrium $\mathbf{x}\in M$ of (2) may dependent on the Riemannian metric $g$ — and the answer is surprisingly “yes”, cf. [57]. However, certain properties such as being a strict local minimum or an isolated critical point are obviously not up to the choice of the metric, and thus the (asymptotic) stability of those equilibria is independent of the Riemannian metric as the next theorem shows.
Theorem 2.4.
  1. (a)
    Every strict local minimum of $f$ is a stable equilibrium of (2).
  2. (b)
    Every strict local minimum of $f$ which is additionally an isolated critical point is an asymptotically stable equilibrium of (2).
Both assertions follow immediately from classical stability theory by taking $f$ as Lyapunov function, cf. [28, 29]. Handling non-isolated critical points is much more subtle. A first hint on how to approach this issue is obtained by Corollary 2.2 which could be restated (in a slightly weaker version) as follows:
For Morse functions with compact sublevel sets every solution of the corresponding gradient system converges (for $t\to\infty$) to a single critical point.
This suggests to work with Morse–Bott functions when it comes to non-isolated critical point. Recall that a smooth function $f:M\to\mathbb{R}$ is a called Morse–Bott function, when the set $C$ of critical points is a closed submanifolds of $M$ such that the tangent space of ${\sf T}_{\mathbf{x}}C$ coincides with the kernel of the Hessian operator for all $\mathbf{x}\in C$. Note that $C$ is allowed to have several connected components with possibly different dimensions, [28, p. 366], [48, Def. 2.41] or [5]. Thus Morse functions are particular Morse-Bott functions, where $C$ consists of 00-dimensional mainfolds (i.e. isolated points).
Now, the above concept allows a generalization of the Morse-Palais Lemma to Morse-Bott functions which is often called Morse-Bott Lemma, see [4]. It yields a local normal form of Morse-Bott functions near their critical points, which reads in local coordinates as follows
$\tilde{f}(\mathbf{x},\mathbf{y},\mathbf{z})=\|\mathbf{x}\|^{2}-\|\mathbf{y}\|^{2}\,,$
with $\mathbf{x}\in\mathbb{R}^{n_{+}}$, $\mathbf{y}\in\mathbb{R}^{n_{-}}$, and $\mathbf{z}\in\mathbb{R}^{n_{0}}$, where $n_{+}$, $n_{-}$ and $n_{0}$ are the number of positive, negative and zero eigenvalues of the Hessian operator. This finally allows to prove that solutions of the respective gradient systems convergence to a single critical point.
Theorem 2.5.
Let $f:M\to\mathbb{R}$ be a Morse–Bott function on a Riemannian manifold $(M,g)$ with compact sublevel sets. Then every solution of the gradient flow (2) converges to a single critical point. Moreover for almost every initial point the flow converges to a local minimum.
Finally, we recall another very powerful result for analyzing the convergence of gradient systems which is based on Łojasiewicz’s celebrated gradient estimate [39]. Let $f:M\to\mathbb{R}$ be real analytic, ${\mathbf{x}}_{0}\in M$ a critical point of $f$ and assume w.l.o.g. $f({\mathbf{x}}_{0})=0$. Then near ${\mathbf{x}}_{0}\in M$ one has the estimate
$\|\operatorname{grad}(\mu\circ f)({\mathbf{x}})\|_{\mathbf{x}}\geq c\,,$ (4)
where $\mu:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a strictly increasing $C^{1}$-function and $c>0$ some positive constant, cf. [33, Cor. 1.1.25]. In the literature, one usually find $\mu(r):=r^{1-\theta}$ with $\theta\in(0,1)$. Eq. (4) easily allows to bound the length of any trajectory of (2) whose $\omega$-limit set is non-empty. Hence one gets the following result.
Theorem 2.6.
[39]. Let $(M,g)$ and $f:M\to\mathbb{R}$ be are real analytic. Then every non-empty $\omega$-limit set of (2) consists only of a single critical point.
2.4 The Exponential Map and Numerical Gradient Descent
Finally, we approach the problem of discretization of (2) resulting in a convergent gradient descent method. The ideas presented here can be traced back to [12, 13] and [55, 56]. Let
$\exp_{\mathbf{x}}:{\sf T}_{\mathbf{x}}M\to M$ (5)
denote the Riemannian exponential map at ${\mathbf{x}}\in M$, i.e. $t\to\exp_{\mathbf{x}}(t\xi)$ is the unique geodesic with initial value ${\mathbf{x}}\in M$ and initial velocity $\xi\in{\sf T}_{\mathbf{x}}M$. Here, we assume for simplicity that $(M,g)$ is (geodesically) complete, i.e. (5) is well-defined for the entire tangent space ${\sf T}_{\mathbf{x}}M$. Probably the simplest discretization scheme given by
${\mathbf{x}}_{k+1}:=\exp_{{\mathbf{x}}_{k}}(-\eta_{k}\operatorname{grad}f({\mathbf{x}}_{k})),$ (6)
can be seen as a “natural” generalization of the explicit Euler method. Here $\eta_{k}>0$ denotes an appropriate “step size” [^2] Note that the “actual” step size results form the modulus of $\eta\operatorname{grad}f({\mathbf{x}}_{k})$. which may depend on $k\in\mathbb{N}$.
In order to guarantee convergence of (6) to the set of critical points, it is sufficient to apply the Armijo rule, see [40]. An alternative to Armijo’s rule provides the step-size selection suggested by [14], see also [28]. Moreover, for compact Riemannian manifolds even a sufficiently small constant step size $\eta>0$ guarantees convergence:
Theorem 2.7.
For compact Riemannian manifolds every trajectory of the discretized gradient descent (6) (with constant but small enough step size) converges to the set of critical points. Moreover, the set of initial points converging to strict saddle points has measure zero—in the sense of Sec. 2.1 here.
The second part of the above statement can be found in [37, Cor. 6] [^3] There the authors consider Riemannian manifolds embedded in the Euclidean space and retractions instead of an intrinsic Riemannian exponential function, but this does not restrict the generality of the result, cf. the Nash embedding theorem [7, Thm. 46]. Thus, for a function with compact sublevel sets, gradient descent (6) behaves similar as its continuous counter-part, cf. Cor. 2.2, i.e. it converges almost surely to a local minimum if the step size is chosen small enough.
Deeper results which yield convergence to a single critical point are more subtle to derive. Here we present one result in this direction which is again based on the analyticity of the cost function $f$ and on Łojasiewicz’s inequaliy.
Theorem 2.8.
[33] If $(M,g)$ and $f$ are real analytic, and the step sizes are chosen according to a version of the first Wolfe–Powell condition for Riemannian manifolds, then pointwise convergence holds.
Remark 2.9.
It should be mentioned that gradient descent algorithms are usually studied as exact algorithms, not as numerical algorithms where real numbers are represented as floating point values and arithmetic is not exact. Numerical effects can be important, for instance cancellation effects from computing gradients in a naive way. Nevertheless, this paper will ignore these numerical issues.
Finally, in order to determine the largest admissible step size of our algorithm we need a notion of Lipschitz continuity of vector fields on Riemannian manifolds. Care has to be taken here since tangent vectors in different tangent spaces cannot be compared by default. Certainly, one can define local Lipschitz continuity via local charts but this does not allow to assign a meaningful Lipschitz constant to the vector field. A natural and intrinsic way to do this is to define Lipschitz continuity via parallel transport along the unique connecting geodesic, as is done in [24]. Here, we use the following equivalent approach via Riemannian normal coordinates which result from any orthogonal coordinate system on $T_{\mathbf{x}}M$ in combination with the Riemannian exponential map $\exp_{\mathbf{x}}$.
Definition 1.
Let $(M,g)$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be a smooth function. We say that $f$ is $\ell$-smooth around $\mathbf{x}\in M$ if there exist normal coordinates $\tilde{f}$ around $\mathbf{x}$ in which $\nabla\tilde{f}$ is $\ell$-Lipschitz.
3 Algorithm and Convergence
We will study the following randomly projected gradient descent algorithm. Given a Riemannian manifold $M$ of dimension $N$, a smooth function $f:M\to\mathbb{R}$, and a step size $\eta>0$ (with upper bound to be determined), the update rule is given by
$\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})),\quad g(\mathbf{x},\mathbf{u})=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle\,\mathbf{u},\quad\mathsf{S}_{\mathbf{x}_{i}}M\ni\mathbf{u}_{i}\overset{\mathrm{i.i.d}}{\sim}\mathcal{P}(\mathbf{x}_{i}),$ (7)
where $\mathsf{S}_{\mathbf{x}}M$ denotes the unit sphere in the tangent space $\mathsf{T}_{\mathbf{x}}M$ and $\mathcal{P}(\mathbf{x}_{i})$ some probability measure on the unit sphere $\mathsf{S}_{\mathbf{x}_{i}}M$. Intuitively, the gradient is projected onto a randomly chosen direction $\mathbf{u}$ at each step. Throughout we will consider two cases.
  • Either $\mathcal{P}(\mathbf{x}_{i})=\mathcal{U}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes the uniform (rotation invariant) probability distribution on the unit sphere, also called Haar measure;
  • or $\mathcal{P}(\mathbf{x}_{i})=\mathcal{D}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes a finite probability distribution on the unit sphere in some sense approximating a uniform distribution.
Let us clarify the case of a finite probability distribution. In this case we have $k$ continuous vector fields $\xi_{j}$ on $M$ and continuous weights $p_{j}:M\to\mathbb{R}$ with $p_{j}>0$ and $\sum_{j_{1}}^{k}p_{j}\equiv 1$. Then $\mathbf{u}_{i}$ takes value $\xi_{j}(\mathbf{x}_{i})$ with probability $p_{j}(\mathbf{x}_{i})$. Moreover we assume that at every point $\mathbf{x}$, the the tangent vectors $\xi_{1}(\mathbf{x}),\dots,\xi_{k}(\mathbf{x})$ span the entire tangent space even after removing all vectors collinear with some given direction $\xi_{j}(\mathbf{x})$.
Our goal is to analyze the convergence behaviour of this algorithm, and in particular to show that it converges almost surely to a local minimum of $f$. A deterministic version of the above described algorithm was analysed in [1, Def. 4.2.1, Thm. 4.3.1] under the heading “gradient-related methods”.
3.1 Basic Properties
We start with some simple properties of the function $g$ defined in (7).
Lemma 3.1.
Let $\mathbf{x}\in M$ be given. If $\mathbf{u}\in\mathsf{S}_{\mathbf{x}}M$, then it holds that
$\langle\operatorname{grad}f(\mathbf{x}),g(\mathbf{x},\mathbf{u})\rangle=\|g(\mathbf{x},\mathbf{u})\|^{2}.$
Proof.
Using the definition of $g(\mathbf{x},\mathbf{u})$ and the fact that $\mathbf{u}$ is a unit vector we immediately obtain $\langle\operatorname{grad}f(\mathbf{x}),g(\mathbf{x},\mathbf{u})\rangle=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle^{2}=\|g(\mathbf{x},\mathbf{u})\|^{2}.$
Note that if $\mathbf{u}$ is uniformly distributed on the sphere, then
[…]
Remark 3.2.
The image of the function $u\mapsto g(\mathbf{x},\mathbf{u})$ is a hypersphere in $\mathsf{T}_{\mathbf{x}}M$ with center $\tfrac{1}{2}\operatorname{grad}f(\mathbf{x})$ and containing the origin. Let $\mathbf{x}\in M$ be fixed and consider the map $g_{\mathbf{x}}:\mathsf{S}_{\mathbf{x}}M\to\mathsf{T}_{\mathbf{x}}M$ given by $\mathbf{u}\mapsto g(\mathbf{x},\mathbf{u})$. To simplify formulas, we choose an orthonormal basis $(e_{1},\ldots,e_{N})$ in $\mathsf{T}_{\mathbf{x}}M$ such that $\operatorname{grad}f(\mathbf{x})=\|\operatorname{grad}f(\mathbf{x})\|e_{N}$. In these coordinates the map $g_{\mathbf{x}}$ is given by
$g_{\mathbf{x}}(u_{1},\ldots,u_{N})=u_{N}\|\operatorname{grad}f(\mathbf{x})\|\,(u_{1},\ldots,u_{N}).$
A straightforward computation shows that $\|\,(u_{N}u_{1},\ldots,u_{N}^{2}-\tfrac{1}{2})\|=\tfrac{1}{2},$ and so the image of $g_{\mathbf{x}}$ lies on a sphere with center $\tfrac{1}{2}\operatorname{grad}f({\mathbf{x}})$ and passing through the origin. This induces a probability measure on the image. Consider the function
$\tilde{g}({\mathbf{x}},\mathbf{v})=\tfrac{1}{2}(\operatorname{grad}f({\mathbf{x}})+\mathbf{v}\|\operatorname{grad}f({\mathbf{x}})\|),$
and let $v$ be distributed on the unit sphere according to a probability measure $\mathcal{V}(\mathsf{S}_{x}M)$ such that $g$ and $\tilde{g}$ induce the same probability measure. The measure $\mathcal{V}$ is not uniform on the sphere, but it is still invariant under rotations preserving $\operatorname{grad}f({\mathbf{x}})$. We see that $v_{N}\sim 2u_{N}^{2}-1$.
Note that since the standard deviation of $u_{N}^{2}$ is larger than the expected value, Chebyshev’s inequality cannot be applied to obtain useful concentration bounds. We will use a different method in Lemma A.2.
Corollary 3.3.
If $\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}]=\mathds{1}/N$, which is satisfied for the Haar measure, then it holds that
$\mathbb{E}[g(\mathbf{x},\mathbf{u})]=\frac{\operatorname{grad}f(\mathbf{x})}{N}$
for any $\mathbf{x}\in M$.
Proof.
This is a simple computation: $\mathbb{E}[g(\mathbf{x},\mathbf{u})]=\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}\operatorname{grad}f(\mathbf{x})]=\frac{1}{N}\operatorname{grad}f(\mathbf{x})$. ∎
The corollary shows that in this case the projected gradient is of size $\operatorname{grad}f(\mathbf{x})/N$, and hence it is rather small for large $N$. Intuitively this happens because in high dimensions, a uniformly random unit vector $u$ will be close to orthogonal to the gradient with high probability — after all, there are $N-1$ dimensions which are orthogonal to the gradient.
For large dimension $N$ there exist good approximations of these distributions. In fact, the distribution of a coordinate of a uniformly random unit vector approximately follows a normal distribution, and hence it’s square approximates a $\chi^{2}_{1}$ distribution. See Lemma A.2 for a precise result.
Now that we better understand the iteration rule, we want to understand by what amount the objective function value is likely to decrease after a certain number of iterations. It will be useful to use normal coordinate charts at the current point $\mathbf{x}_{i}$. These charts satisfy that the metric at the origin is trivial, and that geodesics passing through the origin are straight and uniformly parametrized, see for instance [35, Prop. 5.24]. In particular, when using a normal coordinate chart about $\mathbf{x}_{i}$, the random variable $\mathbf{x}_{i+1}$, conditioned on $\mathbf{x}_{i}$, is still distributed on a hypersphere.
Lemma 3.4.
Let $(M,g)$ be compact and let $f:M\to\mathbb{R}$ be $\ell$-smooth at $\mathbf{x}_{i}$ and with injectivity radius $\operatorname{inj}_{M}(\mathbf{x}_{i})$. Then for $\eta\leq\min(1/\ell,\operatorname{inj}_{M}(\mathbf{x}_{i}))$ it holds that
$f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\leq-\eta\Big{(}1-\frac{\ell\eta}{2}\Big{)}\langle\operatorname{grad}f(\mathbf{x}_{i}),g(\mathbf{x}_{i},\mathbf{u}_{i})\rangle\leq 0.$
In particular the function value cannot increase.
Proof.
We choose a normal coordinate chart about $\mathbf{x}_{i}$ and denote the coordinates by $\mathbf{\tilde{x}}$ and the function in coordinates by $\tilde{f}$. In particular $\mathbf{\tilde{x}}_{i}=0$. We may and do assume that $\nabla\tilde{f}$ is $\ell$-Lipschitz. Note that at the origin $\operatorname{grad}\tilde{f}(\mathbf{x}_{i})=\nabla\tilde{f}(\mathbf{\tilde{x}}_{i})$ and the coordinate representation of $\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i}))$ is given by $\mathbf{\tilde{x}}_{i+1}=-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})$. Then by [46, Lemma 1.2.3] and Lemma 3.1 it holds that
$f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})$ $=\tilde{f}(\mathbf{\tilde{x}}_{i+1})-\tilde{f}(\mathbf{\tilde{x}}_{i})$
$\leq\langle\nabla\tilde{f}(\mathbf{\tilde{x}}_{1}),\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\rangle+\frac{\ell}{2}\|\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\|^{2}$
$\leq-\eta\langle\operatorname{grad}f(\mathbf{x}_{i}),g(\mathbf{x}_{i},\mathbf{u}_{i})\rangle+\frac{\ell\eta^{2}}{2}\|g(\mathbf{x}_{i},\mathbf{u}_{i})\|^{2}$
$=-\eta\Big{(}1-\frac{\ell\eta}{2}\Big{)}\langle\operatorname{grad}f(\mathbf{x}_{i}),g(\mathbf{x}_{i},\mathbf{u}_{i})\rangle,$
as desired. ∎
This result shows that, under the assumptions of Corollary 3.3, we obtain
$\mathbb{E}[f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})]\leq-\frac{\eta}{N}\Big{(}1-\frac{\ell\eta}{2}\Big{)}\|\operatorname{grad}f(\mathbf{x}_{i})\|^{2}.$
More generally the previous result shows that the only way for the algorithm to stop improving is for the gradient to vanish: $\operatorname{grad}f(\mathbf{x}_{i})\to 0$.
Corollary 3.5.
If $f(\mathbf{x}_{i})\to c$ then $c$ is a critical value and $\mathbf{x}_{i}$ converges almost surely to the critical set of $c$.
Proof.
We will show that $\Pr(f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0\wedge\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0)=0$, which immediately implies that if the values $f(\mathbf{x}_{i})$ converge, then the gradients $\operatorname{grad}f(\mathbf{x}_{i})$ converge to 00 almost surely. Note that if $\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0$, then there is some $\varepsilon>0$ such that $\|\operatorname{grad}f(\mathbf{x}_{i})\|>\varepsilon$ on some infinite subsequence. Consider the case of the Haar measure. By the above, if $f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0$, then $u_{N}^{(i)}\to 0$ on the same subsequence. Since the $u_{N}^{(i)}$ are i.i.d (see Lemma A.1) and independent of the $\mathbf{x}_{i}$ the probability that this happens is 00, and this concludes the proof. The case of discrete probability distributions follows immediately form Lemma A.3. ∎
3.2 Escaping Saddle Points
The main difficulty in proving almost sure convergence of the randomly projected gradient descent algorithm is to show that it does not get stuck in a saddle point.
Recall from Lemma 3.4 that the function value cannot increase. Hence, once we reach the critical value corresponding to some saddle point, without sitting in the saddle point itself, we say that we have “passed” the saddle point as now it is impossible to converge to the saddle point in question.
We start by proving the result in a simplified “isotropic” case, before treating the general case as a perturbation.
Lemma 3.6.
Consider the function $f:\mathbb{R}^{n}\to\mathbb{R}$ defined by
$f(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{n-p-q})=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}),$
with $1\leq p,q$ and $p+q\leq n$ as well as $a_{i},b_{j}>0$. Then $f$ is $\ell$-smooth with $\ell=2\max_{i,j}(a_{i},b_{j})$. Moreover let $0<\eta\leq\tfrac{1}{\ell}$ and
$\theta=\arctan\Big{(}\frac{b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}}{a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}}\Big{)}\in\big{[}0,\tfrac{\pi}{2}\big{]}.$ (8)
Then there exist constants $\varepsilon,\delta>0$ such that for all  [^4] The value $\tfrac{\pi}{3}$ is arbitrary and could be replaced by any other value in $(0,\frac{\pi}{2})$ without changing the proof. $\theta\in[0,\tfrac{\pi}{3}]$ it holds that
$\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta.$
Proof.
The negative gradient of $f$ is
$,\displaystyle-\nabla f(x_{1},\ldots,x_{p},y_{1},$$\ldots,y_{q},z_{1},\ldots,z_{n-p-q})$
$=-2(a_{1}x_{1},\ldots,a_{p}x_{p},-b_{1}y_{1},\ldots,-b_{q}y_{q},0,\ldots,0)$
and the value of $\ell$ follows immediately. A key property of this simplified setting is that the gradient is a linear vector field on $\mathbb{R}^{n}$, and so the entire situation is invariant under scaling. From the definition of $\theta$ it is clear that $f$ vanishes if and only if $\theta=\tfrac{\pi}{4}$.
First note that due to the scaling invariance of the algorithm and of (8), we may focus on an initial point $\mathbf{x}_{i}$ on the unit sphere $S^{n-1}$. After executing one step, we obtain a random improvement of our angle, namely $\theta_{i+1}-\theta_{i}$ which depends on $\mathbf{u}_{i}$ appearing in the update rule. Considering this improvement as a function of $\mathbf{x}_{i}$ and $\mathbf{u}_{i}$, i.e.,
$h(\mathbf{x}_{i},\mathbf{u}_{i}):=\theta_{i}-\theta_{i},$
is is clear that it is a continuous function. The optimal improvement achievable for a given $\mathbf{x}_{i}$ is found by maximizing over $\mathbf{u}_{i}$. One can show that the resulting function is still continuous in $\mathbf{x}_{i}$. Moreover we claim that it is strictly positive on the sphere. Indeed, if $\theta_{i}>0$, then this follows from the choice of $\eta$. If we perform a single (deterministic) negative gradient step with step size $\eta$, none of the coordinates will change sign, and whenever $\theta\in(0,\tfrac{\pi}{2})$ its value will strictly increase. If $\theta_{i}=0$, it is clear that the value of $\theta$ will almost surely strictly increase, as the set of points $\mathbf{x}_{i+1}$ with $\theta_{i+1}=0$ form a strict subspace of $\mathbb{R}^{n}$. Taken together this shows that there is some $\varepsilon>0$ such that the optimal improvement is at least $2\varepsilon$ for every initial $\mathbf{x}_{i+1}$.
It follows that for any given initial state $\mathbf{x}_{i+1}$, the probability that $\theta_{i+1}-\theta_{i}\geq\varepsilon$ is strictly positive. Indeed, $\mathbf{x}_{i+1}$ takes values on a sphere, and $\theta(\cdot)$ is a continuous function on this sphere. The preimage of $(\varepsilon,+\infty)$ on the sphere is a non-empty open set and hence has a strictly positive probability. Moreover one can show that this probability depends continuously on the initial state $\mathbf{x}_{i}$, and again we obtain a positive minimum $\delta$ on the sphere by compactness. This proves that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$ independently of the initial state.
Now consider the discrete case. Again we need to prove that there exist constants $\varepsilon,\delta>0$ such that for all $\theta\in[0,\tfrac{\pi}{3}]$ it holds that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$.
The only cases where $h(\mathbf{x},\mathbf{u})=0$ (note that it cannot be negative) with probability one is when all $v_{j}(\mathbf{x})$ for $j=1,\ldots,k$ are either orthogonal to $\operatorname{grad}(\mathbf{x})$ or collinear to $\mathbf{x}$. By assumption, this cannot happen. Hence $\max_{j}h(\mathbf{x},\mathbf{v}_{j})$ is non-negative and continuous, and on the compact set $\{\mathbf{x}\in S^{n-1}:\theta(\mathbf{x})\leq\pi/3\}$ it is even strictly positive. If we denote the minimum by $\varepsilon$ and $\delta=\frac{1}{k}$, then the result follows. ∎
Corollary 3.7.
In the same setting as above, there exist $\varepsilon,\delta>0$ such that, with $N=\lceil\tfrac{\pi}{3\varepsilon}\rceil$, we have that
$\Pr(\theta_{N}\geq\tfrac{\pi}{3})\geq\delta^{N}.$
It follows that for any $m\in\mathbb{N}$
$\Pr(\theta_{mN}<\tfrac{\pi}{3})\leq\Pr(\theta_{N}<\tfrac{\pi}{3})^{m}\leq(1-\delta^{N})^{m},$
which goes to 00 as $m\to\infty$.
Proof.
It follows immediately from Lemma 3.6 that in $N$ steps we achieve $\Pr(\theta_{i+N}-\theta_{i}\geq N\varepsilon)\geq\delta^{N}$, or put differently, setting $N=\lceil\tfrac{\pi}{3\varepsilon}\rceil$, we get
$\Pr(\theta_{N}\geq\tfrac{\pi}{3})\geq\Pr(\theta_{N}\geq N\varepsilon)\geq\delta^{N}=\delta^{\lceil\tfrac{\pi}{3\varepsilon}\rceil}>0.$
This proves the first statement. The second statement follows from the Markovianity of the process. ∎
In order to generalize Corollary 3.7 to the desired setting we will treat the effects of the Riemannian metric and the higher order terms of $f$ as a perturbation on the algorithm. We start with the following general perturbation result.
Lemma 3.8.
Let $G$ be a linear vector field on $\mathbb{R}^{n}$ and let $\eta>0$ be given. Moreover let $P$ be a (not necessarily linear) vector field on $\mathbb{R}^{n}$ satisfying $\|P({\mathbf{x}})\|\leq C\|{\mathbf{x}}\|^{2}$ for some $C>0$. Then it holds that
$\|{\mathbf{x}}-\eta G{\mathbf{x}}-({\mathbf{y}}-\eta(G+P){\mathbf{y}})\|\leq\|\mathds{1}+\eta G\|\|{\mathbf{x}}-{\mathbf{y}}\|+C\eta\|{\mathbf{y}}\|^{2},$
where $\|\mathds{1}+\eta G\|$ denotes the operator norm. Hence after $n$ steps, if the states remain in an $R$-ball about the origin,
$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
Proof.
The first inequality follows from the assumptions using the triangle inequality:
$\|{\mathbf{x}}-\eta G{\mathbf{x}}-({\mathbf{y}}-\eta(G+P){\mathbf{y}})\|$ $=\|(\mathds{1}+\eta G)({\mathbf{x}}-{\mathbf{y}})+\eta P({\mathbf{y}})\|$
$\leq\|\mathds{1}+\eta G\|\|{\mathbf{x}}-{\mathbf{y}}\|+C\eta\|{\mathbf{y}}\|^{2}.$
Now using the assumption that $\|{\mathbf{y}}\|\leq R$, and by iterating this result for $n$ steps, we find that
$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|$ $\leq\|\mathds{1}+\eta G\|^{n}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+(\|\mathds{1}+\eta G\|^{n-1}+\cdots+1)CR^{2}$
$\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
This concludes the proof. ∎
In our case, we can always choose coordinates in which the problem locally looks like the setting of Lemma 3.6 with a perturbation as in Lemma 3.8.
Lemma 3.9.
Let $(M,g)$ be a Riemannian manifold and let $f$ be a Morse–Bott function on $M$. If $\mathbf{x}\in M$ is a critical point of $f$, then there exists a chart $\sigma:U\to\mathbb{R}^{n}$ about $\mathbf{x}$ such that, with $\mathbf{x}=(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{n-p-q})$
$\sigma(\mathbf{x})$ $=0,$
$f\circ\sigma^{-1}(\mathbf{x})$ $=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2})+\mathcal{O}(x^{2}),$
$g_{ij}(\mathbf{x})$ $=\delta_{ij}+\mathcal{O}(x),$
where $\delta_{ij}$ denotes the Kronecker symbol.
Proof.
Choose coordinates about $\mathbf{x}$ such that the metric is Euclidean at the $x$ (e.g. normal coordinates), then use orthogonal transformation to diagonalize the Hessian of $f$ at $x$. ∎
Now we can combine the three previous results.
Lemma 3.10.
Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function. For any strict saddle point ${\mathbf{x}}\in\mathrm{Crit}(f)$ with value $f({\mathbf{x}})=c$ there exists $N>0$ and a neighborhood $U$ of ${\mathbf{x}}$ such that
$\Pr(f({\mathbf{x}}_{i+N})<c\,|\,{\mathbf{x}}_{i}\in U)>\delta.$
Proof.
This follows immediately from Corollary 3.7, Lemma 3.8 and Lemma 3.9. ∎
Lemma 3.11.
Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function. Further assume that $f$ has compact sublevel sets and that $\mathbf{x}_{0}\in M$ is not a critical point of $f$. Then the probability that $\mathbf{x}_{i}$ converges to a critical submanifold of a strict saddle points as $i\to\infty$ is zero.
Proof.
Consider some critical value $c$ of $f$ and let $C\subset M$ be a connected component of the corresponding critical submanifold made up of strict saddle points, cf. [48, Def. 2.41]. Note that $C$ is compact. We will show that the probability that $\mathbf{x}_{i}$ converges to $C$ is zero. Indeed let ${\mathbf{z}}\in C$ be arbitrary. Then, by Lemma 3.10, there exists an open set $U\subset M$ containing ${\mathbf{z}}$ such that $\Pr(f(\mathbf{x}_{i+N})<c\,|\,\mathbf{x}_{i}\in U)>\delta$ for some $\delta>0$ and $N>0$.
Moreover, by compactness we can find a finite collection of such open sets covering $C$. Hence it suffices to show that the probability that $\mathbf{x}_{i}$ has an accumulation point in $C\cap U$ is zero. Consider a sequence $\mathbf{x}_{i}$ and let $\mathbf{z}\in C\cap U$ be an accumulation point. Consider two balls $B_{R_{1}}(\mathbf{z}),B_{R_{2}}(\mathbf{z})$ centered on $\mathbf{z}$ with $0<R_{2}<R_{1}$ and both balls contained in $U$. Choosing $R_{1}$ and $R_{2}/R_{1}$ small enough we can guarantee that $\mathbf{x}_{0}\notin B_{R_{1}}(\mathbf{z})$ and that any realization needs at least $N$ steps to traverse the spherical shell $B_{R_{1}}\setminus B_{R_{2}}$. Now choosing a sequence of strictly decreasing $R_{k}$ with $R_{1}$ and all $R_{2k}/R_{2k-1}$ small enough as above we see that the sequence must traverse all shells $B_{R_{2k-1}}\setminus B_{R_{2k}}$
This shows that any sequence with accumulation point in $C\cap U$ must, infinitely often, spend $N$ consecutive steps in $U$. By Lemma 3.10 the probability that this happens is zero. Here we used the fact that by Markovianity the behaviour of each sequence of $N$ consecutive steps behaves independently. This concludes the proof. ∎
3.3 Almost Sure Convergence
The results proven so far guarantee that the algorithm converges almost surely to the set of local minima:
Proposition 3.12.
Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with compact sublevel sets. Assume that $\mathbf{x}_{0}$ is not a critical point. Then for stepsize $\eta\leq 1/\ell$ the randomized gradient descent algorithm converges almost surely to the set of local minima.
Proof.
Since by Lemma 3.4 the function value cannot increase, and since by compactness of sublevel sets the function is lower bounded, the function value must (surely) converge to some value $f^{\star}$. By Corollary 3.5, $f^{\star}$ is almost surely a critical value. By Lemma 3.11 the algorithm almost surely does not converge towards a strict saddle point. Hence we converge almost surely to the set of local minima. ∎
In order to prove convergence to a single local minimum, we first need the following lemma.
Lemma 3.13.
Let $\mathcal{E}_{i}$ be probability distributions with values in the interval $[0,1]$ and let $E_{i}\sim\mathcal{E}_{i}$ for $i\in\mathbb{N}$. Let $\varepsilon>0$. We assume that there is $q>0$ such that
$\Pr(E_{i}\geq\varepsilon\,|\,E_{1}=e_{1},\ldots,E_{i-1}=e_{i-1})\geq q$
for all $i\in\mathbb{N}$ and for all $e_{1},\ldots,e_{i-1}\in[0,1]$. Then for any $\alpha\in(0,q\ln\tfrac{1}{1-\varepsilon})$ it holds that almost surely
$\prod_{i=1}^{n}(1-E_{i})\leq e^{-\alpha n}$
for $n$ large enough.
Proof.
First recall the following basic fact. Consider a sequence of i.i.d. biased coin tosses $X_{i}\in\{0,1\}$ with $\Pr(X_{i}=1)=q$. By the strong law of large numbers [30, Thm. 20.1] the average $\overline{X}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}X_{i}$ converges almost surely to the expectation $q$. Hence for any $p\in(0,q)$, there almost surely exists some $n_{0}$ large enough such that $\overline{X}_{n}\geq p$ for all $n>n_{0}$.
Let any value $n$, and index set $I=\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,n\}$ (ordered increasingly) be given. Then
$\Pr(E_{i}\geq\varepsilon,\,\,\forall i\in I)$ $=\Pr(E_{i_{k}}\geq\varepsilon\wedge\ldots\wedge E_{i_{1}}\geq\varepsilon)$
$=\Pr(E_{i_{k}}\geq\varepsilon\,|\,E_{i_{k-1}}\geq\varepsilon\wedge\ldots\wedge E_{i_{1}}\geq\varepsilon)\ldots\Pr(E_{i_{1}}\geq\varepsilon)$
$\geq q^{k}$
$=\Pr(X_{i_{k}}=1,\ldots X_{1}=1)$
and thus we have the following: For all $p\in(0,q)$, there almost surely exists some $n_{0}$ large enough such that for all $n>n_{0}$ it holds that at least $pn$ of the $E_{i}$ are greater than or equal to $\varepsilon$. In this case
$\prod_{i=1}^{n}(1-E_{i})\leq(1-\varepsilon)^{np}=e^{-\alpha n},$
with $\alpha=p\ln\tfrac{1}{1-\varepsilon}$. ∎
Theorem 3.14.
Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with $f$ has compact sublevel sets. Further assume that ${\mathbf{x}}_{0}$ is not a critical point of $f$. Then, for stepsize $\eta\leq 1/\ell$, the randomly projected gradient descent algorithm converges almost surely to a local minimum.
Proof.
The idea is the proof is as follows: first we show that almost surely the distance to the set of local minima decreases exponentially. This then implies that the size of each step decreases exponentially as well, and hence the algorithm converges absolutely to a local minimum.
Proposition 3.12 shows that ${\mathbf{x}}_{i}$ converges almost surely to some connected set of local minima denoted $C$. Let ${\mathbf{z}}\in C$ be such a local minimum, and let $H({\mathbf{z}})$ denote the Hessian at ${\mathbf{z}}$, and let $a_{\min}({\mathbf{z}})$ denote the smallest non-zero eigenvalue of $H({\mathbf{z}})$. By compactness and continuity $a_{\min}({\mathbf{z}})$ has a minimum on $C$, which we denote $a_{\min}$. In a small enough neighborhood of ${\mathbf{z}}$, Lemma 3.9 shows that
$\|\operatorname{grad}f({\mathbf{x}})\|^{2}$ $=\sum_{j}4a_{j}(\mathbf{z})^{2}(x_{j}-z_{j})^{2}+\mathcal{O}(\|{\mathbf{x}}-\mathbf{z}\|^{3})$
$\geq 4a_{\min}(\mathbf{z})f({\mathbf{x}})+\mathcal{O}(\|{\mathbf{x}}-\mathbf{z}\|^{3})$
$\geq 2a_{\min}(\mathbf{z})f({\mathbf{x}}).$
Together with Lemma 3.4 we find
$f({\mathbf{x}}_{i+1})\leq\Big{(}1-((\mathbf{u}_{i})_{N})^{2}\delta 2a_{\min}(\mathbf{z})\Big{)}f({\mathbf{x}}_{i})$
where $\delta=\eta(1-\tfrac{\ell\eta}{2})$.
By Lemma 3.13 this shows that there is some $\alpha>0$ such that $f({\mathbf{x}}_{i})\leq f({\mathbf{x}}_{0})e^{-\alpha i}$ almost surely for $i$ large enough. Again by Lemma 3.4 we have that
$\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|\leq\sqrt{\frac{|f({\mathbf{x}}_{i+1})-f({\mathbf{x}}_{i})|}{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}.$
This shows that
$\sum_{i=0}^{\infty}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|\leq\frac{1}{\sqrt{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\sum_{i=0}^{\infty}\sqrt{f({\mathbf{x}}_{i})}=\sqrt{\frac{f({\mathbf{x}}_{0})}{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\frac{1}{1-e^{\alpha/2}}<\infty,$
and so the total length of the trajectory is finite. Thus ${\mathbf{x}}_{i}$ almost surely converges absolutely to a local minimum of $f$. ∎
4 Quantitative Results in Low Dimensions
The goal of this section is to study the hitting time $\tau$ of the sublevel set of the critical value. By construction, the stochastic process $x^{(i)}$ is a Markov chain, since the step $x^{(i+1)}-x^{(i)}$ depends only on the value of $x^{(i)}$. As such, for small enough step size $\eta$ we can approximate the algorithm by an appropriate Ito stochastic differential equation (SDE).
4.1 Two-Dimensional Euclidean Case
In the simplest case we consider the Morse function $f(x,y)=ax^{2}-by^{2}$ on $\mathbb{R}^{2}$ in the Euclidean metric. Due to the scale invariance we can normalize the state after each iteration and so we obtain a discrete-time stochastic process on the unit circle. For small enough stepsize $0<\eta\ll 1$ the iteration rule is approximately [^5] The formula is obtained by locally approximating the unit circle by its tangent line. given by
$\Delta\phi^{i}=\phi^{i+1}-\phi^{i}=\eta\,\frac{a+b}{2}\sin(2\phi)+u_{2}\eta\sqrt{(a\cos(\phi))^{2}+(b\sin(\phi))^{2}}$ (9)
where $\mathbf{u}=(u_{1},u_{2})$ is chosen uniformly at random from the unit circle and hence
$\mathbb{E}[\Delta\phi]=\eta\,\frac{a+b}{2}\sin(2\phi),\quad\operatorname{Var}(\Delta\phi)=\frac{\eta\sqrt{(a\cos(\phi))^{2}+(b\sin(\phi))^{2}}}{2}$
Using [32, Sec. 6.2] we can approximate this process using the stochastic differential equation (SDE)
$d\Phi_{t}=\frac{a+b}{2}\sin(2\phi)dt+\frac{\sqrt{\eta((a\cos(\phi))^{2}+(b\sin(\phi))^{2})}}{2}dW_{t}$ (10)
For simplicity we consider the simplest case $a=b=1$. Then the limiting SDE reads
$d\Phi_{t}=\sin(2\Phi_{t})dt+\frac{\sqrt{\eta}}{2}dW_{t}$
The approximation is visualized in Figure 3.
Figure 3: We use the constants $\eta=0.01$, $T=10$, $a=b=1$. Left: $50$ realizations of the stochastic process (9) for $T/\eta=1000$ steps. Right: $50$ realizations of the stochastic process (10) for time $T$ and with timestep size $\Delta t=0.001$ (simulated using the Euler-Maruyama scheme [32, Sec. 9.1]).
In order to understand how long it takes the algorithm to pass the saddle point we are interested computing in the hitting time
$\tau=\inf\{t>0:\Phi_{t}=\pm\frac{\pi}{4}\}.$
Close to the origin, the SDE can be linearized as $d\Phi_{t}=2\Phi_{t}dt+\frac{\sqrt{\eta}}{2}dW_{t}$, which is a mean repelling Ornstein–Uhlenbeck process. Away from the origin, the deterministic part dominates and we can solve the (deterministic) ODE $\dot{\Phi}_{t}=\sin(2\Phi_{t})$. First we approximate the hitting time distribution of the repelling Ornstein-Uhlenbeck process.
Lemma 4.1.
Let $X_{t}$ be the solution of the SDE $dX_{t}=\kappa X_{t}dt+\sigma dW_{t}$ with $\kappa,\sigma>0$ and $X_{0}=0$. Setting $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, it holds that $X_{t}\sim\mathcal{N}(0,\tilde{\sigma}(t)^{2})$. If we denote by $\tau_{c}$ the hitting time of ±$\pm c$ (where $c>0$), we find the lower bound $\Pr[\tau_{c}\leq t]\geq\Pr[|X_{t}|\geq c]=1+\operatorname{erf}\big{(}\frac{-c}{\tilde{\sigma}(t)\sqrt{2}}\big{)}$ where $\operatorname{erf}$ denotes the error function.
Proof.
The SDE is linear and hence has the well-known solution
[…]
where $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, see [32, Sec. 4.2 & 4.4]. It is clear that if $|X_{t}|\geq c$ then $\tau_{c}\leq t$, which implies the last statement. ∎
The approximation is very accurate with error smaller than $\Delta t$ if $\kappa\Delta t\geq 1$ and $c\gg\sigma/\sqrt{\kappa}$, as illustrated in Figure 4.
As in the proof of Lemma 4.1 the solution of the SDE with initial condition $X_{0}=c$ is given by
[…]
then
$e^{\kappa t}c-n\tilde{\sigma}(t)\geq c\implies c\geq n\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}\frac{1}{e^{\kappa t}-1}\geq\frac{n\sigma}{\sqrt{2\kappa}}\max\Big{(}1,\sqrt{\tfrac{2}{\kappa t}}\Big{)}.$
Hence for the error in $\tau_{c}$ to be small compared to $\Delta t$, we need $c\gg\frac{\sigma}{\sqrt{\kappa}}$ if $\kappa\Delta t$ is big, and $c\gg\frac{\sigma}{\kappa\sqrt{\Delta t}}$ if $\kappa\Delta t$ is small.
Figure 4: We used the constants $\kappa=2$, $\sigma=3$ and $c=10$. Shown are the histograms of the c.d.f. and p.d.f of $\tau_{c}$ computed using timesteps of size $0.001$ (orange) and the lower bound for the c.d.f. and the resulting p.d.f. obtained in Lemma 4.1 (blue).
Now if we choose $c$ small enough that the linearization of $\sin$ is accurate but large enough that the approximation of Lemma 4.1 is good (which is always possible if $\eta$ is small enough), then we can approximate $\tau$ as follows.
We fix some value $c$ and note that the ODE $\dot{\phi}_{t}=\sin(2\phi_{t})$ with initial condition $\phi_{0}=c$ has the solution $\phi_{t}=\arctan(e^{2t}\tan(c))$, and hence the hitting time of $\pi/4$ is $\tilde{\tau}_{c}=-\tfrac{1}{2}\ln(\tan(c))$. Hence we have approximately $\tau\simeq\tau_{c}+\tilde{\tau}_{c}$. This is illustrated in Figure 5 for realistic values.
Figure 5: Using $\eta=0.01$ we plot the c.d.f. of the hitting time $\tau$ of $\tfrac{\pi}{4}$. In orange for the difference equation, in blue for the stochastic differential equation ($\Delta t=0.001$), and in red for the analytic approximation.
5 Ground State Optimization
In quantum information science an important task relates to finding the eigenstate (‘ground state’) and the smallest eigenvalue (‘ground-state energy’) of a Hermitian operator $A$ that describes the energy of the system (i.e. the Hamiltonian). For this setting, we leave the general notation $f:M\to\mathbb{R}$ above and specialise to the cost function over the unitary group (or orbit) $J:U\to\mathbb{R}$ we aim to minimize. For finding the smallest expectation value of $A$ w.r.t. the state $\rho$, the cost function is given by
$J(U):=\operatorname{tr}(AU\rho U^{*})=\langle A,\operatorname{Ad}_{U}(\rho)\rangle,$ (11)
where $\rho\in\mathfrak{pos}_{1}(n)$ is the initial state of the system and $U$ is a unitary transformation that describes a quantum circuit (writing $U^{*}$ for the complex conjugate transpose).
By choosing an appropriate eigenbasis of $A$ we may assume that $A$ is diagonal with eigenvalues in non-increasing order. The function we want to optimize is of the form of Eqn. (11). Many properties of this function can be found in [22] in the more general setting of semisimple Lie groups. We recall the relevant properties here, adapted to our setting:
Proposition 5.1.
Assume that $\rho$ and $A$ are diagonal  [^6] This amounts to a shift in the function $f$ and an appropriate choice of basis.. Then the following holds.
  1. (i)
    $U$ is a critical point of $f$ if and only if $[A,\operatorname{Ad}_{U}(\rho)]=0$. Hence the critical set of $f$ is equal to
    $\mathsf{SU}(n)_{A}\,\mathcal{S}_{n}\,\mathsf{SU}(n)_{\rho},$
    where $\mathcal{S}_{n}$ is the set of $n$ by $n$ permutation matrices. In particular it is a disjoint union of finitely many compact connected submanifolds.
  2. (ii)
    The Hessian at a critical point $U$ is given by
    $\frac{d^{2}}{dt^{2}}\langle A,\operatorname{Ad}_{e^{itH}}(\rho)\rangle=\sum_{i>j}2(a_{i}-a_{j})(\rho_{i}-\rho_{j})|\langle i|\operatorname{Ad}_{U}(H)|j\rangle|^{2},$
    and in particular $f$ is Morse–Bott. [^7] Note that the function $f$ is never Morse, since the maximal torus in $\mathsf{SU}(n)$ stabilizes $A$ and $\rho$.
  3. (iii)
    There is only one local minimal (resp. maximal) value.
Proof.
i: See Lemma 1.1 and Propositions 1.2 and 1.3 of [22]. ii: See Proposition 1.4 and Corollary 1.5 of [22]. iii: See Remark 1.6 of [22]. ∎
Definition 2.
[26][Thm. 3] A unitary representation $\pi:G\to U(n)$, $g\mapsto\pi(g)=:U_{g}$ a finite group $G$ is called a unitary $t$-design with $t\in\mathbb{N}_{0}$ if one of the following equivalent conditions is satisfied:
  1. (a)
    For all polynomials $p\in\mathbb{C}_{t}[Z,\bar{Z}]$ one has the equality
[…]
    where $\mathbb{C}_{t}[Z,\bar{Z}]$ denotes the set of all polynomials in $Z:=(z_{ij})_{i,j=1,\dots n}$ and $\bar{Z}:=(\bar{z}_{ij})_{i,j=1,\dots n}$, which are $t$ homogeneous in $Z$ and $\bar{Z}$., i.e. [^8] This is equivalent to $p(\lambda Z,\bar{\lambda}\bar{Z})=|\lambda|^{2t}p(Z,\bar{Z})$. with $p(\lambda Z,\mu\bar{Z})=\lambda^{t}\mu^{t}p(Z,\bar{Z})$.
  2. (b)
    For all $H\in\mathbb{C}^{n^{t}\times n^{t}}$ one has the equality
[…]
  3. (c)
    One has the equality
[…]
For $t=2$ one has the further equivalence
  1. (d)
    The tensor square representation $\pi\otimes\pi:G\to U(n)\otimes U(n)$, $g\mapsto U_{g}\otimes U_{g}$ acting on $\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ has exactly two irreducible components, namely $\mathrm{Sym}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ and $\mathrm{Alt}(\mathbb{C}^{n}\otimes\mathbb{C}^{n}):=\mathbb{C}^{n}\wedge\mathbb{C}^{n}$.
Remark 5.2.
A more general notion of $t$-designs [20, 21] (see also [26]) allows any finite subset of $U(n)$ which satisfies (a), (b) or, equivalently,(c). However, since almost all known examples are constructed via group representations, we focus here on this restricted approach which is sometimes called group design. Moreover, the reader should note the following to facts: (i) One can assume without loss of generality that $\pi$ is faithful (i.e., one-to-one) because it is straightforward to show that $[g]\mapsto\pi(g)$, $[g]\in G/\mathrm{ker}\pi$ yields a $t$-design whenever $\pi$ is a $t$-design. (ii) Every $t$-design is also $t^{\prime}$-design for $t^{\prime}\leq t$. This follows readily from condition (b) by choosing $H$ of the form $H^{\prime}\otimes I_{n}\otimes\cdots\otimes I_{n}$.
Lemma 5.3.
Let $\pi:G\to U(n)$ be any unitary representation. Then $\pi$ yields a $2$-design in the sense of Def. 2 if and only if $\pi$ acts irreducibly on $i\mathfrak{su}(n)$ via conjugation, i.e. the representation $\hat{\pi}:G\to U\big{(}i\mathfrak{su}(n)\big{)}$, $\pi(g)\,H:=U_{g}HU_{g}^{*}$ is irreducible.
Proof.
First, $\hat{\pi}$ obviously “extends” to $\tilde{\pi}:G\to U\big{(}\mathbb{C}^{n\times n}\big{)}$, $\tilde{\pi}(g)\,A:=U_{g}AU_{g}^{-1}$. Then a straightforward computation shows that the representations $\tilde{\pi}$ and $g\mapsto U_{g}\otimes\bar{U}_{g}$ are equivalent and thus for simplicity we will use the same symbol for both representations. Next, we investigate the commutant of the representations $U_{g}\otimes\bar{U}_{g}$ and $U_{g}\otimes U_{g}$ in $\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}\cong\mathbb{C}^{n^{2}\times n^{2}}$, i.e. we are interested in the solutions $Z$ of
$(U_{g}\otimes\bar{U}_{g})Z=Z(U_{g}\otimes\bar{U}_{g})\quad\text{for all}\quad g\in G$ (12)
and
$(U_{g}\otimes U_{g})Z=Z(U_{g}\otimes U_{g})\quad\text{for all}\quad g\in G\,,$ (13)
respectively. Eq. (12) and (13) are equivalent to $Ad_{U_{g}\otimes\bar{U}_{g}}(Z)=Z$ and $Ad_{U_{g}\otimes U_{g}}(Z)=Z$, respectively. Finally, a tedious but straightforward computation shows that the partial transposed operator $\Phi:\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}$, $A\otimes B\mapsto A\otimes B^{T}$ yield an intertwining map for $Ad_{U_{g}\otimes\bar{U}_{g}}$ and $Ad_{U_{g}\otimes U_{g}}$, i.e.
$Ad_{U_{g}\otimes\bar{U}_{g}}\circ\Phi=\Phi\circ Ad_{U_{g}\otimes U_{g}}\,.$
This implies $\Phi$ maps the commutator of $\tilde{\pi}$ to the commutator of $\pi\otimes\pi$ and therefore the dimensions of commutators (and consequently the number of irreducible subspaces) coincide. Thus $\mathrm{Sym}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ and $\mathrm{Alt}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ are the only irreducible subspaces of $\pi\otimes\pi$ if and only if $\mathbb{C}\,I_{n}$ and $\mathrm{sl}_{0}(n)$ are the only irreducible subspaces to $\tilde{\pi}$ (or, equivalently, $i\mathfrak{su}(n)$ is the only irreducible subspace of $\hat{\pi}$). ∎
Remark 5.4.
Note that the above result does not mean that the tensor square representation $\pi\otimes\pi$ and $\tilde{\pi}$ are equivalent and, in fact, they are not—as easy examples demonstrate. A similar result for Lie-algebra representations was elaborated on in [65] as follow-up on [64] and in particular on [17].
Lemma 5.5.
Let $\pi:G\to\mathrm{U}(n)$ be a unitary $t$-design with $t\geq 2$ and let $H_{s}\in i\mathfrak{su}(n)$ be a traceless Hermitian matrix. Then any subset of $\{U_{g}H_{s}U_{g}^{*}\}_{g\in G}$, where one element is removed, still spans $i\mathfrak{su}(n)$.
Proof.
By Lemma 5.3 we know that the $\hat{\pi}$-invariant subspace spanned by $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ has to coincide with $i\mathfrak{su}(n)$ and thus $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ contains a basis of $i\mathfrak{su}(n)$. Moreover, by the defining property (b) it is straightforward to see that every $1$- and therefore also every $2$-design has to act irreducibly on $\mathbb{C}^{n}$. Thus we conclude
$\sum_{g\in G}U_{g}H_{s}U_{g}^{*}=0$
because the left hand side of the above equation belongs to the commutant of $\pi$ and has trace zero. Now if there is some $g_{0}\in G$ such that $\{U_{g}H_{s}U_{g}^{*}:g\in G\,,g\neq g_{0}\}$ does not span $i\mathfrak{su}(n)$, then $U_{g_{0}}H_{s}U_{g_{0}}^{*}$ does not lie in the span of the others, and hence $\sum_{g\in G}U_{g}H_{s}U_{g}^{*}\neq 0$, which contradicts the above. ∎
6 Conclusion, Discussion, and Outlook
We have analyzed the convergence properties of a recently introduced Haar-randomly projected gradient descent algorithm for cost functions taking the form of a smooth Morse–Bott function (with compact sublevel sets) on a Riemannian manifold. For making it efficient in quantum optimizations, one can approximate the Haar-random projections via unitary 2-designs. For both scenarios we have proven that (i) the respective algorithm almost surely escapes saddle points (Lem. 3.11) and (ii) it almost surely converges to a local minimum (Thm. 3.14). Moreover we have studied the time required by the algorithm to pass a saddle point in a simple two-dimensional setting (Sec. 4.1). Note that unlike adiabatic ground-state preparation strategies (that rest on first preparing some desired initial state to find the ground state of some target Hamiltonian), our randomized Riemannian gradient flow algorithm does not require knowledge of the initial state, as the algorithm converges for almost all initial points. Moreover, as in the quantum setting for ground state problems the critical points just comprise saddles and global extrema, here our result implies almost sure convergence to the global minimum.
However, our approach inherits a key problem already arising without projections: the overall speed of convergence to globally optimal solutions may be slow [43], since it depends on the scaling of the magnitude of the (projected) Riemannian gradient. For Riemannian optimizations over the unitary group, the gradient magnitude converges to its expectation value (being zero), while the variance is inversely proportional to the dimension $d=2^{n}$ [44]. Thus the probability for a gradient magnitude larger than the noise level of a quantum computer decays exponentially with the number of qubits $n$. Well known as barren plateaux problem in quantum optimization [44], such exponentially flat regions constitute a main challenge to all variational quantum algorithms in high dimensions on quantum computers.
The algorithmic steps analyzed here are entirely modular. In view of future applications the (approximate) random projections w.r.t. the Haar measure may well be replaced by selections from other problem-adapted measures without sacrificing the convergence properties. In particular, one may wish to select the measures such that the projections of the gradient do not subside in numerical noise prematurely and thus circumvent the notorious barren-plateaux problem. Moreover, one may consider random projections into subspaces by techniques of compressed sensing and shadow tomography to better approximate the full gradient flow. Likewise, approximations with tensor-network methods (such as, e.g., [60, 61]) may be envisaged as long as one remains on a Riemannian manifold. Needless to say the techniques presented here lend themselves to be taken over to higher-order quasi Newton methods (like L-BFGS) also on quantum computers.
Thus we anticipate that the convergence guaranteed for (randomized) Riemannian gradient flows w.r.t. Morse–Bott type smooth cost functions will encourage wide application in and beyond quantum optimization.
Acknowledgements
E.M. and T.S.H. are supported by the Munich Center for Quantum Science and Technology (MCQST) and the Munich Quantum Valley (MQV) with funds from the Agenda Bayern Plus. C.A. acknowledges support from the National Science Foundation (Grant No. 2231328) and Knowledge Enterprise at Arizona State University.
Appendix
Appendix A Some Technical Results
The probability distributions of $u_{N}$ and $u_{N}^{2}$ are important for understanding the behaviour of the algorithm. Fortunately they can be described quite easily using the beta distribution. Recall that the p.d.f of the beta distribution with parameters $a,b$ is given by $f(x)=\mathrm{const}\cdot x^{a-1}(1-x)^{b-1}.$
Lemma A.1.
If $u\sim\mathcal{U}(S^{N-1})$ and $u_{N}$ is the last coordinate, then
$u_{N}\sim\mathcal{B}_{N}:=2\,\mathrm{Beta}\left(\frac{N-1}{2},\frac{N-1}{2}\right)-1,\quad u_{N}^{2}\sim\mathcal{B}_{N}^{2}:=\mathrm{Beta}\left(\frac{1}{2},\frac{N-1}{2}\right)$
where $\mathrm{Beta}(a,b)$ denotes the beta distribution, and hence
$\mathbb{E}[u_{N}^{2}]=\frac{1}{N},\quad\operatorname{Var}(u_{N}^{2})=\frac{2(N-1)}{N^{2}(N+2)}.$
Proof.
[…]
Since $u_{N}$ is symmetrically distributed around 00, it is easy to compute its p.d.f. starting from that of $u_{N}^{2}$, and one obtains the form given above. The expectation and variance of $u_{N}^{2}$ follow easily from the formula for the moments of a beta distributed variable $X\sim\mathrm{Beta}(a,b)$:
$\mathbb{E}[X^{k}]=\prod_{r=0}^{k-1}\frac{a+r}{a+b+r}.$
This concludes the proof. ∎
Lemma A.2.
For $N\geq 5$ it holds that
$d_{K}(\sqrt{N}\mathcal{B}_{N},\mathcal{N}(0,1))\leq\frac{1}{N}$
where $d_{K}$ denotes the Kolmogorov distance (the supremum distance between the cumulative distribution functions). This shows that $u_{N}^{2}$ is almost distributed according to the distribution $1/N\chi^{2}_{1}$, more precisely
$d_{K}(N(\mathcal{B}_{N})^{2},\chi^{2}_{1})\leq\frac{2}{N}.$
If $\Phi$ denotes the cumulative distribution function of a standard normal distribution, then for any $k\geq 0$ it holds that
$\Pr\Big{(}u^{2}_{N}\geq\frac{1}{k^{2}N}\Big{)}\geq 2\Big{(}1-\Phi(\tfrac{1}{k})-\frac{1}{N}\Big{)}.$
Proof.
The first inequality follows from [51, p. 20] and [50, Thm. 1.2]. It is easy to see that taking the square of both distributions at most doubles the Kolmogorov distance, and this yields the second inequality. Finally we compute
$\Pr\Big{(}u^{2}_{N}\geq\frac{1}{k^{2}N}\Big{)}=2\big{(}1-\Pr\big{(}u_{N}\leq\tfrac{1}{k\sqrt{N}}\big{)}\big{)}\geq 2\big{(}1-\Phi\big{(}\tfrac{1}{k}\big{)}-\frac{1}{N}\big{)},$
proving the third inequality. ∎
Lemma A.3.
Let $M$ be a compact manifold and let $\{v_{j}\}_{j=1}^{k}$ be a set of continuous vector fields on $M$ which span the tangent space at every point. Now let any vector field $v$ on $M$ be given. Then it holds that there is some $\varepsilon>0$ such that
$\min_{x\in M}\max_{i}\langle v_{i},v\rangle^{2}=\varepsilon.$
Proof.
The functions $\langle v_{j},v\rangle^{2}$ defined on $M$ are smooth and non-negative. Hence their maximum over $j$ is still a continuous non-negative function on $M$. Since the $v_{j}$ span each tangent space, this function is even strictly positive, and by compactness of $M$, there is a positive global minimum. ∎

### v2

@@
 Abstract
-We analyze the convergence properties of gradient descent algorithms on Riemannian manifolds.
-We study randomization of the tangent space directions of Riemannian gradient flows for minimizing smooth cost functions (of Morse–Bott type) to obtain convergence to local optima.
-We prove that through randomly projecting Riemannian gradients according to the Haar measure, convergence to local optima can be obtained almost surely despite the existence of saddle points.
-As an application we consider ground state preparation through quantum optimization over the unitary group.
-In this setting one can efficiently approximate the Haar-random projections by implementing unitary 2-designs on quantum computers.
-We prove that the respective algorithm almost surely converges to the global minimum that corresponds to the ground state of a desired Hamiltonian.
-Finally, we discuss the time required by the algorithm to pass a saddle point in a simple two-dimensional setting.
+We analyze convergence of gradient-descent methods on Riemannian manifolds.
+In particular, we study randomization of Riemannian gradient algorithms for minimizing smooth cost functions (of Morse–Bott type).
+We prove that randomized gradient descent methods, where the Riemannian gradient is replaced by a random projection of it, converge to a single local optimum almost surely despite the existence of saddle points.
+We consider both uniformly distributed and discrete random projections.
+We also discuss the time required to pass a saddle point.
+As a major application, we consider ground-state preparation through quantum optimization over the unitary group.
+In mathematical terms our randomized algorithm applied to the trace function $U\to\operatorname{tr}(AU\rho U^{*})$ almost surely converges to its global minimum.
+The minimum corresponds to the smallest eigenvalue (ground state) of the selfadjoint operator $A$ (Hamiltonian) if $\rho$ is a rank-one projector (pure state).
+In this setting, one can efficiently replace the uniform random projections by implementing so-called discrete unitary 2-designs.
 keywords:
@@
 1 Introduction
-Both gradient systems and numerical linear algebra come with a long history as well as a vibrant research activity.
+Both gradient systems and numerical linear algebra come with a long history, as well as a vibrant research activity.
 For instance, gradient systems have been advanced lately from Hilbert manifolds [56] to general metric spaces (e.g., in the context of optimal mass transport [2, 53]).
-In contrast, numerical linear algebra has faced factorization problems of huge tensor structures [6] (tensor SVD) using “randomization strategies” for their classical algorithms in order to handle these amounts of data.
-Only a few decades ago a systematic study of their interplay started by analyzing the QR-algorithm and interior point methods from a (Riemannian) geometric point of view, cf.
+In contrast, numerical linear algebra has faced factorization problems of large tensor structures [8] (tensor SVD) using “randomization strategies” for their classical algorithms in order to handle these huge amounts of data.
+A few decades ago, a systematic study of their interplay started by analyzing the QR-algorithm and interior point methods from a (Riemannian) geometric point of view, cf.
 the collection of [13].
-In this development the seminal work of [12, 13] on the so-called “double-bracket flow” on orbits of the orthogonal group (and likewise of the unitary group in [14]), followed by independent similar ideas of [16], has sparked a plethora of applications.
-For an early overview on optimization techniques on Riemannian manifolds (including higher-order methods) see [56] in [11].
-A mathematical account in all detail can be found in the monograph by [28] including discretization schemes, suitable step sizes, and proofs of convergence.
-For the latter, the authors heavily exploited the Morse-Bott structure of the cost function (i.e.
-the particular stucture of the Hessian at possibly non-isolated critical points, cf.
+In this development, the seminal work of [15, 16] on the so-called “double-bracket flow” on orbits of the orthogonal group (and likewise of the unitary group in [17]), followed by independent similar ideas of [19], has sparked a plethora of applications.
+For an early overview on optimization techniques on Riemannian manifolds (including higher-order methods) see [65] and [13].
+A detailed mathematical account can be found in the monograph by [33], including discretization schemes, suitable step sizes, and proofs of convergence.
+For the latter, the authors heavily exploited the Morse–Bott structure of the respective cost functions (i.e.
+the particular structure of the Hessian at possibly non-isolated critical points, cf.
 Subsec.
-2.3 and [22] to obtain convergence to a single critical point.
-Based on these developments, higher-order methods with applications to further cost functions on classical matrix manifolds are treated in [1].
-In [54] these ideas have been extended to flows and their discretization schemes on more general homogeneous spaces.
-There they were also connected to applications in quantum dynamics as well as to applications in numerical (multi-)linear algebra like higher-rank tensor approximations—as elaborated by [19] to complement standard methods (like higher-order powers of [34] or quasi Newton methods of [53]).
+2.3 here and [26]) to obtain convergence to a single critical point.
+Based on these developments, higher-order methods with applications to a variety of cost functions on classical matrix manifolds are treated in [1] or, more recently, by [14].
+In [63], the above ideas have been extended to flows and their discretization schemes on more general homogeneous spaces.
+There, the authors also established connections to applications in quantum dynamics as well as to applications in numerical (multi-)linear algebra, such as higher-rank tensor approximations — as elaborated by [23] to complement standard methods (such as higher-order powers of [40] or quasi-Newton methods of [62]).
 Independently, in the physics community, [72] (re)devised gradient flows of Brockett type, e.g., to (band-)diagonalize Hamiltonians, which the monograph by [37] elaborated to address further quantum many-body applications.
-Often gradient-flow approaches for deriving optimization schemes on Riemannian manifolds hinge on the computability of the Riemannian exponential ensuring to take the gradient from the tangent space back on the manifold.
-To be precise, this is crucial whenever the manifold cannot be identified with its tangent spaces such as for the unitary group or unitary orbits.
-With these stipulations, Riemannian optimization techniques constitute a versatile toolbox, since a cost function can readily be tailored to the particular application such as finding extremal eigen- or singular values (i.e.
-problems otherwise addressed by variational approaches).
+Often, gradient-flow approaches for deriving optimization schemes on Riemannian manifolds hinge on the computability of the Riemannian exponential, ensuring to take the gradient from the tangent space back to the manifold.
+To be precise, this is crucial whenever the manifold cannot be identified with its tangent spaces, such as for the unitary group and its orbits.
+But, in particular in this case, analytical expressions and efficient numerical approximations for the Riemannian exponential are well known.
+With these stipulations, Riemannian optimization constitutes a versatile toolbox, since cost functions can readily be tailored to the particular application such as finding extremal eigen- or singular values.
 These instances also connect to other branches of mathematics such as numerical and $C$-numerical ranges and their extremal values, the numerical and $C$-numerical radii (see [31, 45]).
 Over the last decade, techniques from Riemannian optimization have also been adopted in quantum information science.
-For example, in quantum computing, hybrid quantum-classical algorithms [8], such as variational quantum algorithms (VQAs) [15, 59], have been developed to solve ground state problems [49] (i.e.
-typically minimal eigenvalue problems), combinatorial optimization [23], and quantum machine learning problems [9].
-In VQAs a fixed parameterized quantum circuit, i.e.
-a parameterized set of unitary transformations (usually implemented by a fixed set of unitary gates), is iteratively optimized in tandem with a classical computer to minimize a cost function whose global optima encode the solution(s) to the desired problem.
-More precisely, a quantum processor, including a measurement device, is used to output the current (e.g., expectation) value of the cost function in each iterative step.
-Interleaved, the classical computer component provides the optimization routine that determines the parameter update for the next iteration on the quantum processor.
-Since quantum computers can, in principle, handle some classically exponentially complex problems, the hope is that VQAs will be able to solve problems of large size a solely classical algorithm would be incapable of.
-However, since the optimization problems VQAs aim to solve are typically non-convex, which can often be traced back to the quantum circuit parameterization [41, 36], the iterative search for the optimal parameters can get stuck in suboptimal solutions [10].
-To overcome this challenge, adaptive quantum algorithms have been designed.
-Instead of fixing a quantum circuit and picking a parameterization as in VQAs, adaptive algorithms successively grow a quantum circuit informed by measurement data from the quantum processor.
-In this approach the cost function is minimized while the quantum circuit is grown, see [25, 63, 42, 43].
-One of the most prominent examples of such a strategy is the so-called ADAPT-VQE algorithm [25], which was originally proposed to solve ground state problems in chemistry, and was further developed later to solve combinatorial optimization problems on quantum computers [58].
-Unfortunately, since typically in adaptive quantum algorithms the circuit growth is informed by gradient estimates performed by the quantum computer, adaptive quantum algorithms face similar challenges as VQAs, i.e., the adaptive circuit growth can get stuck when gradients vanish.
-This issue has recently been addressed by identifying adaptive quantum algorithms as quantum computer implementations of Riemannian gradient flows on the special unitary group $\text{SU}(d)$ of dimension $d=2^{n}$ where $n$ is the number of qubits of the quantum computer [63].
-Since implementing Riemannian gradient methods on quantum computers requires in general exponential resources (i.e., the number of gates and/or the number of iterations of the corresponding quantum circuit growths exponentially in $n$).
-[63] proposed projecting the Riemannian gradient into smaller dimensional subspaces that scale polynomially in $n$, which in turn yields scalable quantum computer implementations.
-However, the proposed “dimension reduction” strategy of the Riemannian gradient to obtain efficient quantum computer implementations comes with a cost: While it can be shown that in the quantum computing setting (i.e., the Riemannian gradient flow on the special unitary group for grounds state problems) the full gradient flow converges almost surely to the problem solution given by the global minimizer, despite the existence of strict saddle points (i.e., saddle points at which the Hessian has at least one negative eigenvalue) [37], for the projected versions no convergence guarantees can be made.
-Indeed, numerical simulations suggest that existing versions of gradient algorithms that are efficiently implementable on quantum computers also suffer from converging to suboptimal solutions.
-Inspired by the success of classical randomized gradient descent [52], this problem has most recently been addressed by some of the authors in [43] via “random projections” of the Riemannian gradient, as depicted in Fig.
-1, which are efficiently implementable on quantum computers.
-An efficient quantum computer implementation of these random directions can be achieved through exact and approximate unitary 2-designs [21].
-The numerical experiments presented in [43] indicate that such a randomization procedure of the Riemannian gradient indeed yields (almost surely) convergence to the problem solution.
-In this paper we consider the randomized gradient algorithm presented in [43] in the more general setting of Morse–Bott cost functions on Riemannian manifolds.
-For both continuous and discrete probability distributions of the random gradient directions we prove that the algorithm almost surely converges to a local minimum even.
-Finally we show how the result can be applied to quantum optimization tasks, such as ground state preparation.
-In the quantum setting there are no local optima, thereby we prove that the randomized gradient algorithm converges almost surely to the global minimum.
-Figure 1: Schematic representation of the randomized gradient descent algorithm on a Riemannian manifold $\mathcal{M}$.
-At a point $X_{k}$ on $\mathcal{M}$ the Riemannian gradient $\operatorname{grad}J(X_{k})$ of a cost function $J:\mathcal{M}\to\mathbb{R}$ lies in the tangent space $T_{X_{k}}\mathcal{M}$ depicted as plane spanned by $\Omega_{k}$ and its orthocomplement $\Omega_{k}^{\perp}$.
-Instead of iteratively following the full gradient in order to optimize $J$, here in each step we consider a projection of $\text{grad}J(X_{k})$ onto a randomly chosen tangent-space direction $\Omega_{k}$.
+For example, in quantum computing, hybrid quantum-classical algorithms [10], such as variational quantum algorithms (VQAs) [18, 69], have been developed to solve ground-state problems [58], combinatorial optimization problems [27], and quantum machine learning problems [11].
+[73] proposed projecting the Riemannian gradient into smaller dimensional subspaces to obtain scalable [^1] Sec.
+5 gives a precise meaning of ‘scalable quantum computers’ quantum-computer implementations.
+However, the “dimension reduction” strategy comes at a cost: While the full gradient flow converges almost surely to some local minimum in spite of strict saddle points222i.e., saddle points at which the Hessian has at least one negative eigenvalue [44], for the projected versions no such guarantees can be made.
+Indeed, numerical simulations see convergence to artificial local minima induced by the projection into a fixed subspace.
+Inspired by the success of classical randomized gradient descent [61], the problem of artificial minima has recently been circumvented by some of the authors in [50] via “random projections” of the Riemannian gradient.
+(Details on different optimization algorithms and their efficient implementations on quantum computers are postponed to Section 5.)
+In this paper we consider the randomized gradient algorithm presented in [50] in the more general setting of the large class [^3] Indeed, on a compact manifold, the set of Morse functions, which is a subset of the Morse–Bott functions, forms an open dense subset of the space of smooth functions [3].
+In this sense, Morse–Bott functions are generic.
+of Morse–Bott cost functions on Riemannian manifolds.
+For both continuous and discrete probability distributions of the random gradient directions, we prove that in fact the algorithm almost surely converges to a local minimum.
+This is a significant improvement on the state-of-the-art, where thus far (see [32]) only convergence to the set of critical points has been proven without the Morse–Bott assumption.
+Related papers have established convergence to local optima for perturbed and stochastic gradient descent [36], also generalized to Riemannian manifolds [21] and [66].
+Finally, we show how our result can be applied to quantum optimization tasks, such as ground state preparation: It is by construction that in the specific quantum setting, there are no local optima, whereby we prove that our randomized gradient algorithm then converges almost surely to a single global minimum.
+Figure 1: Schematic representation of a randomized gradient descent algorithm on a Riemannian manifold $M$.
+At a point $\mathbf{x}_{i}$ on $M$, the Riemannian gradient $\operatorname{grad}f(\mathbf{x}_{i})$ of a cost function $f:M\to\mathbb{R}$ lies in the tangent space ${\sf T}_{\mathbf{x}_{i}}M$ depicted as the plane spanned by $\mathbf{u}_{i}$ and its orthocomplement.
+Instead of iteratively following the full gradient in order to optimize $f$, here in each step we consider a projection of $\text{grad}f(\mathbf{x}_{i})$ onto a randomly chosen tangent-space direction $\mathbf{u}_{i}$.
 This is key for efficient implementations of gradient flows in large-scale quantum optimization problems.
-In the case of the special unitary group the randomized directions are given by (traceless) skew-Hermitian operators $\Omega_{k}=iH_{k}$.
+In the case of the special unitary group the randomized directions are given by (traceless) skew-Hermitian operators $\mathbf{u}_{i}=iH_{i}$.
 The corresponding quantum circuit consists of a sequence of unitary transformations as shown in Fig.
-2.
-Figure 2: Schematic implementation of the discretized Riemannian gradient scheme by a quantum algorithm in case the manifold is $\mathcal{M}=SU(2^{n})$ or $\mathcal{M}=Ad_{SU(2^{n})}$.
-For both instances the dimension of the problem scales exponentially with the number $n$ of quantum particles (here qubits).
-Each quantum circuit $U_{k}$ is adaptively turned into $U_{k+1}=\exp\left(\frac{-i\eta\,dJ(0)}{d\theta_{k}}H_{k}\right)U_{k}$ by moving in a random tangent-space direction $iH_{k}$ whose corresponding projection $\frac{dJ_{k}(0)}{d\theta_{k}}=\langle\text{grad}J(U_{k}),iH_{k}U_{k}\rangle$ is measured on a quantum device, where $\eta$ denotes the learning rate and we define the cost-function value at the $k$-th iteration as $J_{k}(x)\equiv J(x,U_{k})$ where $J(x,U_{k})=\text{tr}\{A\exp(-ixH_{k})U_{k}\rho U_{k}^{*}\exp(ixH_{k})\}$ (see Sec.
-5).
-The measurement projection is done on the quantum computer, and the statistics of repeated measurements for the expectation value is done on the classical computer.
-In this way a randomized gradient flow as in Fig.
-1 is implemented.
-In this work we show that for sufficiently small $\eta$ such a (discretized) flow converges almost surely to the global optimum.
+5, where the cost function is then denoted by $J$.
 Structure and Main Results
@@
 2 we introduce the fundamental notions from Riemannian geometry to describe gradient flows (Sec.
-2.1), recollecting some well-known results on the convergence of gradient flows (Sec.
-2.2), and their algorithmic implementation as gradient descents (Sec.
+2.1), recollecting some well-known results on the convergence of such gradient flows (Secs.
+2.2 and 2.3), and their algorithmic implementation as gradient descents (Sec.
 2.4).
 In Sec.
-3 we describe a gradient descent algorithm where in each step the full gradient of a high-dimensional problem is projected on just a randomly chosen direction of the tangent space, considering both continuous and discrete distributions.
-The main main result is obtained in Secs.
-3.2 and 3.3, where we prove that a randomly projected gradient descent algorithm to a smooth Morse–Bott function (with compact sublevel sets) on a Riemannian manifold comes with two vitally beneficial properties: (i) it almost surely escapes saddle points—as is shown in Lem.
-3.11 and as a consequence (ii) it almost surely converges a local minimum—as shown in Thm.
-3.14.
-In Sec.
-4 the case of a two-dimensional saddle point is studied approximately using analytical methods and also using numerical simulation.
-We obtain good approximations for the time necessary to pass the saddle point.
-Finally Sec. 5 shows how the previous results can be applied in quantum optimization to the problem of ground state preparation.
+3 we describe a gradient descent algorithm where, in each step, the full gradient of a high-dimensional problem is projected on just a randomly chosen direction of the tangent space as illustrated in Fig.1.
+We consider both uniformly distributed444with respect to the Haar measure and discrete random directions.
+The main results are obtained in Secs.
+3.2 and 3.3, where we prove that a randomly projected gradient descent algorithm of a smooth Morse–Bott cost function (with compact sublevel sets) on a Riemannian manifold comes with two vitally beneficial properties: (i) it almost surely escapes saddle points — as is shown in Lem.
+3.14, and as a consequence (ii) it almost surely converges to a single local minimum — as shown in Thm.
+3.17.
+In Sec. 4, the case of a two-dimensional saddle point is studied using analytical methods and numerical simulation. We obtain good approximations for the time necessary to pass the saddle point.
+Finally, Sec.
+5 shows how the previous results can be applied in quantum optimization: by construction, the important problem class of ground-state preparation comes with the particular advantage that our algorithm achieves almost sure convergence to a single global minimum.
+Moreover, in these cases the algorithm can be efficiently implemented on quantum computers.
 2 Riemannian Gradient Flow and Descent
 2.1 Basic Concepts
-First we recall some basic notions and notations from Riemannian geometry.
-Let $M$ denote a finite dimensional smooth manifold of dimension $n$ with tangent and cotangent bundles ${\sf T}M$ and ${\sf T}^{*}M$, respectively.
-Moreover, let $M$ be endowed with a Riemannian metric $g$, i.e.
-a smoothly varying scalar product $g(\mathbf{x})=\langle\,\cdot,\cdot\,\rangle_{\mathbf{x}}$ on each tangent space ${\sf T}_{\mathbf{x}}M$, $\mathbf{x}\in M$.
-The pair $(M,g)$ is called a Riemannian manifold.
-Every Riemannian manifold is equipped with a Riemannian density $\mu$ induced by the metric $g$, see [35, Prop.
+[…]
+Every Riemannian manifold is equipped with a Riemannian density $\mu$ induced by the metric $\mathsf{g}$, see [42, Prop.
 2.44].
 This density can be used to integrate functions, to induce a measure on $M$ and to define sets of measure zero.
-Equivalently (but without introduction a Riemannian density), one could say that $S\subset M$ has measure zero if there is an atlas such that $S\cap{\rm dom}\,\sigma$ has Lebesgue measure zero in every chart $\sigma$.
-In particular any submanifold of dimension strictly smaller than $N$ has measure zero.
+Equivalently (but without introduction of a Riemannian density), one could say that $S\subset M$ has measure zero if there is an atlas such that $S\cap{\rm dom}\,\sigma$ has Lebesgue measure zero in every chart $\sigma$.
+In particular, any submanifold of dimension strictly smaller than $N$ has measure zero.
 2.2 Gradient Flows and Asymptotic Behaviour (I)
@@
 Then the gradient of $f$ at $\mathbf{x}\in M$, denoted by $\operatorname{grad}f(\mathbf{x})$, is the unique vector in ${\sf T}_{\mathbf{x}}M$ determined by the identity
-$df(\mathbf{x})\xi=\langle\operatorname{grad}f(\mathbf{x}),\xi\rangle_{\mathbf{x}}$ (1)
+$df(\mathbf{x})\xi=\langle\operatorname{grad}f(\mathbf{x}),\xi\rangle$ (1)
 for all $\xi\in{\sf T}_{\mathbf{x}}M$.
@@
 and its flow are referred to as the gradient system and the gradient flow of $f$, respectively.
-In local coordinates the above reads as follows: Let $\sigma:U\to\mathbb{R}^{n}$ be a chart about $\mathbf{x}\in M$, denote $\mathbf{y}=\sigma(\mathbf{x})$, and let $\tilde{f}:=f\circ\sigma^{-1}$ be the chart representation of $f$.
-Moreover let $\tilde{g}:=(\sigma^{-1})^{*}g$ be the chart representation of the metric $g$ on $\sigma(U)$, i.e.
-$\tilde{g}(\mathbf{y})=\sum_{i,j=1}^{n}\tilde{g}_{ij}(\mathbf{y})\,dy^{i}dy^{j}$, where $\tilde{G}(y):=\big{(}\tilde{g}_{ij}(\mathbf{y})\big{)}$ is a positive definite matrix varying smoothly in $\mathbf{y}\in\sigma(U)$.
-Finally let $\tilde{F}:=(\sigma^{-1})^{*}F$ be the chart representation of the gradient vector field $\operatorname{grad}f$.
+In local coordinates, the above reads as follows: Let $\sigma:U\to\mathbb{R}^{N}$ be a chart about $\mathbf{x}\in M$, denote $\mathbf{y}=\sigma(\mathbf{x})$, and let $\tilde{f}:=f\circ\sigma^{-1}$ be the chart representation of $f$.
+Moreover, let $\tilde{\mathsf{g}}:=\sigma_{*}\mathsf{g}$ be the chart representation (where $\sigma_{*}$ denotes the pushforward) of the metric $\mathsf{g}$ on $\sigma(U)$, i.e.
+$\tilde{\mathsf{g}}(\mathbf{y})=\sum_{i,j=1}^{n}\tilde{\mathsf{g}}_{ij}(\mathbf{y})\,dy^{i}dy^{j}$, where $\tilde{G}(\mathbf{y}):=\big{(}\tilde{\mathsf{g}}_{ij}(\mathbf{y})\big{)}$ is a positive definite matrix varying smoothly in $\mathbf{y}\in\sigma(U)$.
+Finally, let $\tilde{F}:=\sigma_{*}F$ be the chart representation of the gradient vector field $-\operatorname{grad}f$.
 Then it holds that
-$\tilde{F}^{j}(\mathbf{y})=\sum_{i=1}^{n}\tilde{g}^{ij}(\mathbf{y})\frac{\partial\tilde{f}}{\partial y_{i}}(\mathbf{y}),$
-where $\tilde{g}^{ij}(\mathbf{y})$ are the entries of the inverse of $\tilde{G}(\mathbf{y})$ and $\frac{\partial}{\partial y_{i}}$ denotes the $i$-th partial derivative.
+$\tilde{F}^{j}(\mathbf{y})=-\sum_{i=1}^{n}\tilde{\mathsf{g}}^{ij}(\mathbf{y})\frac{\partial\tilde{f}}{\partial y_{i}}(\mathbf{y}),$
+where $\tilde{\mathsf{g}}^{ij}(\mathbf{y})$ are the entries of the inverse of $\tilde{G}(\mathbf{y})$ and $\frac{\partial}{\partial y_{i}}$ denotes the $i$-th partial derivative.
 Any convergence analysis of gradient systems is based on the following two observations: (i) the critical points of $f$ are the equilibria of (2), (ii) $f$ constitutes a Lyapunov function for (2), i.e.
 $f$ is monotonically decreasing along solutions.
-Despite their simplicity they immediately allow several non-trivial statements about the asymptotic behavior of (2) which can be characterized by its $\omega$-limit sets
+Despite their simplicity, they immediately allow several non-trivial statements about the asymptotic behavior of (2), which can be characterized by its $\omega$-limit sets
 $\omega(\mathbf{x}_{0}):=\bigcap_{s>0}\overline{\big{\{}\varphi(t,\mathbf{x}_{0})\;|\;t>s\big{\}}}\,,$
@@
 If f has compact sublevel sets, i.e.
-if the sets $\{\mathbf{x}\in M:f(\mathbf{x})\leq c\}$ are compact for all $c\in\mathbb{R}$, then every solution of (2) exists for $t\geq 0$ and its $\omega$-limit set is a non-empty, compact and connected subset of the set of critical points of $f$.
-Although, Proposition 2.1 shows that $\omega(\mathbf{x})$ is contained in the set of critical points of $f$ it does not guarantee convergence to a single critical point and, indeed, there are smooth gradient systems the trajectories of which exhibit a non-trivial convergence behavior to the set of critical points (cf.
+if the sets $\{\mathbf{x}\in M:f(\mathbf{x})\leq c\}$ are compact for all $c\in\mathbb{R}$, then every trajectory of (2) converges to the set of critical points.
+More precisely, every trajectory of (2) exists for all $t\geq 0$ and its $\omega$-limit set is a non-empty, compact and connected subset of the set of critical points of $f$.
+Although Prop.
+2.1 shows that $\omega(\mathbf{x}_{0})$ is contained in the set of critical points of $f$, it does not guarantee convergence to a single critical point and, indeed, there are smooth gradient systems whose trajectories exhibit a non-trivial convergence behavior to the set of critical points (cf.
 “Mexican hat counter-example” by [22]).
-Yet for isolated critical points one has the following trivial consequence of Proposition 2.1.
+Yet for isolated critical points, one has the following trivial consequence of Prop.
+2.1.
 Corollary 2.2.
 If f has compact sublevel sets and all critical points are isolated, then any solution of (2) converges (for $t\to\infty$) to a single critical point of $f$.
-Moreover if additionally all saddle points are strict111see Subsec.
+Moreover, if additionally all saddle points are strict,555see Subsec.
 2.3 below then for almost all initial points (in the sense of Sec.
-2.1) the flow converges to a local minimum.
+2.1), the flow converges to a local minimum.
 Continua of critical points are much more subtle to handle. Some enhanced conditions guaranteeing convergence to a single critical point will be briefly discussed in the next subsection.
@@
 As in the Euclidian case, linearizing the vector field $F$ at its equilibria sheds light on its local stability.
-Clearly, since $F=\operatorname{grad}f$ constitutes a gradient vector field, its linearization at an equilibrium $\mathbf{x}\in M$ is given by the Hessian $\operatorname{H}_{f}(\mathbf{x})$ of $f$ at $\mathbf{x}\in M$.
+Clearly, since $F=-\operatorname{grad}f$ constitutes a gradient vector field, its linearization at an equilibrium $\mathbf{x}\in M$ is given by the Hessian $\operatorname{H}_{f}(\mathbf{x})$ of $f$ at $\mathbf{x}\in M$.
 In general, if $M$ is non-Euclidian, the computation of $\operatorname{H}_{f}(\mathbf{x})$ can be rather involved.
@@
 $\operatorname{H}_{f}(\mathbf{x}_{0})(\xi,\eta):=\sum_{i,j=1}^{n}\operatorname{\tilde{H}}(f\circ\sigma^{-1})\big{(}\sigma(\mathbf{x}_{0})\big{)}_{ij}(D\sigma(\mathbf{x}_{0})\xi)_{i}(D\sigma(\mathbf{x}_{0})\eta)_{j}\,,$ (3)
-where $\sigma$ is any chart around $\mathbf{x}_{0}\in M$ and $\operatorname{\tilde{H}}(f\circ\sigma^{-1})$ denotes the ordinary Hesse matrix of the chart representation $f\circ\sigma^{-1}$.
+where $\sigma:U\to\mathbb{R}^{n}$ is any chart around $\mathbf{x}_{0}\in M$ and $\operatorname{\tilde{H}}(f\circ\sigma^{-1})$ denotes the ordinary Hesse matrix of the chart representation $f\circ\sigma^{-1}$.
 It is straightforward to show that (3) is independent of $\sigma$.
@@
 Proposition 2.3.
-Let $(M,g)$ be a Riemannian manifold and let $\mathbf{x}$ be a critical point of $f:M\to\mathbb{R}$.
+Let $M$ be a Riemannian manifold and let $\mathbf{x}$ be a critical point of the smooth function $f:M\to\mathbb{R}$.
 If $\operatorname{H}_{f}(\mathbf{x})$ is positive definite, then $\mathbf{x}$ is a strict local minimum of $f$.
-Now the question arises whether the (asymptotic) stability of an equilibrium $\mathbf{x}\in M$ of (2) may dependent on the Riemannian metric $g$ — and the answer is surprisingly “yes”, cf.
+Now the question arises whether the (asymptotic) stability of an equilibrium $\mathbf{x}\in M$ of (2) may dependent on the Riemannian metric $\mathsf{g}$ — and the answer is surprisingly “yes”, cf.
 [67].
@@
 Theorem 2.4.
-  1. (a)
+Let $M$ be a Riemannian manifold and $f:M\to\mathbb{R}$ a smooth function. Then the following hold:
+  1. (i)
     Every strict local minimum of $f$ is a stable equilibrium of (2).
-  2. (b)
+  2. (ii)
     Every strict local minimum of $f$ which is additionally an isolated critical point is an asymptotically stable equilibrium of (2).
@@
 Handling non-isolated critical points is much more subtle.
-A first hint on how to approach this issue is obtained by Corollary 2.2 which could be restated (in a slightly weaker version) as follows:
+A first hint on how to approach this issue is obtained by Cor.
+2.2 which could be restated (in a slightly weaker version) as follows:
 For Morse functions with compact sublevel sets every solution of the corresponding gradient system converges (for $t\to\infty$) to a single critical point.
-This suggests to work with Morse–Bott functions when it comes to non-isolated critical point.
-Recall that a smooth function $f:M\to\mathbb{R}$ is a called Morse–Bott function, when the set $C$ of critical points is a closed submanifolds of $M$ such that the tangent space of ${\sf T}_{\mathbf{x}}C$ coincides with the kernel of the Hessian operator for all $\mathbf{x}\in C$.
-Note that $C$ is allowed to have several connected components with possibly different dimensions, [28, p.
-366], [48, Def.
+This suggests to work with Morse–Bott functions when it comes to non-isolated critical points.
+Recall that a smooth function $f:M\to\mathbb{R}$ is a called Morse–Bott function, if the set $C$ of critical points is a closed submanifold of $M$, such that its tangent space ${\sf T}_{\mathbf{x}}C$ coincides with the kernel of the Hessian operator for all $\mathbf{x}\in C$.
+Note that $C$ is allowed to have several connected components with possibly different dimensions, [33, p.
+366], [57, Def.
 2.41] or [7].
-Thus Morse functions are particular Morse-Bott functions, where $C$ consists of 00-dimensional mainfolds (i.e.
+Thus Morse functions are particular Morse–Bott functions, where $C$ consists only of 00-dimensional mainfolds (i.e.
 isolated points).
-Now, the above concept allows a generalization of the Morse-Palais Lemma to Morse-Bott functions which is often called Morse-Bott Lemma, see [4].
-It yields a local normal form of Morse-Bott functions near their critical points, which reads in local coordinates as follows
+Now, the above concept allows a generalization of the Morse–Palais Lemma to Morse–Bott functions which is often called Morse–Bott Lemma, see [6].
+It yields a local normal form of Morse–Bott functions near their critical points, which reads in local coordinates as follows
 $\tilde{f}(\mathbf{x},\mathbf{y},\mathbf{z})=\|\mathbf{x}\|^{2}-\|\mathbf{y}\|^{2}\,,$
-with $\mathbf{x}\in\mathbb{R}^{n_{+}}$, $\mathbf{y}\in\mathbb{R}^{n_{-}}$, and $\mathbf{z}\in\mathbb{R}^{n_{0}}$, where $n_{+}$, $n_{-}$ and $n_{0}$ are the number of positive, negative and zero eigenvalues of the Hessian operator.
-This finally allows to prove that solutions of the respective gradient systems convergence to a single critical point.
-Theorem 2.5.
-Let $f:M\to\mathbb{R}$ be a Morse–Bott function on a Riemannian manifold $(M,g)$ with compact sublevel sets.
+with $\mathbf{x}\in\mathbb{R}^{n_{+}}$, $\mathbf{y}\in\mathbb{R}^{n_{-}}$, and $\mathbf{z}\in\mathbb{R}^{n_{0}}$, where $n_{+}$, $n_{-}$ and $n_{0}$ are the numbers of positive, negative and zero eigenvalues of the Hessian operator respectively.
+Here $\|\cdot\|^{2}$ denotes the usual Euclidean norm.
+This finally allows to prove that solutions of the respective gradient system converge to a single critical point.
+Theorem 2.5 (Thm. 2.3 in [4]).
+Let $f:M\to\mathbb{R}$ be a Morse–Bott function on a Riemannian manifold $M$ with compact sublevel sets.
 Then every solution of the gradient flow (2) converges to a single critical point.
-Moreover for almost every initial point the flow converges to a local minimum.
+Moreover, for almost every initial point, the flow converges to a local minimum.
 Finally, we recall another very powerful result for analyzing the convergence of gradient systems which is based on Łojasiewicz’s celebrated gradient estimate [46].
-Let $f:M\to\mathbb{R}$ be real analytic, ${\mathbf{x}}_{0}\in M$ a critical point of $f$ and assume w.l.o.g.
-$f({\mathbf{x}}_{0})=0$.
-Then near ${\mathbf{x}}_{0}\in M$ one has the estimate
-$\|\operatorname{grad}(\mu\circ f)({\mathbf{x}})\|_{\mathbf{x}}\geq c\,,$ (4)
+Let $f:M\to\mathbb{R}$ be real analytic, ${\mathbf{x}}_{0}\in M$ a critical point of $f$, and assume without loss of generality that $f({\mathbf{x}}_{0})=0$.
+Then, near ${\mathbf{x}}_{0}$, one has the estimate
+$\|\operatorname{grad}(\mu\circ f)({\mathbf{x}})\|\geq c\,,$ (4)
 where $\mu:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a strictly increasing $C^{1}$-function and $c>0$ some positive constant, cf.
-[33, Cor.
+[39, Cor.
 1.1.25].
-In the literature, one usually find $\mu(r):=r^{1-\theta}$ with $\theta\in(0,1)$.
+In the literature, one usually finds $\mu(r):=r^{1-\theta}$ with $\theta\in(0,1)$.
 Eq.
@@
 Hence one gets the following result.
-Theorem 2.6.
-[39]. Let $(M,g)$ and $f:M\to\mathbb{R}$ be are real analytic. Then every non-empty $\omega$-limit set of (2) consists only of a single critical point.
+Theorem 2.6 ([46]).
+Let $(M,\mathrm{g})$ and $f:M\to\mathbb{R}$ be are real analytic. Then every non-empty $\omega$-limit set of (2) consists only of a single critical point.
 2.4 The Exponential Map and Numerical Gradient Descent
-Finally, we approach the problem of discretization of (2) resulting in a convergent gradient descent method. The ideas presented here can be traced back to [12, 13] and [55, 56]. Let
+Finally, we approach the problem of discretization of (2) resulting in a convergent gradient descent method (cp. Fig.1). The ideas presented here can be traced back to [15, 16] and [64, 65]. Let
 $\exp_{\mathbf{x}}:{\sf T}_{\mathbf{x}}M\to M$ (5)
@@
 $t\to\exp_{\mathbf{x}}(t\xi)$ is the unique geodesic with initial value ${\mathbf{x}}\in M$ and initial velocity $\xi\in{\sf T}_{\mathbf{x}}M$.
-Here, we assume for simplicity that $(M,g)$ is (geodesically) complete, i.e.
+Here, we assume for simplicity that $(M,\mathsf{g})$ is (geodesically) complete, i.e.
 (5) is well-defined for the entire tangent space ${\sf T}_{\mathbf{x}}M$.
@@
 which may depend on $k\in\mathbb{N}$.
+Remark 2.7.
+It should be mentioned that gradient descent algorithms are usually studied as exact algorithms, not as numerical algorithms where real numbers are represented as floating point values and arithmetic is not exact.
+Numerical effects can be important, for instance cancellation effects from computing gradients in a naive way [54].
+Nevertheless, this paper will ignore these numerical issues.
 In order to guarantee convergence of (6) to the set of critical points, it is sufficient to apply the Armijo rule, see [47].
@@
 Theorem 2.8.
-For compact Riemannian manifolds every trajectory of the discretized gradient descent (6) (with constant but small enough step size) converges to the set of critical points.
-Moreover, the set of initial points converging to strict saddle points has measure zero—in the sense of Sec.
-2.1 here.
-The second part of the above statement can be found in [37, Cor.
-6] [^3] There the authors consider Riemannian manifolds embedded in the Euclidean space and retractions instead of an intrinsic Riemannian exponential function, but this does not restrict the generality of the result, cf.
-the Nash embedding theorem [7, Thm.
+If $f$ has compact sublevel sets, every trajetory of the discretized gradient descent (6) (with constant but small enough step size) converges to the set of critical points. Moreover, if additionally
+  1. (i)
+    all critical points are isolated, then any trajectory of (6) converges to a single critical point of $f$.
+  2. (ii)
+    all saddle points are strict, then the set of initial points converging to strict saddle points has measure zero (in the sense of Sec. 2.1).
+Part (ii) of the above statement can be found in [44, Cor.
+6].
+[^7] There the authors consider Riemannian manifolds embedded in the Euclidean space and retractions instead of an intrinsic Riemannian exponential function, but this does not restrict the generality of the result, cf.
+the Nash embedding theorem [9, Thm.
 46].
-Thus, for a function with compact sublevel sets, gradient descent (6) behaves similar as its continuous counter-part, cf.
-Cor.
+Thus, for a function with compact sublevel sets, gradient descent (6) behaves similarly to its continuous counter-part, cf.
+Prop.
+2.1 and Cor.
 2.2, i.e.
 it converges almost surely to a local minimum if the step size is chosen small enough.
-Deeper results which yield convergence to a single critical point are more subtle to derive.
+Deeper results, which yield convergence to a single critical point, are more subtle to derive.
 Here we present one result in this direction which is again based on the analyticity of the cost function $f$ and on Łojasiewicz’s inequaliy.
 Theorem 2.9.
-[33] If $(M,g)$ and $f$ are real analytic, and the step sizes are chosen according to a version of the first Wolfe–Powell condition for Riemannian manifolds, then pointwise convergence holds.
-Remark 2.9.
-It should be mentioned that gradient descent algorithms are usually studied as exact algorithms, not as numerical algorithms where real numbers are represented as floating point values and arithmetic is not exact.
-Numerical effects can be important, for instance cancellation effects from computing gradients in a naive way.
-Nevertheless, this paper will ignore these numerical issues.
-Finally, in order to determine the largest admissible step size of our algorithm we need a notion of Lipschitz continuity of vector fields on Riemannian manifolds.
+[39] If $(M,\mathsf{g})$ and $f$ are real analytic, and the step sizes are chosen according to a version of the first Wolfe–Powell condition for Riemannian manifolds, then pointwise convergence holds.
+Finally, in order to determine the largest admissible step size of our algorithm, we need a notion of Lipschitz continuity of vector fields on Riemannian manifolds.
 Care has to be taken here since tangent vectors in different tangent spaces cannot be compared by default.
@@
 A natural and intrinsic way to do this is to define Lipschitz continuity via parallel transport along the unique connecting geodesic, as is done in [28].
-Here, we use the following equivalent approach via Riemannian normal coordinates which result from any orthogonal coordinate system on $T_{\mathbf{x}}M$ in combination with the Riemannian exponential map $\exp_{\mathbf{x}}$.
-Definition 1.
-Let $(M,g)$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be a smooth function.
-We say that $f$ is $\ell$-smooth around $\mathbf{x}\in M$ if there exist normal coordinates $\tilde{f}$ around $\mathbf{x}$ in which $\nabla\tilde{f}$ is $\ell$-Lipschitz.
+Definition 2.10.
+Let $(M,\mathsf{g})$ be a complete Riemannian manifold and let $X$ be a vector field on $M$.
+We say that $X$ is $\ell$-Lipschitz if for all $\mathbf{x},\mathbf{y}\in M$ with [^8] Here $r_{M}$ stands for the injectivity radius of $M$.
+$d(x,y)\leq r_{M}$ it holds that [^9] Here $\Pi_{\mathbf{y},\mathbf{x}}$ denotes the parallel transport from $\mathbf{y}$ to $\mathbf{x}$ along the unique length minimizing geodesic connecting the two points.
+$\|X(\mathbf{x})-\Pi_{\mathbf{y},\mathbf{x}}X(\mathbf{y})\|_{\mathbf{x}}\leq\ell\,d(\mathbf{x},\mathbf{y})$.
+If $f:M\to\mathbb{R}$ is a differentiable function, then we say that $f$ is $\ell$-smooth if $\operatorname{grad}(f)$ is $\ell$-Lipschitz.
+Note that this definition is slightly broader than the definition given in [28].
 3 Algorithm and Convergence
-We will study the following randomly projected gradient descent algorithm.
-Given a Riemannian manifold $M$ of dimension $N$, a smooth function $f:M\to\mathbb{R}$, and a step size $\eta>0$ (with upper bound to be determined), the update rule is given by
-$\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})),\quad g(\mathbf{x},\mathbf{u})=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle\,\mathbf{u},\quad\mathsf{S}_{\mathbf{x}_{i}}M\ni\mathbf{u}_{i}\overset{\mathrm{i.i.d}}{\sim}\mathcal{P}(\mathbf{x}_{i}),$ (7)
-where $\mathsf{S}_{\mathbf{x}}M$ denotes the unit sphere in the tangent space $\mathsf{T}_{\mathbf{x}}M$ and $\mathcal{P}(\mathbf{x}_{i})$ some probability measure on the unit sphere $\mathsf{S}_{\mathbf{x}_{i}}M$.
+We will study the following randomly projected gradient descent algorithm (cp.
+Fig.
+1).
+Given a Riemannian manifold $M$ of dimension $N$, an initial state $\mathbf{x}_{0}\in M$, a smooth function $f:M\to\mathbb{R}$, and a step size $\eta>0$ (with upper bound to be determined), the update rule is given by
+$\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})),\,\,\,g(\mathbf{x},\mathbf{u})=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle\,\mathbf{u},\,\,\,\mathsf{S}_{\mathbf{x}_{i}}M\ni\mathbf{u}_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{P}(\mathbf{x}_{i}),$ (7)
+where $\mathsf{S}_{\mathbf{x}}M$ denotes the unit sphere in the tangent space $\mathsf{T}_{\mathbf{x}}M$ and $\mathcal{P}(\mathbf{x}_{i})$ some probability distribution on the unit sphere $\mathsf{S}_{\mathbf{x}_{i}}M$.
+(Recall that “i.i.d.” stands for “independent and identically distributed”).
 Intuitively, the gradient is projected onto a randomly chosen direction $\mathbf{u}$ at each step.
-Throughout we will consider two cases.
-  • Either $\mathcal{P}(\mathbf{x}_{i})=\mathcal{U}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes the uniform (rotation invariant) probability distribution on the unit sphere, also called Haar measure;
-  • or $\mathcal{P}(\mathbf{x}_{i})=\mathcal{D}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes a finite probability distribution on the unit sphere in some sense approximating a uniform distribution.
-Let us clarify the case of a finite probability distribution.
-In this case we have $k$ continuous vector fields $\xi_{j}$ on $M$ and continuous weights $p_{j}:M\to\mathbb{R}$ with $p_{j}>0$ and $\sum_{j_{1}}^{k}p_{j}\equiv 1$.
-Then $\mathbf{u}_{i}$ takes value $\xi_{j}(\mathbf{x}_{i})$ with probability $p_{j}(\mathbf{x}_{i})$.
-Moreover we assume that at every point $\mathbf{x}$, the the tangent vectors $\xi_{1}(\mathbf{x}),\dots,\xi_{k}(\mathbf{x})$ span the entire tangent space even after removing all vectors collinear with some given direction $\xi_{j}(\mathbf{x})$.
-Our goal is to analyze the convergence behaviour of this algorithm, and in particular to show that it converges almost surely to a local minimum of $f$.
-A deterministic version of the above described algorithm was analysed in [1, Def.
+Throughout this paper we will consider two cases:
+  1. (I)
+    Either $\mathcal{P}(\mathbf{x}_{i})=\mathcal{U}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes the uniform (rotationally invariant) probability distribution on the unit sphere, also called Haar measure;
+  2. (II)
+    or $\mathcal{P}(\mathbf{x}_{i})=\mathcal{D}(\mathsf{S}_{\mathbf{x}_{i}}M)$ denotes a finite probability distribution on the unit sphere accessing all possible directions in the sense of Assumption 3.1.
+Assumption 3.1 (Discrete case).
+Given $k$ continuous vector fields $\xi_{1},\ldots,\xi_{k}$ and $k$ continuous weights $p_{1},\ldots,p_{k}$ on $M$ with $p_{j}\geq 0$ and $\sum_{j=1}^{k}p_{j}\equiv 1$, such that $p_{j}\xi_{j}/\|\xi_{j}\|$ is continuous, where the expression is set to zero when it is undefined.
+This implies for instance that $p_{j}(\mathbf{x})=0$ whenever $\xi_{j}(\mathbf{x})=0$.
+Then $\mathbf{u}_{i}$ takes value $\xi_{j}(\mathbf{x}_{i})/\|\xi_{j}(\mathbf{x}_{i})\|$ with probability $p_{j}(\mathbf{x}_{i})$.
+Moreover, we suppose that at every point $\mathbf{x}$, the tangent vectors $p_{1}(\mathbf{x})\xi_{1}(\mathbf{x}),\dots,p_{k}(\mathbf{x})\xi_{k}(\mathbf{x})$ span the entire tangent space even after removing any single vector $p_{j}(\mathbf{x})\xi_{j}(\mathbf{x})$ and, if present, all its scalar multiples.
+The following technical lemma will be useful later.
+Lemma 3.2.
+Let $K$ be a compact subset of a manifold $M$, and let $\{v_{j}\}_{j=1}^{k}$ be a set of continuous vector fields on $K$ which span the tangent space at every point.
+Then, for any continuous, non-vanishing vector field $v$ on $K$ there is some $\varepsilon>0$ such that
+${\min_{\mathbf{x}\in K}\max_{j=1,\ldots,k}|\langle v_{j}(\mathbf{x}),v(\mathbf{x})\rangle|=\varepsilon.}$
+Proof.
+The functions $\mathbf{x}\mapsto|\langle v_{j}(\mathbf{x}),v(\mathbf{x})\rangle|$ defined on $K$ are continuous and non-negative.
+Hence their maximum over $j$ is still a continuous non-negative function on $K$.
+Since the $v_{j}$ span each tangent space, this function is even strictly positive, and by compactness of $K$, there is a positive global minimum.
+∎
+Note that a related condition is used in [32, Thm. 5.2] to guarantee convergence to the set of critical points.
+Our goal is to analyze the convergence behavior of this algorithm, and in particular, to show that it converges almost surely to a local minimum of $f$.
+A deterministic version of the above described algorithm (in the sense that all “perturbations” from the classical gradient descent are deterministic and not random) was analyzed in [1, Def.
 4.2.1, Thm.
@@
 Proof.
-Using the definition of $g(\mathbf{x},\mathbf{u})$ and the fact that $\mathbf{u}$ is a unit vector we immediately obtain $\langle\operatorname{grad}f(\mathbf{x}),g(\mathbf{x},\mathbf{u})\rangle=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle^{2}=\|g(\mathbf{x},\mathbf{u})\|^{2}.$
-Note that if $\mathbf{u}$ is uniformly distributed on the sphere, then
+Using the definition of $g(\mathbf{x},\mathbf{u})$ and the fact that $\mathbf{u}$ is a unit vector, we immediately obtain $\langle\operatorname{grad}f(\mathbf{x}),g(\mathbf{x},\mathbf{u})\rangle=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle^{2}=\|g(\mathbf{x},\mathbf{u})\|^{2}.$
+Corollary 3.4.
+Note that if $\mathbf{u}$ is uniformly distributed (w.r.t. the Haar measure) on the sphere, then
 […]
@@
 Remark 3.5.
-The image of the function $u\mapsto g(\mathbf{x},\mathbf{u})$ is a hypersphere in $\mathsf{T}_{\mathbf{x}}M$ with center $\tfrac{1}{2}\operatorname{grad}f(\mathbf{x})$ and containing the origin.
-Let $\mathbf{x}\in M$ be fixed and consider the map $g_{\mathbf{x}}:\mathsf{S}_{\mathbf{x}}M\to\mathsf{T}_{\mathbf{x}}M$ given by $\mathbf{u}\mapsto g(\mathbf{x},\mathbf{u})$.
+The image of the function $\mathbf{u}\mapsto g(\mathbf{x},\mathbf{u})$ is a hypersphere in $\mathsf{T}_{\mathbf{x}}M$ with center $\tfrac{1}{2}\operatorname{grad}f(\mathbf{x})$ and intersecting the origin: Indeed, let $\mathbf{x}\in M$ be fixed and consider the map $g_{\mathbf{x}}:\mathsf{S}_{\mathbf{x}}M\to\mathsf{T}_{\mathbf{x}}M$ given by $\mathbf{u}\mapsto g(\mathbf{x},\mathbf{u})$.
 To simplify formulas, we choose an orthonormal basis $(e_{1},\ldots,e_{N})$ in $\mathsf{T}_{\mathbf{x}}M$ such that $\operatorname{grad}f(\mathbf{x})=\|\operatorname{grad}f(\mathbf{x})\|e_{N}$.
-In these coordinates the map $g_{\mathbf{x}}$ is given by
+In such coordinates the map $g_{\mathbf{x}}$ is given by
 $g_{\mathbf{x}}(u_{1},\ldots,u_{N})=u_{N}\|\operatorname{grad}f(\mathbf{x})\|\,(u_{1},\ldots,u_{N}).$
@@
 $\tilde{g}({\mathbf{x}},\mathbf{v})=\tfrac{1}{2}(\operatorname{grad}f({\mathbf{x}})+\mathbf{v}\|\operatorname{grad}f({\mathbf{x}})\|),$
-and let $v$ be distributed on the unit sphere according to a probability measure $\mathcal{V}(\mathsf{S}_{x}M)$ such that $g$ and $\tilde{g}$ induce the same probability measure.
+and let $\mathbf{v}$ be distributed on the unit sphere according to a probability measure $\mathcal{V}(\mathsf{S}_{x}M)$ such that $g$ and $\tilde{g}$ induce the same probability measure.
 The measure $\mathcal{V}$ is not uniform on the sphere, but it is still invariant under rotations preserving $\operatorname{grad}f({\mathbf{x}})$.
 We see that $v_{N}\sim 2u_{N}^{2}-1$.
+More details on the distribution of $u_{N}$ and $u_{N}^{2}$ can be found in Lems.
+A.1 and A.2.
 Note that since the standard deviation of $u_{N}^{2}$ is larger than the expected value, Chebyshev’s inequality cannot be applied to obtain useful concentration bounds.
-We will use a different method in Lemma A.2.
+For large dimension $N$, there exist good approximations of these distributions.
+In fact, the distribution of a coordinate of a uniformly random unit vector approximately follows a normal distribution, and hence it’s square approximates a $\chi^{2}_{1}$ distribution.
+See Lem.
+A.2 for a precise result.
 Corollary 3.6.
-If $\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}]=\mathds{1}/N$, which is satisfied for the Haar measure, then it holds that
+If $\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}]=\mathrm{I}_{N}/N$, which is satisfied for the Haar measure, then it holds that
 $\mathbb{E}[g(\mathbf{x},\mathbf{u})]=\frac{\operatorname{grad}f(\mathbf{x})}{N}$
@@
 Proof.
-This is a simple computation: $\mathbb{E}[g(\mathbf{x},\mathbf{u})]=\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}\operatorname{grad}f(\mathbf{x})]=\frac{1}{N}\operatorname{grad}f(\mathbf{x})$. ∎
+We compute: $\mathbb{E}[g(\mathbf{x},\mathbf{u})]=\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}\operatorname{grad}f(\mathbf{x})]=\frac{1}{N}\operatorname{grad}f(\mathbf{x})$. ∎
 The corollary shows that in this case the projected gradient is of size $\operatorname{grad}f(\mathbf{x})/N$, and hence it is rather small for large $N$.
-Intuitively this happens because in high dimensions, a uniformly random unit vector $u$ will be close to orthogonal to the gradient with high probability — after all, there are $N-1$ dimensions which are orthogonal to the gradient.
-For large dimension $N$ there exist good approximations of these distributions.
-In fact, the distribution of a coordinate of a uniformly random unit vector approximately follows a normal distribution, and hence it’s square approximates a $\chi^{2}_{1}$ distribution.
-See Lemma A.2 for a precise result.
+Intuitively, this happens because in high dimensions, a uniformly random unit vector $u$ will be close to orthogonal to the gradient with high probability — after all, there are $N-1$ dimensions which are orthogonal to the gradient.
 Now that we better understand the iteration rule, we want to understand by what amount the objective function value is likely to decrease after a certain number of iterations.
 It will be useful to use normal coordinate charts at the current point $\mathbf{x}_{i}$.
-These charts satisfy that the metric at the origin is trivial, and that geodesics passing through the origin are straight and uniformly parametrized, see for instance [35, Prop.
+These charts satisfy that the metric at the origin is trivial, and that geodesics passing through the origin are straight and uniformly parametrized, see for instance [42, Prop.
 5.24].
@@
 Lemma 3.7.
-Let $(M,g)$ be compact and let $f:M\to\mathbb{R}$ be $\ell$-smooth at $\mathbf{x}_{i}$ and with injectivity radius $\operatorname{inj}_{M}(\mathbf{x}_{i})$.
-Then for $\eta\leq\min(1/\ell,\operatorname{inj}_{M}(\mathbf{x}_{i}))$ it holds that
+Let $(M,\mathsf{g})$ be compact with injectivity radius $r_{M}$ and let $f:M\to\mathbb{R}$ be $\ell$-smooth. Then for $\eta\leq\min(1/\ell,r_{M})$ it holds that
 $f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\leq-\eta\Big{(}1-\frac{\ell\eta}{2}\Big{)}\langle\operatorname{grad}f(\mathbf{x}_{i}),g(\mathbf{x}_{i},\mathbf{u}_{i})\rangle\leq 0.$
-In particular the function value cannot increase.
+In particular the function value is surely non-increasing.
 Proof.
@@
 In particular $\mathbf{\tilde{x}}_{i}=0$.
-We may and do assume that $\nabla\tilde{f}$ is $\ell$-Lipschitz.
-Note that at the origin $\operatorname{grad}\tilde{f}(\mathbf{x}_{i})=\nabla\tilde{f}(\mathbf{\tilde{x}}_{i})$ and the coordinate representation of $\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i}))$ is given by $\mathbf{\tilde{x}}_{i+1}=-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})$.
-Then by [46, Lemma 1.2.3] and Lemma 3.1 it holds that
+Note that at the origin $\operatorname{grad}\tilde{f}(\tilde{\mathbf{x}}_{i})=\nabla\tilde{f}(\mathbf{\tilde{x}}_{i})$ and the coordinate representation of $\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i}))$ is given by $\mathbf{\tilde{x}}_{i+1}=-\eta\tilde{g}(\mathbf{x}_{i},\mathbf{u}_{i})$, where $\tilde{g}(\mathbf{x}_{i},\mathbf{u}_{i})$ is the chart representation of $g(\mathbf{x}_{i},\mathbf{u}_{i})$.
+Then, by the proof of [55, Lemma 1.2.3] and Lem.
+3.3, it holds that
 $f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})$ $=\tilde{f}(\mathbf{\tilde{x}}_{i+1})-\tilde{f}(\mathbf{\tilde{x}}_{i})$
-$\leq\langle\nabla\tilde{f}(\mathbf{\tilde{x}}_{1}),\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\rangle+\frac{\ell}{2}\|\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\|^{2}$
+$\leq\langle\nabla\tilde{f}(\mathbf{\tilde{x}}_{i}),\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\rangle+\frac{\ell}{2}\|\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\|^{2}$
 $\leq-\eta\langle\operatorname{grad}f(\mathbf{x}_{i}),g(\mathbf{x}_{i},\mathbf{u}_{i})\rangle+\frac{\ell\eta^{2}}{2}\|g(\mathbf{x}_{i},\mathbf{u}_{i})\|^{2}$
@@
 as desired. ∎
-This result shows that, under the assumptions of Corollary 3.3, we obtain
-$\mathbb{E}[f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})]\leq-\frac{\eta}{N}\Big{(}1-\frac{\ell\eta}{2}\Big{)}\|\operatorname{grad}f(\mathbf{x}_{i})\|^{2}.$
-More generally the previous result shows that the only way for the algorithm to stop improving is for the gradient to vanish: $\operatorname{grad}f(\mathbf{x}_{i})\to 0$.
+This result shows that, under the assumptions of Cor. 3.6, we obtain the conditional expectation
+$\mathbb{E}[f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\,|\,\mathbf{x}_{i}]\leq-\frac{\eta}{N}\Big{(}1-\frac{\ell\eta}{2}\Big{)}\|\operatorname{grad}f(\mathbf{x}_{i})\|^{2}.$
+More generally the previous result shows that the only way for the algorithm to stop improving is for the gradient to vanish: $\operatorname{grad}f(\mathbf{x}_{i})\to 0$ as $i\to\infty$.
 Corollary 3.8.
-If $f(\mathbf{x}_{i})\to c$ then $c$ is a critical value and $\mathbf{x}_{i}$ converges almost surely to the critical set of $c$.
+Under the assumptions of Lem. 3.7, the function values converge (surely) to some value $c$. Moreover, the following assertions hold almost surely
+  1. (i)
+    $c$ is a critical value of $f$.
+  2. (ii)
+    $\mathbf{x}_{i}$ converges to the critical set of $c$, i.e. to the set of critical point with $f(\mathbf{x})=c$.
 Proof.
-We will show that $\Pr(f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0\wedge\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0)=0$, which immediately implies that if the values $f(\mathbf{x}_{i})$ converge, then the gradients $\operatorname{grad}f(\mathbf{x}_{i})$ converge to 00 almost surely.
-Note that if $\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0$, then there is some $\varepsilon>0$ such that $\|\operatorname{grad}f(\mathbf{x}_{i})\|>\varepsilon$ on some infinite subsequence.
-Consider the case of the Haar measure.
-By the above, if $f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0$, then $u_{N}^{(i)}\to 0$ on the same subsequence.
-Since the $u_{N}^{(i)}$ are i.i.d (see Lemma A.1) and independent of the $\mathbf{x}_{i}$ the probability that this happens is 00, and this concludes the proof.
-The case of discrete probability distributions follows immediately form Lemma A.3.
-∎
+By Lem.
+3.7 the function values are non-increasing and bounded below, hence they converge to some value $c$.
+We will show that $\Pr(\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0)=0$, which immediately implies the result.
+Since $f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0$, Lem.
+3.7 shows that $\langle\operatorname{grad}f(\mathbf{x}_{i}),g(\mathbf{x}_{i},\mathbf{u}_{i})\rangle\to 0$.
+Note that if $\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0$, then there is some $m\in\mathbb{N}$ such that $\|\operatorname{grad}f(\mathbf{x}_{i})\|>\tfrac{1}{m}$ on some infinite subsequence.
+Let $z_{j}^{m}=\langle\mathbf{u}_{i(j)},\operatorname{grad}f(\mathbf{x}_{i(j)})\rangle/\|\operatorname{grad}f(\mathbf{x}_{i(j)})\|$ where $i(j)$ is the $j$-th index such that $\|\operatorname{grad}f(\mathbf{x}_{i(j)})\|\geq\tfrac{1}{m}$.
+Note that
+$\langle\operatorname{grad}f(\mathbf{x}_{i(j)}),g(\mathbf{x}_{i(j)},\mathbf{u}_{i(j)})\rangle=(z_{j}^{m}\|\operatorname{grad}f(\mathbf{x}_{i(j)})\|)^{2}\,.$
+Hence, if the gradients do not converge to 00, then, for some $m$, it must hold that $|z_{j}^{m}|\to 0$.
+We claim that there exist $10<\varepsilon,\delta<$ 1 such that for all $\mathbf{x}\in M$ with $f(\mathbf{x})\leq f(\mathbf{x}_{0})$ it holds that
+$\Pr(|\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle|\leq\varepsilon\|\operatorname{grad}f(\mathbf{x})\|)\leq\delta\,.$
+If $\mathbf{u}\sim\mathcal{U}(\mathsf{S}_{\mathbf{x}}M)$ this follows from Lem. A.1, if $\mathbf{u}\sim\mathcal{D}(\mathsf{S}_{\mathbf{x}}M)$ this follows from Lem. 3.2.
+Given $m,n,k\in\mathbb{N}$ with $1/n<\varepsilon$, and using the chain rule, Markovianity, and the claim above, we compute that
+$\Pr(\forall j>k:|z_{j}^{m}|\leq\tfrac{1}{n})=\prod_{j=k}^{\infty}\Pr(|z_{j}^{m}|\leq\tfrac{1}{n}\;\big{|}\;\forall l\in\{j,\ldots,k\}:|z_{l}^{m}|\leq\tfrac{1}{n})=0\,.$
+For any fixed value of $n$, we can use a union bound over all $m,k\geq 1$ and obtain
+$\Pr(\exists m,k\,\forall j>k:|z_{j}^{m}|\leq\tfrac{1}{n})=0\,,$
+which implies that
+$\Pr(\exists m:|z_{j}^{m}|\to 0\text{ as }j\to\infty)=0\,,$
+as desired. ∎
 3.2 Escaping Saddle Points
-The main difficulty in proving almost sure convergence of the randomly projected gradient descent algorithm is to show that it does not get stuck in a saddle point.
-Recall from Lemma 3.4 that the function value cannot increase.
-Hence, once we reach the critical value corresponding to some saddle point, without sitting in the saddle point itself, we say that we have “passed” the saddle point as now it is impossible to converge to the saddle point in question.
+The main difficulty in proving almost sure convergence to a local minimum of the randomly projected gradient descent algorithm is to show that it does not get stuck in a saddle point.
+Recall from Lem.
+3.7 that the function value cannot increase.
+Hence, once we cross the critical value corresponding to some saddle point, we say that we have passed the saddle point, as now it is impossible to converge to the saddle point in question.
 We start by proving the result in a simplified “isotropic” case, before treating the general case as a perturbation.
@@
 Consider the function $f:\mathbb{R}^{N}\to\mathbb{R}$ defined by
-$f(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{n-p-q})=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}),$
-with $1\leq p,q$ and $p+q\leq n$ as well as $a_{i},b_{j}>0$. Then $f$ is $\ell$-smooth with $\ell=2\max_{i,j}(a_{i},b_{j})$. Moreover let $0<\eta\leq\tfrac{1}{\ell}$ and
-$\theta=\arctan\Big{(}\frac{b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}}{a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}}\Big{)}\in\big{[}0,\tfrac{\pi}{2}\big{]}.$ (8)
-Then there exist constants $\varepsilon,\delta>0$ such that for all  [^4] The value $\tfrac{\pi}{3}$ is arbitrary and could be replaced by any other value in $(0,\frac{\pi}{2})$ without changing the proof.
+$f(x,y_{,}z)=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}),$
+with $x\in\mathbb{R}^{p}$, $y\in\mathbb{R}^{q}$, $z\in\mathbb{R}^{N-p-q}$, and $p,q\geq 1$ as well as $a_{i},b_{j}>0$.
+Then $f$ is $\ell$-smooth with $\ell=2\max\{a_{1},\dots,a_{p},b_{1},\dots,b_{q}\}$.
+Moreover, let $0<\eta\leq\tfrac{1}{\ell}$ and define
+$\theta:=\arctan\Big{(}\frac{b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}}{a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}}\Big{)}\in\big{[}0,\tfrac{\pi}{2}\big{]}\quad\text{for}\quad(x_{1},\dots,x_{p},y_{1},\dots,y_{q})\neq 0.$ (8)
+Then there exist constants $\varepsilon,\delta>0$ such that for all  [^10] The value $\tfrac{\pi}{3}$ is arbitrary and could be replaced by any other value in $(\tfrac{\pi}{4},\frac{\pi}{2})$ without changing the proof.
 $\theta\in[0,\tfrac{\pi}{3}]$ it holds that
@@
 A key property of this simplified setting is that the gradient is a linear vector field on $\mathbb{R}^{N}$, and so the entire situation is invariant under scaling.
-From the definition of $\theta$ it is clear that $f$ vanishes if and only if $\theta=\tfrac{\pi}{4}$.
-First note that due to the scaling invariance of the algorithm and of (8), we may focus on an initial point $\mathbf{x}_{i}$ on the unit sphere $S^{n-1}$.
+Moreover, from the definition of $\theta$ it is clear that $f$ vanishes if and only if $\theta=\tfrac{\pi}{4}$ or $(x_{1},\dots,x_{p},y_{1},\dots,y_{q})=0$.
+Due to the scaling invariance of the algorithm and of (8), we may focus on an initial point $\mathbf{x}_{i}=(x_{i},y_{i},z_{i})$ on the unit sphere $S^{N-1}$.
 After executing one step, we obtain a random improvement of our angle, namely $\theta_{i+1}-\theta_{i}$ which depends on $\mathbf{u}_{i}$ appearing in the update rule.
-Considering this improvement as a function of $\mathbf{x}_{i}$ and $\mathbf{u}_{i}$, i.e.,
-$h(\mathbf{x}_{i},\mathbf{u}_{i}):=\theta_{i}-\theta_{i},$
-is is clear that it is a continuous function.
+First let us consider this improvement in the uniform Case I as a function of $\mathbf{x}_{i}$ and $\mathbf{u}_{i}$, i.e.,
+$h(\mathbf{x}_{i},\mathbf{u}_{i}):=\theta_{i+1}-\theta_{i}\,.$
+It is clear that this is a continuous function.
 The optimal improvement achievable for a given $\mathbf{x}_{i}$ is found by maximizing over $\mathbf{u}_{i}$.
 One can show that the resulting function is still continuous in $\mathbf{x}_{i}$.
-Moreover we claim that it is strictly positive on the sphere.
+Moreover we claim that it is strictly positive on the compact set $K_{\pi/3}:=\{\mathbf{x}\in S^{n-1}:\theta(\mathbf{x})\leq\pi/3\}$.
 Indeed, if $\theta_{i}>0$, then this follows from the choice of $\eta$.
-If we perform a single (deterministic) negative gradient step with step size $\eta$, none of the coordinates will change sign, and whenever $\theta\in(0,\tfrac{\pi}{2})$ its value will strictly increase.
-If $\theta_{i}=0$, it is clear that the value of $\theta$ will almost surely strictly increase, as the set of points $\mathbf{x}_{i+1}$ with $\theta_{i+1}=0$ form a strict subspace of $\mathbb{R}^{n}$.
-Taken together this shows that there is some $\varepsilon>0$ such that the optimal improvement is at least $2\varepsilon$ for every initial $\mathbf{x}_{i+1}$.
-It follows that for any given initial state $\mathbf{x}_{i+1}$, the probability that $\theta_{i+1}-\theta_{i}\geq\varepsilon$ is strictly positive.
-Indeed, $\mathbf{x}_{i+1}$ takes values on a sphere, and $\theta(\cdot)$ is a continuous function on this sphere.
-The preimage of $(\varepsilon,+\infty)$ on the sphere is a non-empty open set and hence has a strictly positive probability.
-Moreover one can show that this probability depends continuously on the initial state $\mathbf{x}_{i}$, and again we obtain a positive minimum $\delta$ on the sphere by compactness.
-This proves that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$ independently of the initial state.
-Now consider the discrete case.
+If we perform a single (deterministic) negative gradient step with step size $\eta$, none of the coordinates will change sign, and whenever $\theta_{i}\in(0,\tfrac{\pi}{3}]$ its value will strictly increase.
+If $\theta_{i}=0$, it is clear that the value of $\theta_{i}$ will almost surely strictly increase, as the set of points $\mathbf{x}_{i+1}$ with $\theta_{i+1}=0$ form a strict subspace of $\mathbb{R}^{n}$.
+Taken together this shows that there is some $\varepsilon>0$ such that the optimal improvement is at least $2\varepsilon$ for every initial $\mathbf{x}_{i}\in K_{\pi/3}$.
+It follows that for any given initial state $\mathbf{x}_{i}\in K_{\pi/3}$, the probability that $\theta_{i+1}-\theta_{i}\geq\varepsilon$ is strictly positive.
+Indeed, $h(\mathbf{x}_{i},\cdot)$ is a continuous function on this sphere and its preimage of $(\varepsilon,+\infty)$ on the sphere is a non-empty open set and hence has a strictly positive probability (with respect to the Haar measure).
+Moreover one can show that this probability depends continuously on the initial state $\mathbf{x}_{i}$, and again we obtain a positive minimum $\delta$ on the $K_{\theta}$ by compactness.
+This proves that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$ independently of the initial state $\mathbf{x}\in K_{\pi/3}$.
+Next consider the discrete Case II.
 Again we need to prove that there exist constants $\varepsilon,\delta>0$ such that for all $\theta\in[0,\tfrac{\pi}{3}]$ it holds that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$.
 The only cases where $h(\mathbf{x},\mathbf{u})=0$ (note that it cannot be negative) with probability one is when all $v_{j}(\mathbf{x})$ for $j=1,\ldots,k$ are either orthogonal to $\operatorname{grad}(\mathbf{x})$ or collinear to $\mathbf{x}$.
-By assumption, this cannot happen.
-Hence $\max_{j}h(\mathbf{x},\mathbf{v}_{j})$ is non-negative and continuous, and on the compact set $\{\mathbf{x}\in S^{n-1}:\theta(\mathbf{x})\leq\pi/3\}$ it is even strictly positive.
-If we denote the minimum by $\varepsilon$ and $\delta=\frac{1}{k}$, then the result follows.
+By Assumption 3.1, this cannot happen.
+Hence $\max_{j}h(\mathbf{x},\mathbf{v}_{j})$ is non-negative and continuous, and on the compact set $K_{\pi/3}$ it is even strictly positive.
+This concludes the proof.
 ∎
+As a corollary, we are able to show that after sufficiently many steps, the algorithm passes the saddle point with high probability.
 Corollary 3.10.
@@
 $\Pr(\theta_{n}\geq\tfrac{\pi}{3})\geq\delta^{n}.$
-It follows that for any $m\in\mathbb{N}$
+It follows that for any $m\in\mathbb{N}$, it holds that
 $\Pr(\theta_{mn}<\tfrac{\pi}{3})\leq\Pr(\theta_{n}<\tfrac{\pi}{3})^{m}\leq(1-\delta^{n})^{m},$
@@
 Proof.
-It follows immediately from Lemma 3.6 that in $N$ steps we achieve $\Pr(\theta_{i+N}-\theta_{i}\geq N\varepsilon)\geq\delta^{N}$, or put differently, setting $N=\lceil\tfrac{\pi}{3\varepsilon}\rceil$, we get
+It follows immediately from Lem.
+3.9 that in $n$ steps we achieve $\Pr(\theta_{i+n}-\theta_{i}\geq n\varepsilon)\geq\delta^{n}$, or put differently, setting $n=\lceil\tfrac{\pi}{3\varepsilon}\rceil$, we get
 $\Pr(\theta_{n}\geq\tfrac{\pi}{3})\geq\Pr(\theta_{n}\geq n\varepsilon)\geq\delta^{n}=\delta^{\lceil\tfrac{\pi}{3\varepsilon}\rceil}>0.$
 This proves the first statement. The second statement follows from the Markovianity of the process. ∎
-In order to generalize Corollary 3.7 to the desired setting we will treat the effects of the Riemannian metric and the higher order terms of $f$ as a perturbation on the algorithm.
+In order to generalize Cor.
+3.10 to the desired setting we will treat the effects of the Riemannian metric and the higher order terms of $f$ as a perturbation on the algorithm.
 We start with the following general perturbation result.
@@
 Let $G$ be a linear vector field on $\mathbb{R}^{n}$ and let $\eta>0$ be given.
-Moreover let $P$ be a (not necessarily linear) vector field on $\mathbb{R}^{n}$ satisfying $\|P({\mathbf{x}})\|\leq C\|{\mathbf{x}}\|^{2}$ for some $C>0$.
+Moreover let $P$ be any other vector field on $\mathbb{R}^{n}$ satisfying $\|P({\mathbf{x}})\|\leq C\|{\mathbf{x}}\|^{2}$ for some $C>0$.
 Then it holds that
 $\|{\mathbf{x}}-\eta G{\mathbf{x}}-({\mathbf{y}}-\eta(G+P){\mathbf{y}})\|\leq\|\mathds{1}+\eta G\|\|{\mathbf{x}}-{\mathbf{y}}\|+C\eta\|{\mathbf{y}}\|^{2},$
-where $\|\mathds{1}+\eta G\|$ denotes the operator norm. Hence after $n$ steps, if the states remain in an $R$-ball about the origin,
-$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
+where $\|\mathds{1}+\eta G\|$ denotes the operator norm.
+Let ${\mathbf{x}}_{i+1}={\mathbf{x}}_{i}-\eta G{\mathbf{x}}_{i}$ and ${{\mathbf{y}}_{i+1}}={\mathbf{y}}_{i}-\eta(G+P){\mathbf{y}}_{i}$.
+Hence after $n$ steps, if the states remain in an $R$-ball about the origin, then
+$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i}-{\mathbf{y}}_{i}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
 Proof.
@@
 Now using the assumption that $\|{\mathbf{y}}\|\leq R$, and by iterating this result for $n$ steps, we find that
-$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|$ $\leq\|\mathds{1}+\eta G\|^{n}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+(\|\mathds{1}+\eta G\|^{n-1}+\cdots+1)CR^{2}$
-$\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
+$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|$ $\leq\|\mathds{1}+\eta G\|^{n}\|{\mathbf{x}}_{i}-{\mathbf{y}}_{i}\|+(\|\mathds{1}+\eta G\|^{n-1}+\cdots+1)CR^{2}$
+$\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i}-{\mathbf{y}}_{i}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
 This concludes the proof. ∎
-In our case, we can always choose coordinates in which the problem locally looks like the setting of Lemma 3.6 with a perturbation as in Lemma 3.8.
+In our case, we can always choose coordinates in which the problem locally looks like the setting of Lem. 3.9 with a perturbation as in Lem. 3.11.
 Lemma 3.12.
-Let $(M,g)$ be a Riemannian manifold and let $f$ be a Morse–Bott function on $M$.
-If $\mathbf{x}\in M$ is a critical point of $f$, then there exists a chart $\sigma:U\to\mathbb{R}^{n}$ about $\mathbf{x}$ such that, with $\mathbf{x}=(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{n-p-q})$
-$\sigma(\mathbf{x})$ $=0,$
-$f\circ\sigma^{-1}(\mathbf{x})$ $=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2})+\mathcal{O}(x^{2}),$
-$g_{ij}(\mathbf{x})$ $=\delta_{ij}+\mathcal{O}(x),$
+Let $(M,\mathrm{g})$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be a Morse–Bott function on $M$.
+If $\mathbf{x}_{0}\in M$ is a critical point of $f$, then there exists a chart $\sigma:U\to\mathbb{R}^{n}$ about $\mathbf{x}_{0}$ such that, with $\tilde{\mathbf{x}}=\sigma(\mathbf{x})=(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{N-p-q})$,
+$\sigma(\mathbf{x}_{0})$ $=0,$
+$f\circ\sigma^{-1}(\tilde{\mathbf{x}})$ $=c+a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2})+\mathcal{O}(\|\tilde{\mathbf{x}}\|^{2}),$
+$\mathrm{g}_{ij}(\tilde{\mathbf{x}})$ $=\delta_{ij}+\mathcal{O}(\|\tilde{\mathbf{x}}\|),$
 where $\delta_{ij}$ denotes the Kronecker symbol.
 Proof.
-Choose coordinates about $\mathbf{x}$ such that the metric is Euclidean at the $x$ (e.g. normal coordinates), then use orthogonal transformation to diagonalize the Hessian of $f$ at $x$. ∎
+Choose coordinates about $\mathbf{x}_{0}$ such that the metric is Euclidean at $\tilde{\mathbf{x}}_{0}$ (e.g.
+normal coordinates), then use an orthogonal transformation to diagonalize the Hessian of $f$ at $\tilde{\mathbf{x}}_{0}$.
+∎
 Now we can combine the three previous results.
+Corollary 3.13.
+Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function on $M$.
+For any strict saddle point ${\mathbf{x}}\in M$ with value $f({\mathbf{x}})=c$ there exists $N>0$ and a neighborhood $U$ of ${\mathbf{x}}$ such that
+${\Pr(f({\mathbf{x}}_{i+N})<c\,|\,{\mathbf{x}}_{i},\ldots,{\mathbf{x}}_{i+N}\in U)}>\delta.$
+Proof.
+This follows immediately from Cor. 3.10, Lem. 3.11 and Lem. 3.12. ∎
 Lemma 3.14.
-Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function.
-For any strict saddle point ${\mathbf{x}}\in\mathrm{Crit}(f)$ with value $f({\mathbf{x}})=c$ there exists $N>0$ and a neighborhood $U$ of ${\mathbf{x}}$ such that
-$\Pr(f({\mathbf{x}}_{i+N})<c\,|\,{\mathbf{x}}_{i}\in U)>\delta.$
+Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function on $M$.
+Further assume that $f:M\to\mathbb{R}$ has compact sublevel sets and that $\mathbf{x}_{0}\in M$ is not a critical point of $f$.
+Then the probability that $\mathbf{x}_{i}$ converges to the set of strict saddle points as $i\to\infty$ is zero.
 Proof.
-This follows immediately from Corollary 3.7, Lemma 3.8 and Lemma 3.9. ∎
-Lemma 3.11.
-Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function.
-Further assume that $f$ has compact sublevel sets and that $\mathbf{x}_{0}\in M$ is not a critical point of $f$.
-Then the probability that $\mathbf{x}_{i}$ converges to a critical submanifold of a strict saddle points as $i\to\infty$ is zero.
-Proof.
-Consider some critical value $c$ of $f$ and let $C\subset M$ be a connected component of the corresponding critical submanifold made up of strict saddle points, cf.
-[48, Def.
+Let $C$ be the set of all strict saddle points $\mathbf{x}$ with $f(\mathbf{x})\leq f(\mathbf{x}_{0})$.
+By assumption on $f$ the set $C$ is compact; actually, $C$ consists of finitely many connected compact submanifolds, cf.
+[57, Def.
 2.41].
-Note that $C$ is compact.
 We will show that the probability that $\mathbf{x}_{i}$ converges to $C$ is zero.
-Indeed let ${\mathbf{z}}\in C$ be arbitrary.
-Then, by Lemma 3.10, there exists an open set $U\subset M$ containing ${\mathbf{z}}$ such that $\Pr(f(\mathbf{x}_{i+N})<c\,|\,\mathbf{x}_{i}\in U)>\delta$ for some $\delta>0$ and $N>0$.
-Moreover, by compactness we can find a finite collection of such open sets covering $C$.
-Hence it suffices to show that the probability that $\mathbf{x}_{i}$ has an accumulation point in $C\cap U$ is zero.
-Consider a sequence $\mathbf{x}_{i}$ and let $\mathbf{z}\in C\cap U$ be an accumulation point.
-Consider two balls $B_{R_{1}}(\mathbf{z}),B_{R_{2}}(\mathbf{z})$ centered on $\mathbf{z}$ with $0<R_{2}<R_{1}$ and both balls contained in $U$.
-Choosing $R_{1}$ and $R_{2}/R_{1}$ small enough we can guarantee that $\mathbf{x}_{0}\notin B_{R_{1}}(\mathbf{z})$ and that any realization needs at least $N$ steps to traverse the spherical shell $B_{R_{1}}\setminus B_{R_{2}}$.
-Now choosing a sequence of strictly decreasing $R_{k}$ with $R_{1}$ and all $R_{2k}/R_{2k-1}$ small enough as above we see that the sequence must traverse all shells $B_{R_{2k-1}}\setminus B_{R_{2k}}$
-This shows that any sequence with accumulation point in $C\cap U$ must, infinitely often, spend $N$ consecutive steps in $U$.
-By Lemma 3.10 the probability that this happens is zero.
-Here we used the fact that by Markovianity the behaviour of each sequence of $N$ consecutive steps behaves independently.
-This concludes the proof.
+Note that by Lem.
+3.7 the function value is (surely) not increasing, and therefore we can focus on $C_{c}:=C\cap f^{-1}$$(c)$ for some $c\leq f(\mathbf{x}_{0})$.
+By compactness of $C_{c}$ we can choose for each $\mathbf{z}\in C_{c}$ balls $B_{r_{z}}(\mathbf{z})$, $B_{R_{z}}(\mathbf{z})$ around $\mathbf{z}$ with $r_{z}<R_{z}$ such that both balls are contained in a neighborhood $U(\mathbf{z})$ of $\mathbf{z}$ as in Cor.
+3.13 and such that any realization $\mathbf{x}_{i}$ needs at least $N$ steps to transverse the spherical shells $B_{R_{z}}(\mathbf{z})\setminus B_{r_{z}}(\mathbf{z})$.
+Since $C_{c}$ is compact we can find a finite covering by balls of the form $B_{r_{z}/2}(\mathbf{z})$, i.e.
+$C_{c}\subset\bigcup_{k=1}^{K}B_{r_{k}/2}(\mathbf{z}_{k})\,,$
+Next define $r_{*}:=\min_{k=1,\dots K}r_{k}/2$.
+Then, by assumption, there exists some $L\in\mathbb{N}$ such that $d(\mathbf{x}_{i},C_{c})<r_{*}$ for all $i\geq L$.
+Hence we conclude $d(\mathbf{x}_{L},\mathbf{z})\leq r_{*}$ for some $\mathbf{z}\in C_{c}$ and moreover $\mathbf{z}\in B_{r_{k}/2}(\mathbf{z}_{k})$ for some $k\in\{1,\dots,K\}$.
+Thus, the triangle inequality implies $\mathbf{x}_{L}\in B_{r_{k}}(\mathbf{z}_{k})$ and therefore we conclude $\mathbf{x}_{L},\mathbf{x}_{L+1},\dots,\mathbf{x}_{L+N}$ is in $U(\mathbf{z}_{k})$.
+Since $\mathbf{x}_{L+N}$ again satisfies $d(\mathbf{x}_{L+N},C_{c})<r_{*}$ we can repeat the above arguments and obtain $\mathbf{x}_{L+N},\mathbf{x}_{L+N+1},\dots,\mathbf{x}_{L+2N}$ are in some $U(\mathbf{z}_{k^{\prime}})$, etc.
+Combining this observation with Cor.
+3.13 we get
+$\Pr(\mathbf{x}_{i}\to C_{c}\;\text{for}\;i\to\infty)\leq\sum_{L=1}^{\infty}{\Pr}_{L}=0\,,$
+where $\Pr_{L}$ is defined by
+[…]
+Above we used the fact that by Markovianity the behavior of each sequence of $N$ consecutive steps behaves independently.
+Since there are only finitely many critical submanifolds contained in $C$ it follows that there are only finitely many $c\leq f(\mathbf{x}_{0})$ such that $C_{c}\neq\emptyset$ and thus we conclude that the probability that $\mathbf{x}_{i}$ converges to $C$ is also zero.
 ∎
 3.3 Almost Sure Convergence
-The results proven so far guarantee that the algorithm converges almost surely to the set of local minima:
+The results proven so far guarantee that the algorithm (cp. Fig. 1) converges almost surely to the set of local minima:
 Proposition 3.15.
@@
 Assume that $\mathbf{x}_{0}$ is not a critical point.
-Then for stepsize $\eta\leq 1/\ell$ the randomized gradient descent algorithm converges almost surely to the set of local minima.
+Then, for stepsize $\eta\leq 1/\ell$, the randomized gradient descent algorithm converges almost surely to the set of local minima.
 Proof.
-Since by Lemma 3.4 the function value cannot increase, and since by compactness of sublevel sets the function is lower bounded, the function value must (surely) converge to some value $f^{\star}$.
-By Corollary 3.5, $f^{\star}$ is almost surely a critical value.
-By Lemma 3.11 the algorithm almost surely does not converge towards a strict saddle point.
-Hence we converge almost surely to the set of local minima.
+Since by Lem.
+3.7 the function value cannot increase, and since by compactness of sublevel sets the function is lower bounded, the function value must (surely) converge to some value $f^{\star}$.
+By Cor.
+3.8, $f^{\star}$ is almost surely a critical value.
+By Lem.
+3.14 the algorithm almost surely does not converge towards the set of strict saddle points.
+Hence it has converge almost surely to the set of local minima.
 ∎
-In order to prove convergence to a single local minimum, we first need the following lemma.
+In order to prove convergence to a single local minimum, we first need the following technical lemma.
 Lemma 3.16.
-Let $\mathcal{E}_{i}$ be probability distributions with values in the interval $[0,1]$ and let $E_{i}\sim\mathcal{E}_{i}$ for $i\in\mathbb{N}$.
-Let $\varepsilon>0$.
+Let $\mathcal{E}_{i}$ be probability distributions over the interval $[0,1]$ and let $E_{i}\sim\mathcal{E}_{i}$ for $i\in\mathbb{N}$.
+Let $\varepsilon>0$ be given.
 We assume that there is $q>0$ such that
-$\Pr(E_{i}\geq\varepsilon\,|\,E_{1}=e_{1},\ldots,E_{i-1}=e_{i-1})\geq q$
-for all $i\in\mathbb{N}$ and for all $e_{1},\ldots,e_{i-1}\in[0,1]$. Then for any $\alpha\in(0,q\ln\tfrac{1}{1-\varepsilon})$ it holds that almost surely
+$\Pr(E_{i}\geq\varepsilon\,|\,E_{1},\ldots,E_{i-1})\geq q$
+for all $i\in\mathbb{N}$. Then for any $\alpha\in(0,q\ln\tfrac{1}{1-\varepsilon})$ it holds that almost surely
 $\prod_{i=1}^{n}(1-E_{i})\leq e^{-\alpha n}$
@@
 biased coin tosses $X_{i}\in\{0,1\}$ with $\Pr(X_{i}=1)=q$.
-By the strong law of large numbers [30, Thm.
+By the strong law of large numbers [35, Thm.
 20.1] the average $\overline{X}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}X_{i}$ converges almost surely to the expectation $q$.
 Hence for any $p\in(0,q)$, there almost surely exists some $n_{0}$ large enough such that $\overline{X}_{n}\geq p$ for all $n>n_{0}$.
-Let any value $n$, and index set $I=\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,n\}$ (ordered increasingly) be given. Then
+Let any value $n$, and index set $I=\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,n\}$ (ordered increasingly) be given. Then, using the law of total probability,
 $\Pr(E_{i}\geq\varepsilon,\,\,\forall i\in I)$ $=\Pr(E_{i_{k}}\geq\varepsilon\wedge\ldots\wedge E_{i_{1}}\geq\varepsilon)$
@@
 $\geq q^{k}$
-$=\Pr(X_{i_{k}}=1,\ldots X_{1}=1)$
+$=\Pr(X_{i_{k}}=1,\ldots X_{i_{1}}=1)$
 and thus we have the following: For all $p\in(0,q)$, there almost surely exists some $n_{0}$ large enough such that for all $n>n_{0}$ it holds that at least $pn$ of the $E_{i}$ are greater than or equal to $\varepsilon$.
@@
 with $\alpha=p\ln\tfrac{1}{1-\varepsilon}$. ∎
+Finally this allows us to prove the main result:
 Theorem 3.17.
-Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with $f$ has compact sublevel sets.
+Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with compact sublevel sets.
 Further assume that ${\mathbf{x}}_{0}$ is not a critical point of $f$.
@@
 Proof.
-The idea is the proof is as follows: first we show that almost surely the distance to the set of local minima decreases exponentially.
-This then implies that the size of each step decreases exponentially as well, and hence the algorithm converges absolutely to a local minimum.
-Proposition 3.12 shows that ${\mathbf{x}}_{i}$ converges almost surely to some connected set of local minima denoted $C$.
+The idea of the proof is as follows: first we show that almost surely the distance to the set of local minima decreases exponentially.
+Then this implies that the size of each step decreases exponentially as well, and hence the algorithm converges absolutely to a local minimum.
+Prop.
+3.15 shows that ${\mathbf{x}}_{i}$ converges almost surely to some set of local minima denoted $C$.
 Let ${\mathbf{z}}\in C$ be such a local minimum, and let $H({\mathbf{z}})$ denote the Hessian at ${\mathbf{z}}$, and let $a_{\min}({\mathbf{z}})$ denote the smallest non-zero eigenvalue of $H({\mathbf{z}})$.
+Assume w.l.o.g.
+that $f(\mathbf{z})=0$.
 By compactness and continuity $a_{\min}({\mathbf{z}})$ has a minimum on $C$, which we denote $a_{\min}$.
-In a small enough neighborhood of ${\mathbf{z}}$, Lemma 3.9 shows that
+In a sufficiently small neighborhood of ${\mathbf{z}}$, working in a chart as provided by Lem.
+3.12, we obtain
 $\|\operatorname{grad}f({\mathbf{x}})\|^{2}$ $=\sum_{j}4a_{j}(\mathbf{z})^{2}(x_{j}-z_{j})^{2}+\mathcal{O}(\|{\mathbf{x}}-\mathbf{z}\|^{3})$
@@
 $\geq 2a_{\min}(\mathbf{z})f({\mathbf{x}}).$
-Together with Lemma 3.4 we find
-$f({\mathbf{x}}_{i+1})\leq\Big{(}1-((\mathbf{u}_{i})_{N})^{2}\delta 2a_{\min}(\mathbf{z})\Big{)}f({\mathbf{x}}_{i})$
-where $\delta=\eta(1-\tfrac{\ell\eta}{2})$.
-By Lemma 3.13 this shows that there is some $\alpha>0$ such that $f({\mathbf{x}}_{i})\leq f({\mathbf{x}}_{0})e^{-\alpha i}$ almost surely for $i$ large enough. Again by Lemma 3.4 we have that
+Together with Lem. 3.7 and Cor. 3.4 we find
+$f({\mathbf{x}}_{i+1})\leq\Big{(}1-((\mathbf{u}_{i})_{N})^{2}\delta 2a_{\min}\Big{)}f({\mathbf{x}}_{i})$
+with $\delta=\eta(1-\tfrac{\ell\eta}{2})<\tfrac{1}{2a_{\min}}$.
+By Lem.
+3.16, applied to $E_{i}=((\mathbf{u}_{i})_{N})^{2}\delta a_{\min}$, this shows that there is some $\alpha>0$ such that $f({\mathbf{x}}_{i})\leq f({\mathbf{x}}_{0})e^{-\alpha i}$ almost surely for $i$ large enough.
+Again by Lem.
+3.7 we have that
 $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|\leq\sqrt{\frac{|f({\mathbf{x}}_{i+1})-f({\mathbf{x}}_{i})|}{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}.$
-This shows that
-$\sum_{i=0}^{\infty}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|\leq\frac{1}{\sqrt{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\sum_{i=0}^{\infty}\sqrt{f({\mathbf{x}}_{i})}=\sqrt{\frac{f({\mathbf{x}}_{0})}{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\frac{1}{1-e^{\alpha/2}}<\infty,$
-and so the total length of the trajectory is finite. Thus ${\mathbf{x}}_{i}$ almost surely converges absolutely to a local minimum of $f$. ∎
+This shows that for $L\in\mathbb{N}$ large enough we can assume that all $\mathbf{x}_{i}$ with $i\geq L$ are contained in a single chart of the same kind as in Lem. 3.7 and thus we conclude
+$\sum_{i=L}^{\infty}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|\leq\frac{1}{\sqrt{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\sum_{i=L}^{\infty}\sqrt{f({\mathbf{x}}_{i})}\leq\sqrt{\frac{f({\mathbf{x}}_{0})}{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\frac{e^{L\alpha/2}}{1-e^{\alpha/2}}<\infty,$
+and so the total length of the trajectory is finite. Thus ${\mathbf{x}}_{i}$ almost surely converges to a local minimum of $f$. ∎
 4 Quantitative Results in Low Dimensions
-The goal of this section is to study the hitting time $\tau$ of the sublevel set of the critical value.
-By construction, the stochastic process $x^{(i)}$ is a Markov chain, since the step $x^{(i+1)}-x^{(i)}$ depends only on the value of $x^{(i)}$.
-As such, for small enough step size $\eta$ we can approximate the algorithm by an appropriate Ito stochastic differential equation (SDE).
+The goal of this section is to study how long it takes to pass a saddle point.
+More precisely, we want to understand the hitting time $\tau$ of the sublevel set of the critical value at a saddle point.
+By construction, the discrete stochastic process ${\mathbf{x}}_{i}$ is a Markov chain, since the step ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ depends only on the value of ${\mathbf{x}}_{i}$, and not on previous steps.
+As such, for small enough step size $\eta$, we can approximate the algorithm by an appropriate Ito stochastic differential equation (SDE).
+As we will see, the accuracy of this approximation improves with smaller step size $\eta$, at the cost of slower convergence.
 4.1 Two-Dimensional Euclidean Case
-In the simplest case we consider the Morse function $f(x,y)=ax^{2}-by^{2}$ on $\mathbb{R}^{2}$ in the Euclidean metric.
-Due to the scale invariance we can normalize the state after each iteration and so we obtain a discrete-time stochastic process on the unit circle.
-For small enough stepsize $0<\eta\ll 1$ the iteration rule is approximately [^5] The formula is obtained by locally approximating the unit circle by its tangent line.
-given by
-$\Delta\phi^{i}=\phi^{i+1}-\phi^{i}=\eta\,\frac{a+b}{2}\sin(2\phi)+u_{2}\eta\sqrt{(a\cos(\phi))^{2}+(b\sin(\phi))^{2}}$ (9)
-where $\mathbf{u}=(u_{1},u_{2})$ is chosen uniformly at random from the unit circle and hence
-$\mathbb{E}[\Delta\phi]=\eta\,\frac{a+b}{2}\sin(2\phi),\quad\operatorname{Var}(\Delta\phi)=\frac{\eta\sqrt{(a\cos(\phi))^{2}+(b\sin(\phi))^{2}}}{2}$
-Using [32, Sec. 6.2] we can approximate this process using the stochastic differential equation (SDE)
-$d\Phi_{t}=\frac{a+b}{2}\sin(2\phi)dt+\frac{\sqrt{\eta((a\cos(\phi))^{2}+(b\sin(\phi))^{2})}}{2}dW_{t}$ (10)
-For simplicity we consider the simplest case $a=b=1$. Then the limiting SDE reads
-$d\Phi_{t}=\sin(2\Phi_{t})dt+\frac{\sqrt{\eta}}{2}dW_{t}$
-The approximation is visualized in Figure 3.
-Figure 3: We use the constants $\eta=0.01$, $T=10$, $a=b=1$.
-Left: $50$ realizations of the stochastic process (9) for $T/\eta=1000$ steps.
-Right: $50$ realizations of the stochastic process (10) for time $T$ and with timestep size $\Delta t=0.001$ (simulated using the Euler-Maruyama scheme [32, Sec.
-9.1]).
-In order to understand how long it takes the algorithm to pass the saddle point we are interested computing in the hitting time
-$\tau=\inf\{t>0:\Phi_{t}=\pm\frac{\pi}{4}\}.$
-Close to the origin, the SDE can be linearized as $d\Phi_{t}=2\Phi_{t}dt+\frac{\sqrt{\eta}}{2}dW_{t}$, which is a mean repelling Ornstein–Uhlenbeck process.
-Away from the origin, the deterministic part dominates and we can solve the (deterministic) ODE $\dot{\Phi}_{t}=\sin(2\Phi_{t})$.
-First we approximate the hitting time distribution of the repelling Ornstein-Uhlenbeck process.
+[…]
+[…]
+where $\mathbf{u}=(u_{1},u_{2})$ is chosen uniformly at random from the unit circle and hence using Lem. A.1 we find that
+$\mathbb{E}[\Delta\phi_{i}]=\eta\,\sin(2\phi_{i}),\quad\operatorname{Var}(\Delta\phi_{i})=\frac{\eta}{2}\,.$
+Using [38, Sec. 6.2] we can approximate this process using the stochastic differential equation (SDE)
+$d\Phi_{t}=\sin(2\Phi_{t})dt+\frac{\sqrt{\eta}}{2}dW_{t}\,,$ (10)
+where $W_{t}$ denotes a standard Brownian motion. The approximation is visualized in Fig. 2.
+Figure 2: We use the step size $\eta=0.01$, total time $T=10$, and initial state $\phi_{0}=\Phi_{0}=0$.
+Left: $50$ realizations of the discrete-time stochastic process (9) for $T/\eta=1000$ steps.
+Right: $50$ realizations of the continuous-time stochastic process (10) for time $T=10$, simulated using the Euler–Maruyama scheme [38, Sec.
+9.1] with time step size $\Delta t=0.001$.
+In order to understand how long it takes the algorithm to pass the saddle point, we are interested in computing the hitting time
+$\tau=\inf\{t\geq 0:\Phi_{t}=\pm\tfrac{\pi}{4}\},$
+where we always assume that $\Phi_{0}=0$.
+Close to 00, the SDE can be linearized as $d\Phi_{t}=2\Phi_{t}dt+\tfrac{\sqrt{\eta}}{2}dW_{t}$, which is a mean-repelling Ornstein–Uhlenbeck process.
+Away from 00, the deterministic part dominates and we can solve the (deterministic) ODE $\dot{\Phi}_{t}=\sin(2\Phi_{t})$.
+First we approximate the hitting time distribution of the mean-repelling Ornstein–Uhlenbeck process.
+More precisely, we find a lower bound on the cumulative distribution function (c.d.f.) of $\tau_{c}$, the hitting time of ±$\pm c$.
 Lemma 4.1.
@@
 Setting $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, it holds that $X_{t}\sim\mathcal{N}(0,\tilde{\sigma}(t)^{2})$.
-If we denote by $\tau_{c}$ the hitting time of ±$\pm c$ (where $c>0$), we find the lower bound $\Pr[\tau_{c}\leq t]\geq\Pr[|X_{t}|\geq c]=1+\operatorname{erf}\big{(}\frac{-c}{\tilde{\sigma}(t)\sqrt{2}}\big{)}$ where $\operatorname{erf}$ denotes the error function.
+If we denote by $\tau_{c}$ the hitting time of ±$\pm c$ (where $c>0$), we find the lower bound $\Pr[\tau_{c}\leq t]\geq\Pr[|X_{t}|\geq c]=1+\operatorname{erf}\big{(}\frac{-c}{\tilde{\sigma}(t)\sqrt{2}}\big{)}$ where $\operatorname{erf}(\cdot)$ denotes the error function.
 Proof.
@@
 […]
-where $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, see [32, Sec. 4.2 & 4.4]. It is clear that if $|X_{t}|\geq c$ then $\tau_{c}\leq t$, which implies the last statement. ∎
-The approximation is very accurate with error smaller than $\Delta t$ if $\kappa\Delta t\geq 1$ and $c\gg\sigma/\sqrt{\kappa}$, as illustrated in Figure 4.
-As in the proof of Lemma 4.1 the solution of the SDE with initial condition $X_{0}=c$ is given by
+where $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, see [38, Sec. 4.2 & 4.4]. It is clear that if $|X_{t}|\geq c$ then $\tau_{c}\leq t$, which implies the last statement. ∎
+The approximation is very accurate, as illustrated in Fig. 3. The following remark gives some details on the quality of the approximation.
+Figure 3: Lower bound for the hitting time $\tau_{c}$ of the mean-repelling Ornstein–Uhlenbeck process, as obtained in Lem.
+4.1.
+We used the constants $\kappa=2$, $\sigma=3$ and $c=10$.
+Shown are the histograms of the cumulative distribution function (c.d.f.) and probability density function (p.d.f.) of $\tau_{c}$ for the SDE $dX_{t}=\kappa X_{t}dt+\sigma dW_{t}$ computed using time steps of size $0.001$ (orange), and the analytical lower bound for the c.d.f.
+and the resulting p.d.f.
+obtained in Lem.
+4.1 (blue).
+Remark 4.2.
+As in the proof of Lem. 4.1 the solution of the SDE with initial condition $X_{0}=c$ is given by
 […]
@@
 $e^{\kappa t}c-n\tilde{\sigma}(t)\geq c\implies c\geq n\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}\frac{1}{e^{\kappa t}-1}\geq\frac{n\sigma}{\sqrt{2\kappa}}\max\Big{(}1,\sqrt{\tfrac{2}{\kappa t}}\Big{)}.$
-Hence for the error in $\tau_{c}$ to be small compared to $\Delta t$, we need $c\gg\frac{\sigma}{\sqrt{\kappa}}$ if $\kappa\Delta t$ is big, and $c\gg\frac{\sigma}{\kappa\sqrt{\Delta t}}$ if $\kappa\Delta t$ is small.
-Figure 4: We used the constants $\kappa=2$, $\sigma=3$ and $c=10$.
-Shown are the histograms of the c.d.f.
-and p.d.f of $\tau_{c}$ computed using timesteps of size $0.001$ (orange) and the lower bound for the c.d.f.
-and the resulting p.d.f.
-obtained in Lemma 4.1 (blue).
-Now if we choose $c$ small enough that the linearization of $\sin$ is accurate but large enough that the approximation of Lemma 4.1 is good (which is always possible if $\eta$ is small enough), then we can approximate $\tau$ as follows.
-We fix some value $c$ and note that the ODE $\dot{\phi}_{t}=\sin(2\phi_{t})$ with initial condition $\phi_{0}=c$ has the solution $\phi_{t}=\arctan(e^{2t}\tan(c))$, and hence the hitting time of $\pi/4$ is $\tilde{\tau}_{c}=-\tfrac{1}{2}\ln(\tan(c))$.
+Hence, for the error in $\tau_{c}$ to be small compared to some given $t>0$, we need $c\gg\frac{\sigma}{\sqrt{\kappa}}$ if $\kappa t$ is big, and $c\gg\frac{\sigma}{\kappa\sqrt{t}}$ if $\kappa t$ is small.
+Now if we choose $c$ small enough that the linearization of $\sin(\cdot)$ is accurate, but large enough that the approximation of Lem.
+4.1 is good (which is always possible if $\eta$ is small enough), then we can approximate $\tau$ as follows.
+We fix some value $c$ and note that the ODE $\dot{\phi}(t)=\sin(2\phi(t))$ with initial condition $\phi(0)=c$ has the solution $\phi(t)=\arctan(e^{2t}\tan(c))$, and hence the hitting time of $\pi/4$ is $\tilde{\tau}_{c}=-\tfrac{1}{2}\ln(\tan(c))$.
 Hence we have approximately $\tau\simeq\tau_{c}+\tilde{\tau}_{c}$.
-This is illustrated in Figure 5 for realistic values.
-Figure 5: Using $\eta=0.01$ we plot the c.d.f.
+This is illustrated in Fig.
+4 for realistic values.
+Figure 4: Using $\eta=0.05$ we plot the c.d.f.
 of the hitting time $\tau$ of $\tfrac{\pi}{4}$.
-In orange for the difference equation, in blue for the stochastic differential equation ($\Delta t=0.001$), and in red for the analytic approximation.
-5 Ground State Optimization
-In quantum information science an important task relates to finding the eigenstate (‘ground state’) and the smallest eigenvalue (‘ground-state energy’) of a Hermitian operator $A$ that describes the energy of the system (i.e.
-the Hamiltonian).
-For this setting, we leave the general notation $f:M\to\mathbb{R}$ above and specialise to the cost function over the unitary group (or orbit) $J:U\to\mathbb{R}$ we aim to minimize.
-For finding the smallest expectation value of $A$ w.r.t.
-the state $\rho$, the cost function is given by
-$J(U):=\operatorname{tr}(AU\rho U^{*})=\langle A,\operatorname{Ad}_{U}(\rho)\rangle,$ (11)
-where $\rho\in\mathfrak{pos}_{1}(n)$ is the initial state of the system and $U$ is a unitary transformation that describes a quantum circuit (writing $U^{*}$ for the complex conjugate transpose).
-By choosing an appropriate eigenbasis of $A$ we may assume that $A$ is diagonal with eigenvalues in non-increasing order.
+In orange for the difference equation (9), in blue for the stochastic differential equation (10) (with time step $\Delta t=0.001$), and in red for the analytic approximation as described above.
+5 Major Application:
+Ground-State Optimization in Quantum Physics
+Among the Riemannian optimization techniques adopted in quantum information science, hybrid quantum-classical algorithms [10] such as variational quantum algorithms (VQAs) [18, 69], have been developed to solve ground-state (smallest eigenvalue) problems [58].
+From the perspective of quantum information and computation, in VQAs, a fixed parameterized quantum circuit, i.e.
+a parameterized set of unitary transformations (usually implemented by a fixed set of unitary gates), is iteratively optimized in tandem with a classical computer to minimize a cost function whose global optima encode the solution(s) to the desired problem.
+More precisely, a quantum computer, including a measurement device, is used to output the current (e.g., expectation) value of the cost function in each iterative step.
+Interleaved, a classical computer component provides the optimization routine that determines the parameter update for the next iteration on the quantum computer.
+Since quantum computers can, in principle, handle some classically exponentially complex problems, the hope is that VQAs will be able to solve problems of large size that a solely classical algorithm would be incapable of.
+However, since the optimization problems VQAs aim to solve are highly non-linear, which can often be traced back to the quantum circuit parametrization [48, 43], the iterative search for the optimal parameters can get stuck in suboptimal solutions [12].
+To overcome this challenge, adaptive quantum algorithms have been designed.
+Instead of fixing a quantum circuit and picking a parametrization as in VQAs, adaptive algorithms successively grow a quantum circuit based on measurement data from the quantum computer.
+In this approach, the cost function is minimized while the quantum circuit is grown, see [29, 73, 49, 50].
+One of the most prominent examples of such a strategy is the so-called adapt-vqe algorithm [29], which was originally proposed to solve ground state problems in chemistry, and was further developed later to solve combinatorial optimization problems on quantum computers [68].
+Unfortunately, since typically in adaptive quantum algorithms the circuit growth is informed by gradient estimates performed by the quantum computer, adaptive quantum algorithms face similar challenges as VQAs, i.e.
+the adaptive circuit growth can get stuck when gradients (asymptotically) vanish.
+This issue has recently been addressed by identifying adaptive quantum algorithms as quantum-computer implementations of Riemannian gradient flows on the special unitary group $\text{SU}(d)$ of dimension $d=2^{n}$ where $n$ is the number of qubits of the quantum computer [73].
+Since implementing Riemannian gradient methods on quantum computers requires in general exponential resources (i.e.
+the number of gates or the number of iterations of the corresponding quantum circuit grows exponentially in $n$), [73] proposed projecting the Riemannian gradient into smaller dimensional subspaces that scale polynomially in $n$, which in turn yields scalable quantum-computer implementations.
+As already mentioned, “dimension reduction” for efficient quantum-computer implementations comes at the cost of convergence not only to local minima of the cost function but also to artificial minima induced by projections of the full Riemannian gradient into fixed subspaces.
+These artifacts can be avoided by “random projections” of the Riemannian gradient [50].
+The random projections can be implemented on quantum computers as depicted in Figs.
+1 and 5.
+More precisely, an efficient quantum-computer implementation of these random directions can be achieved through exact and approximate unitary 2-designs [25].
+Figure 5: Schematic implementation of the discretized Riemannian gradient scheme by a quantum algorithm in case the manifold is $M=\mathrm{SU}(2^{n})$.
+Thus the dimension of the problem scales exponentially with the number $n$ of quantum particles (qubits).
+Each quantum circuit $U_{k}\in\mathrm{SU}(2^{n})$ is adaptively turned into $U_{k+1}=\exp\left(\frac{-i\eta\,dJ_{k}(0)}{d\theta}H_{k}\right)U_{k}$ by moving in a random tangent-space direction $iH_{k}$ whose corresponding projection $\frac{dJ_{k}(0)}{d\theta}=\langle\text{grad}J(U_{k}),iH_{k}U_{k}\rangle$ is measured on a quantum device.
+Here $\eta$ denotes the step size and $J_{k}(\theta)$ the cost function $\theta\mapsto J_{k}(\theta):=\mathrm{tr}(A\exp(-i\theta H_{k})U_{k}\rho U_{k}^{*}\exp(i\theta H_{k}))$ of the $k$-th iteration.
+The measurement of $J_{k}(\theta)$ is done on the quantum computer, and the statistics of repeated measurements to obtain the expectation value is done on a classical computer.
+In this way a randomized gradient flow as in Fig.
+1 is implemented.
+In this work we show that for sufficiently small $\eta$ such a (discretized) flow converges almost surely to the global optimum.
+To continue the mathematical perspective of ‘ground state’ optimization, i.e.
+of finding the smallest eigenvalue (‘ground state energy’) of a Hermitian operator $A$ (Hamiltonian) that describes the energy of the system, we leave the general notation $f:M\to\mathbb{R}$ above and specialize to the cost function over the unitary group (or orbit) $J:\mathrm{U}(d)\to\mathbb{R}$ we aim to minimize.
+For finding the smallest eigenvalue or, more generally, the smallest expectation value of $A$ with respect to the initial state $\rho$, the cost function is given by
+$J(U):=\operatorname{tr}(AU\rho U^{*})=:\langle A,\operatorname{Ad}_{U}(\rho)\rangle,$ (11)
+where $\rho$ is a density matrix representing the initial state of the system and $U$ is a unitary transformation that describes a quantum circuit (writing $U^{*}$ for the complex conjugate transpose).
+By choosing an appropriate eigenbasis of $A$, we may assume that $A$ is diagonal with eigenvalues in non-increasing order.
 The function we want to optimize is of the form of Eqn.
@@
 Proposition 5.1.
-Assume that $\rho$ and $A$ are diagonal  [^6] This amounts to a shift in the function $f$ and an appropriate choice of basis.. Then the following holds.
+Assume that $\rho$ and $A$ are diagonal  [^12] This amounts to a shift in the function $J$ which results from an appropriate choice of $U:=U_{0}^{*}V_{0}$ and the identity $\operatorname{tr}(A(U_{0}^{*}V_{0})\rho(U_{0}^{*}V_{0})^{*})=\operatorname{tr}(U_{0}AU_{0}^{*}V_{0}\rho V_{0}^{*})$..
+Then the following hold.
   1. (i)
-    $U$ is a critical point of $f$ if and only if $[A,\operatorname{Ad}_{U}(\rho)]=0$. Hence the critical set of $f$ is equal to
-    $\mathsf{SU}(n)_{A}\,\mathcal{S}_{n}\,\mathsf{SU}(n)_{\rho},$
-    where $\mathcal{S}_{n}$ is the set of $n$ by $n$ permutation matrices. In particular it is a disjoint union of finitely many compact connected submanifolds.
+    $U$ is a critical point of $f$ if and only if $[A,\operatorname{Ad}_{U}(\rho)]=0$. Hence, the critical set of $J$ is equal to
+    $\mathrm{SU}(d)_{A}\,\mathcal{S}_{d}\,\mathrm{SU}(d)_{\rho},$
+    where $\mathcal{S}_{d}$ is the set of $d$ by $d$ permutation matrices and $\mathrm{SU}(d)_{A}$, $\mathrm{SU}(d)_{\rho}$ denote the stabilizer subgroups for the action of conjugation.
+In particular, it is a disjoint union of finitely many compact connected submanifolds.
   2. (ii)
-    The Hessian at a critical point $U$ is given by
+    The Hessian at a critical point $U$ is determined by
     $\frac{d^{2}}{dt^{2}}\langle A,\operatorname{Ad}_{e^{itH}}(\rho)\rangle=\sum_{i>j}2(a_{i}-a_{j})(\rho_{i}-\rho_{j})|\langle i|\operatorname{Ad}_{U}(H)|j\rangle|^{2},$
-    and in particular $f$ is Morse–Bott. [^7] Note that the function $f$ is never Morse, since the maximal torus in $\mathsf{SU}(n)$ stabilizes $A$ and $\rho$.
+    and in particular $J$ is Morse–Bott. [^13] Note that the function $J$ is never Morse, since the maximal torus in $\mathrm{SU}(d)$ stabilizes $A$ and $\rho$.
   3. (iii)
@@
 Proof.
-i: See Lemma 1.1 and Propositions 1.2 and 1.3 of [22]. ii: See Proposition 1.4 and Corollary 1.5 of [22]. iii: See Remark 1.6 of [22]. ∎
-Definition 2.
-[26][Thm.
-3] A unitary representation $\pi:G\to U(n)$, $g\mapsto\pi(g)=:U_{g}$ a finite group $G$ is called a unitary $t$-design with $t\in\mathbb{N}_{0}$ if one of the following equivalent conditions is satisfied:
+i: See Lem. 1.1 and Props. 1.2 and 1.3 of [26]. ii: See Prop. 1.4 and Cor. 1.5 of [26]. iii: See Rmk. 1.6 of [26]. ∎
+A key tool for the experimental implementation of the randomized gradient descent algorithm on quantum hardware are unitary $t$-designs, whose purpose is to efficiently implement random unitaries.
+Definition 5.2 (Thm. 3 in [30]).
+A unitary representation $\pi:G\to\mathrm{U}(d)$, $g\mapsto\pi(g)=:U_{g}$ of a finite group $G$ is called a unitary $t$-design with $t\in\mathbb{N}_{0}$ if one of the following equivalent conditions is satisfied:
   1. (a)
@@
 […]
-    where $\mathbb{C}_{t}[Z,\bar{Z}]$ denotes the set of all polynomials in $Z:=(z_{ij})_{i,j=1,\dots n}$ and $\bar{Z}:=(\bar{z}_{ij})_{i,j=1,\dots n}$, which are $t$ homogeneous in $Z$ and $\bar{Z}$., i.e.
+    where $\mathbb{C}_{t}[Z,\bar{Z}]$ denotes the set of all polynomials in $Z:=(z_{ij})_{i,j=1,\dots d}$ and $\bar{Z}:=(\bar{z}_{ij})_{i,j=1,\dots d}$, which are $t$ homogeneous in $Z$ and $\bar{Z}$., i.e.
 [^14] This is equivalent to $p(\lambda Z,\bar{\lambda}\bar{Z})=|\lambda|^{2t}p(Z,\bar{Z})$.
@@
   2. (b)
-    For all $H\in\mathbb{C}^{n^{t}\times n^{t}}$ one has the equality
+    For all $H\in\mathbb{C}^{d^{t}\times d^{t}}$ one has the equality
 […]
@@
   1. (d)
-    The tensor square representation $\pi\otimes\pi:G\to U(n)\otimes U(n)$, $g\mapsto U_{g}\otimes U_{g}$ acting on $\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ has exactly two irreducible components, namely $\mathrm{Sym}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ and $\mathrm{Alt}(\mathbb{C}^{n}\otimes\mathbb{C}^{n}):=\mathbb{C}^{n}\wedge\mathbb{C}^{n}$.
+    The tensor square representation $\pi\otimes\pi:G\to\mathrm{U}(d)\otimes\mathrm{U}(d)$, $g\mapsto U_{g}\otimes U_{g}$ acting on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ has exactly two irreducible components, namely $\mathrm{Sym}(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ and $\mathrm{Alt}(\mathbb{C}^{d}\otimes\mathbb{C}^{d}):=\mathbb{C}^{d}\wedge\mathbb{C}^{d}$.
 Remark 5.3.
-A more general notion of $t$-designs [20, 21] (see also [26]) allows any finite subset of $U(n)$ which satisfies (a), (b) or, equivalently,(c).
+A more general notion of $t$-designs [24, 25] (see also [30]) allows any finite subset of $\mathrm{U}(d)$ which satisfies (a), (b) or, equivalently,(c).
 However, since almost all known examples are constructed via group representations, we focus here on this restricted approach which is sometimes called group design.
-Moreover, the reader should note the following to facts: (i) One can assume without loss of generality that $\pi$ is faithful (i.e., one-to-one) because it is straightforward to show that $[g]\mapsto\pi(g)$, $[g]\in G/\mathrm{ker}\pi$ yields a $t$-design whenever $\pi$ is a $t$-design.
+Moreover, the reader should note the following to facts: (i) One can assume without loss of generality that $\pi$ is faithful (i.e., one-to-one) because it is straightforward to show that $[g]\mapsto\pi(g)$, $[g]\in G/\operatorname{ker}\pi$ yields a $t$-design whenever $\pi$ is a $t$-design.
 (ii) Every $t$-design is also $t^{\prime}$-design for $t^{\prime}\leq t$.
 This follows readily from condition (b) by choosing $H$ of the form $H^{\prime}\otimes I_{n}\otimes\cdots\otimes I_{n}$.
+Now we need the following technical lemma:
 Lemma 5.4.
-Let $\pi:G\to U(n)$ be any unitary representation.
+Let $\pi:G\to\mathrm{U}(d)$ be any unitary representation.
 Then $\pi$ yields a $2$-design in the sense of Def.
-2 if and only if $\pi$ acts irreducibly on $i\mathfrak{su}(n)$ via conjugation, i.e.
-the representation $\hat{\pi}:G\to U\big{(}i\mathfrak{su}(n)\big{)}$, $\pi(g)\,H:=U_{g}HU_{g}^{*}$ is irreducible.
+5.2 if and only if $\pi$ acts irreducibly on $i\mathfrak{su}(n)$ via conjugation, i.e.
+the representation $\hat{\pi}:G\to U\big{(}i\mathfrak{su}(d)\big{)}$, $\pi(g)\,H:=U_{g}HU_{g}^{*}$ is irreducible.
 Proof.
-First, $\hat{\pi}$ obviously “extends” to $\tilde{\pi}:G\to U\big{(}\mathbb{C}^{n\times n}\big{)}$, $\tilde{\pi}(g)\,A:=U_{g}AU_{g}^{-1}$.
+First, $\hat{\pi}$ obviously “extends” to $\tilde{\pi}:G\to U\big{(}\mathbb{C}^{d\times d}\big{)}$, $\tilde{\pi}(g)\,A:=U_{g}AU_{g}^{-1}$.
 Then a straightforward computation shows that the representations $\tilde{\pi}$ and $g\mapsto U_{g}\otimes\bar{U}_{g}$ are equivalent and thus for simplicity we will use the same symbol for both representations.
-Next, we investigate the commutant of the representations $U_{g}\otimes\bar{U}_{g}$ and $U_{g}\otimes U_{g}$ in $\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}\cong\mathbb{C}^{n^{2}\times n^{2}}$, i.e.
+Next, we investigate the commutant of the representations $U_{g}\otimes\bar{U}_{g}$ and $U_{g}\otimes U_{g}$ in $\mathbb{C}^{d\times d}\otimes\mathbb{C}^{d\times d}\cong\mathbb{C}^{d^{2}\times d^{2}}$, i.e.
 we are interested in the solutions $Z$ of
-$(U_{g}\otimes\bar{U}_{g})Z=Z(U_{g}\otimes\bar{U}_{g})\quad\text{for all}\quad g\in G$ (12)
+$(U_{g}\otimes\bar{U}_{g})Z=Z(U_{g}\otimes\bar{U}_{g})\quad\text{for all}\quad g\in G\,,$ (12)
 and
@@
 Eq.
-(12) and (13) are equivalent to $Ad_{U_{g}\otimes\bar{U}_{g}}(Z)=Z$ and $Ad_{U_{g}\otimes U_{g}}(Z)=Z$, respectively.
-Finally, a tedious but straightforward computation shows that the partial transposed operator $\Phi:\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}$, $A\otimes B\mapsto A\otimes B^{T}$ yield an intertwining map for $Ad_{U_{g}\otimes\bar{U}_{g}}$ and $Ad_{U_{g}\otimes U_{g}}$, i.e.
-$Ad_{U_{g}\otimes\bar{U}_{g}}\circ\Phi=\Phi\circ Ad_{U_{g}\otimes U_{g}}\,.$
+(12) and (13) are equivalent to $\operatorname{Ad}_{U_{g}\otimes\bar{U}_{g}}(Z)=Z$ and $\operatorname{Ad}_{U_{g}\otimes U_{g}}(Z)=Z$, respectively.
+Finally, a tedious but straightforward computation shows that the partial transposed operator $\Phi:\mathbb{C}^{d\times d}\otimes\mathbb{C}^{d\times d}\to\mathbb{C}^{d\times d}\otimes\mathbb{C}^{d\times d}$, $A\otimes B\mapsto A\otimes B^{T}$ yield an intertwining map for $\operatorname{Ad}_{U_{g}\otimes\bar{U}_{g}}$ and $\operatorname{Ad}_{U_{g}\otimes U_{g}}$, i.e.
+$\operatorname{Ad}_{U_{g}\otimes\bar{U}_{g}}\circ\,\Phi=\Phi\circ\operatorname{Ad}_{U_{g}\otimes U_{g}}\,.$
 This implies $\Phi$ maps the commutator of $\tilde{\pi}$ to the commutator of $\pi\otimes\pi$ and therefore the dimensions of commutators (and consequently the number of irreducible subspaces) coincide.
-Thus $\mathrm{Sym}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ and $\mathrm{Alt}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ are the only irreducible subspaces of $\pi\otimes\pi$ if and only if $\mathbb{C}\,I_{n}$ and $\mathrm{sl}_{0}(n)$ are the only irreducible subspaces to $\tilde{\pi}$ (or, equivalently, $i\mathfrak{su}(n)$ is the only irreducible subspace of $\hat{\pi}$).
+Thus $\mathrm{Sym}(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ and $\mathrm{Alt}(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ are the only irreducible subspaces of $\pi\otimes\pi$ if and only if $\mathbb{C}\,I_{d}$ and $\mathfrak{sl}_{0}(d)$ are the only irreducible subspaces to $\tilde{\pi}$ (or, equivalently, $i\mathfrak{su}(d)$ is the only irreducible subspace of $\hat{\pi}$).
 ∎
@@
 Note that the above result does not mean that the tensor square representation $\pi\otimes\pi$ and $\tilde{\pi}$ are equivalent and, in fact, they are not — as easy examples demonstrate.
-A similar result for Lie-algebra representations was elaborated on in [65] as follow-up on [64] and in particular on [17].
-Lemma 5.5.
+A similar result for Lie algebra representations was elaborated on in [75] as follow-up to [74] and in particular to [20].
+The following is the main result of this section, showing that unitary $2$-designs satisfy Assumption 3.1.
+As a consequence, the main convergence result (Thm.
+3.17) applies to the problem considered in this section.
+Proposition 5.6.
 Let $\pi:G\to\mathrm{U}(n)$ be a unitary $t$-design with $t\geq 2$ and let $H_{s}\in i\mathfrak{su}(n)$ be a traceless Hermitian matrix.
@@
 Proof.
-By Lemma 5.3 we know that the $\hat{\pi}$-invariant subspace spanned by $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ has to coincide with $i\mathfrak{su}(n)$ and thus $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ contains a basis of $i\mathfrak{su}(n)$.
-Moreover, by the defining property (b) it is straightforward to see that every $1$- and therefore also every $2$-design has to act irreducibly on $\mathbb{C}^{n}$.
+By Lem.
+5.4 we know that the $\hat{\pi}$-invariant subspace spanned by $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ has to coincide with $i\mathfrak{su}(d)$ and thus $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ contains a basis of $i\mathfrak{su}(d)$.
+Moreover, by the defining property (b) it is straightforward to see that every $1$-design, and therefore also every $2$-design, has to act irreducibly on $\mathbb{C}^{d}$.
 Thus we conclude
@@
 because the left hand side of the above equation belongs to the commutant of $\pi$ and has trace zero.
-Now if there is some $g_{0}\in G$ such that $\{U_{g}H_{s}U_{g}^{*}:g\in G\,,g\neq g_{0}\}$ does not span $i\mathfrak{su}(n)$, then $U_{g_{0}}H_{s}U_{g_{0}}^{*}$ does not lie in the span of the others, and hence $\sum_{g\in G}U_{g}H_{s}U_{g}^{*}\neq 0$, which contradicts the above.
+Now if there is some $g_{0}\in G$ such that $\{U_{g}H_{s}U_{g}^{*}:g\in G\,,g\neq g_{0}\}$ does not span $i\mathfrak{su}(d)$, then $U_{g_{0}}H_{s}U_{g_{0}}^{*}$ does not lie in the span of the others, and hence $\sum_{g\in G}U_{g}H_{s}U_{g}^{*}\neq 0$, which contradicts the above.
 ∎
+Corollary 5.7.
+The randomized gradient descent algorithm introduced in Sec.
+3, applied to the cost function (11) on the unitary group, with randomly chosen directions $\{U_{g}H_{s}U_{g}^{*}\}_{g\in G}$ as in Prop.
+5.6, converges almost surely to a global minimum.
+Moreover, by continuity, this continues to hold if $U_{g}$ is an approximate $2$-design, as long as the approximation is good enough.
+On the Orbit
+Since the map $J:\mathrm{SU}(d)\to\mathbb{R}$, which depends on the initial state $\rho$, is invariant under the stabilizer subgroup of $\rho$, denoted $\mathrm{SU}(d)_{\rho}$, the map can be factored through the corresponding quotient space.
+Equivalently, one may define the map on the unitary orbit of the initial state $\rho$.
+We denote this map by $\bar{J}:\mathrm{SU}(d)\rho\to\mathbb{R}$.
+Working with this function has some theoretical advantages, such as the reduced dimension of the state space.
+Moreover, in many cases the function is even Morse:
+Lemma 5.8.
+If $A$ is non-degenerate, then the function $\bar{J}$ is Morse.
+Proof.
+See [26, Coro. 3.7, Sec. 4]. ∎
+Despite the nice properties of $\bar{J}$, we do not use this formulation, since it is not clear how to efficiently sample random unitaries yielding directions in the tangent space of the orbit.
 6 Conclusion, Discussion, and Outlook
@@
 For both scenarios we have proven that (i) the respective algorithm almost surely escapes saddle points (Lem.
-3.11) and (ii) it almost surely converges to a local minimum (Thm.
-3.14).
-Moreover we have studied the time required by the algorithm to pass a saddle point in a simple two-dimensional setting (Sec.
+3.14) and (ii) it almost surely converges to a local minimum (Thm.
+3.17).
+Moreover, we have studied the time required by the algorithm to pass a saddle point in a simple two-dimensional setting (Sec.
 4.1).
 Note that unlike adiabatic ground-state preparation strategies (that rest on first preparing some desired initial state to find the ground state of some target Hamiltonian), our randomized Riemannian gradient flow algorithm does not require knowledge of the initial state, as the algorithm converges for almost all initial points.
-Moreover, as in the quantum setting for ground state problems the critical points just comprise saddles and global extrema, here our result implies almost sure convergence to the global minimum.
+Furthermore, as in the quantum setting for ground state problems the critical points just comprise saddles and global extrema, here our result implies almost sure convergence to the global minimum.
 However, our approach inherits a key problem already arising without projections: the overall speed of convergence to globally optimal solutions may be slow [50], since it depends on the scaling of the magnitude of the (projected) Riemannian gradient.
@@
 The algorithmic steps analyzed here are entirely modular.
-In view of future applications the (approximate) random projections w.r.t.
+In view of future applications, the (approximate) random projections w.r.t.
 the Haar measure may well be replaced by selections from other problem-adapted measures without sacrificing the convergence properties.
-In particular, one may wish to select the measures such that the projections of the gradient do not subside in numerical noise prematurely and thus circumvent the notorious barren-plateaux problem.
+In particular, one may wish to select the measures such that the projections of the gradient do not subside in numerical noise prematurely and thus circumvent the notorious barren plateaux problem.
 Moreover, one may consider random projections into subspaces by techniques of compressed sensing and shadow tomography to better approximate the full gradient flow.
 Likewise, approximations with tensor-network methods (such as, e.g., [70, 71]) may be envisaged as long as one remains on a Riemannian manifold.
-Needless to say the techniques presented here lend themselves to be taken over to higher-order quasi Newton methods (like L-BFGS) also on quantum computers.
-Thus we anticipate that the convergence guaranteed for (randomized) Riemannian gradient flows w.r.t.
-Morse–Bott type smooth cost functions will encourage wide application in and beyond quantum optimization.
+More specifically, see [52] for the connection of trotterized MERA tensor networks with Riemannian optimization (extended to quasi-Newton) of VQAs in view of hybrid quantum computing.
+Another interesting recent application, which may potentially profit from the random methods described here, is Riemannian quantum circuit optimization based on matrix product operators as suggested by [41].
+Needless to say, the randomized gradient techniques presented here in this paper lend themselves to be taken over to higher-order quasi Newton methods (like L-BFGS) as described in [63] also on quantum computers.
+Thus we anticipate that the convergence guaranteed for (randomized) Riemannian gradient flows with respect to Morse–Bott type smooth cost functions will encourage wide application in and beyond quantum optimization.
 Acknowledgements
@@
 2231328) and Knowledge Enterprise at Arizona State University.
+Conflicts of Interest
+All authors declare that they have no conflicts of interest.
 Appendix
 Appendix A Some Technical Results
-The probability distributions of $u_{N}$ and $u_{N}^{2}$ are important for understanding the behaviour of the algorithm.
-Fortunately they can be described quite easily using the beta distribution.
-Recall that the p.d.f of the beta distribution with parameters $a,b$ is given by $f(x)=\mathrm{const}\cdot x^{a-1}(1-x)^{b-1}.$
+The probability distributions of $u_{N}$ and $u_{N}^{2}$ are important for understanding the behavior of the algorithm.
+Fortunately, they can be described quite easily using the beta distribution.
+Recall that the p.d.f.
+of the beta distribution with parameters $a,b>0$ is given by $f(x)=\mathrm{const}\cdot x^{a-1}(1-x)^{b-1}$.
 Lemma A.1.
-If $u\sim\mathcal{U}(S^{N-1})$ and $u_{N}$ is the last coordinate, then
-$u_{N}\sim\mathcal{B}_{N}:=2\,\mathrm{Beta}\left(\frac{N-1}{2},\frac{N-1}{2}\right)-1,\quad u_{N}^{2}\sim\mathcal{B}_{N}^{2}:=\mathrm{Beta}\left(\frac{1}{2},\frac{N-1}{2}\right)$
+If $\mathbf{u}\sim\mathcal{U}(S^{N-1})$ and $u_{N}$ is the last coordinate, then
+$u_{N}\sim\mathcal{B}_{N}:=2\,\mathrm{Beta}\left(\frac{N-1}{2},\frac{N-1}{2}\right)-1,\quad u_{N}^{2}\sim\mathcal{B}_{N}^{2}:=\mathrm{Beta}\left(\frac{1}{2},\frac{N-1}{2}\right),$
 where $\mathrm{Beta}(a,b)$ denotes the beta distribution, and hence
@@
 For $N\geq 5$ it holds that
-$d_{K}(\sqrt{N}\mathcal{B}_{N},\mathcal{N}(0,1))\leq\frac{1}{N}$
+$d_{K}(\sqrt{N}\mathcal{B}_{N},\mathcal{N}(0,1))\leq\frac{1}{N}\,,$
 where $d_{K}$ denotes the Kolmogorov distance (the supremum distance between the cumulative distribution functions).
@@
 Proof.
-The first inequality follows from [51, p.
-20] and [50, Thm.
+The first inequality follows from [60, p.
+20] and [59, Thm.
 1.2].
@@
 Finally we compute
-$\Pr\Big{(}u^{2}_{N}\geq\frac{1}{k^{2}N}\Big{)}=2\big{(}1-\Pr\big{(}u_{N}\leq\tfrac{1}{k\sqrt{N}}\big{)}\big{)}\geq 2\big{(}1-\Phi\big{(}\tfrac{1}{k}\big{)}-\frac{1}{N}\big{)},$
+$\Pr\big{(}u^{2}_{N}\geq\tfrac{1}{k^{2}N}\big{)}=2\big{(}1-\Pr\big{(}u_{N}\leq\tfrac{1}{k\sqrt{N}}\big{)}\big{)}\geq 2\big{(}1-\Phi\big{(}\tfrac{1}{k}\big{)}-\tfrac{1}{N}\big{)},$
 proving the third inequality. ∎
-Lemma A.3.
-Let $M$ be a compact manifold and let $\{v_{j}\}_{j=1}^{k}$ be a set of continuous vector fields on $M$ which span the tangent space at every point.
-Now let any vector field $v$ on $M$ be given.
-Then it holds that there is some $\varepsilon>0$ such that
-$\min_{x\in M}\max_{i}\langle v_{i},v\rangle^{2}=\varepsilon.$
-Proof.
-The functions $\langle v_{j},v\rangle^{2}$ defined on $M$ are smooth and non-negative.
-Hence their maximum over $j$ is still a continuous non-negative function on $M$.
-Since the $v_{j}$ span each tangent space, this function is even strictly positive, and by compactness of $M$, there is a positive global minimum.
-∎